### v2

# The rationality problem for multinorm one tori

Abstract.
In this paper, we study the rationality problem for multinorm one tori, a natural generalization of norm one tori. For multinorm one tori that split over finite Galois extensions with nilpotent Galois group, we prove that stable rationality and retract rationality are equivalent, and give a criterion for the validity of the above two conditions. This generalizes the result of Endo (2011) on the rationality problem for norm one tori. To accomplish it, we introduce a generalization of character groups of multinorm one tori. Moreover, we establish systematic reduction methods originating in work of Endo (2001) for an investigation of the rationality problem for arbitrary multinorm one tori. In addition, we provide a new example for which the multinorm principle holds.
2020 Mathematics Subject Classification:
11E72, 12F20, 13A50, 14E08, 20C10, 20G15
1. Introduction
Let $k$ be a field and $k^{\operatorname{sep}}$ a fixed separable closure of $k$. In algebraic geometry, a fundamental problem is to determine whether a given algebraic variety over $k$ is rational; that is birationally equivalent to a projective space over $k$. It is also important to determine stable rationality, retract rationality, and unirationality which are weaker notions of rationality. These properties satisfy:
rational $\Rightarrow$ stably rational $\Rightarrow$ retract rational $\Rightarrow$ unirational.
We remark that the direction of the implication cannot be reversed. See Section 2 for detail.
The multinorm one tori primarily studied in this paper are algebraic tori. We recall that an algebraic torus over $k$ is a group $k$-scheme $T$ that satisfies $T\otimes_{k}k^{\operatorname{sep}}\cong(\mathbb{G}_{m,k^{\operatorname{sep}}})^{n}$ for some non-negative integer $n$. Note that an algebraic $k$-torus $T$ is always unirational (see [Vos98, p. 40, Example 21]) and Voskresenskii conjectured that stably rational tori are rational (see [Vos98, p. 68]). In this paper, we focus on studies on the stable rationality and the retract rationality.
The rationality problem is well-understood for tori of small dimensions. It is known by Voskresenskii [Vos67] that all tori of dimension $2$ are rational. Moreover, Kunyavskii [Kun90] solved the rationality problem for $3$-dimensional algebraic $k$-tori. After that, Hoshi–Yamasaki [HY17] classified algebraic $k$-tori of dimensions $4$ and $5$ that are stably rational (resp. retract rational).
In this paper, we restrict our attention to the stable rationality and the retract rationality for multinorm one tori. Let $\mathbf{K}$ be a finite étale algebra over $k$, that is, a finite product of finite separable field extensions of $k$ which are contained in $k^{\operatorname{sep}}$. Then, we set
$T_{\mathbf{K}/k}:=\operatorname{Ker}(\operatorname{N}_{\mathbf{K}/k}\colon\operatorname{Res}_{\mathbf{K}/k}\mathbb{G}_{m}\rightarrow\mathbb{G}_{m}),$
where $\operatorname{Res}_{\mathbf{K}/k}$ is the Weil restriction. We call it the multinorm one torus associated to $\mathbf{K}/k$. If $\mathbf{K}$ is a field, then $T_{\mathbf{K}/k}$ is called the norm one torus. Note that $T_{\mathbf{K}/k}$ has dimension $\dim_{k}(\mathbf{K})-1$, and splits over the Galois closure of the composite field of all factors of $\mathbf{K}$. This means that there is an isomorphism of $L$-tori $T_{\mathbf{K}/k}\otimes_{k}L\cong\mathbb{G}_{m,L}^{\dim_{k}(\mathbf{K})-1}$, where $\mathbf{K}$ is the product of finite separable field extensions $K_{1},\ldots,K_{r}$ of $k$, and $L$ is the Galois closure of the composite field of $K_{1},\ldots,K_{r}$ of $k$.
The rationality problem for norm one tori has been extensively investigated by [EM75], [CS77], [Hür84], [CS87], [LeB95], [CK00], [LL00], [Flo], [End11], [HY17], [HHY20], [HY21], [HY24] and [HY]. On the other hand, the rationality problem for multinorm one tori (especially not norm one tori) has been studied by [Hür84], [End01]. Moreover it has been treated in [CS77], [CK00], [End11]. However, less is known about it compared to the norm one tori.
As a motivation for studying the rationality problem for multinorm one tori, it is expected that this problem has applications to the rationality problem for norm one tori. Consider the norm one torus $T_{K/k}$ associated to a finite separable field extension $K/k$. Let $K^{\prime}/k$ be a finite separable field extension. By definition, there is an isomorphism of $K^{\prime}$-tori
$T_{K/k}\otimes_{k}K^{\prime}\cong T_{(K\otimes_{k}K^{\prime})/K^{\prime}}.$
Here $K\otimes_{k}K^{\prime}$ may not be a finite separable field extension of $K^{\prime}$, however it is a finite étale algebra over $K^{\prime}$. This means that multinorm one tori appear by taking base change of norm one tori. Moreover, if $T_{(K\otimes_{k}K^{\prime})/K^{\prime}}$ is not rational (resp. stably rational; retract rational) over $K^{\prime}$, then one can prove that $T_{K/k}$ is so over $k$. In fact, this approach is used in the proof of [End11, [[I]]] to obtain the non-retract rationality of some norm one tori.
The main theorem of Endo ([End01]) can be stated in our notation as follows:
Theorem 1.1 ([End01, Theorem 2]).
Let $p$ be a prime, $k$ a field, $\mathbf{K}=\prod_{i=1}^{r}K_{i}^{h_{i}}$ a finite étale $k$-algebra with $r\geq 1$ and $h_{i}\geq 1$, and $L$ the Galois closure of the composite field of $K_{1},\ldots,K_{r}$ over $k$. Assume
  • $\operatorname{Gal}(L/k)$ is an elementary $p$-abelian group;
  • $K_{i}\neq K_{j}$ for any $i,j\in\{1,\ldots,r\}$ with $i\neq j$; and
  • $K_{i}:k]=p$ for any $i\in\{1,\ldots,r\}$.
Then the following hold.
  • (1)
    In case of $p\neq 2$, the following are equivalent:
    1. (i)
      $T_{\mathbf{K}/k}$ is stably rational over $k$;
    2. (ii)
      $T_{\mathbf{K}/k}$ is retract rational over $k$;
    3. (iii)
      $r=1$.
  • (2)
    In case of $p=2$, the following are equivalent:
    1. (i)
      $T_{\mathbf{K}/k}$ is stably rational over $k$;
    2. (ii)
      $T_{\mathbf{K}/k}$ is retract rational over $k$;
    3. (iii)
      $r=1$ or $2$.
Theorem 1.1 states that the stable rationality is determined solely by the number $r$ of direct factors, which is both simple and interesting. Moreover, it naturally raises the question of how this extends to more general cases.
1.1. Main theorems
Our first main theorems generalize Theorem 1.1 to the case where $\operatorname{Gal}(L/k)$ is a $p$-group.
Theorem 1.2.
Let $p$ be an odd prime number, $k$ a field, $\mathbf{K}=\prod_{i=1}^{r}K_{i}$ a finite étale $k$-algebra with $r\geq 1$, and $L$ the Galois closure of the composite field of $K_{1},\ldots,K_{r}$ over $k$. Assume
  • $K_{i}\not\subset K_{j}$ for any $i,j\in\{1,\ldots,r\}$ with $i\neq j$; and
  • $[L:k]$ is a power of $p$.
Then the following are equivalent:
  1. (i)
    $T_{\mathbf{K}/k}$ is stably rational over $k$;
  2. (ii)
    $T_{\mathbf{K}/k}$ is retract rational over $k$;
  3. (iii)
    $r=1$ and $L$ is cyclic over $k$.
We denote by $D_{n}$ the dihedral group of order $2n$, that is,
$D_{n}:=\langle\sigma_{n},\tau_{n}\mid\sigma_{n}^{n}=\tau_{n}^{2}=1,\tau_{n}\sigma_{n}\tau_{n}^{-1}=\sigma_{n}^{-1}\rangle.$
Note that there is an isomorphism $D_{2}\cong(C_{2})^{2}$.
Theorem 1.3.
Let $k$ be a field, $\mathbf{K}=\prod_{i=1}^{r}K_{i}$ a finite étale $k$-algebra with $r\geq 1$, and $L$ the Galois closure of the composite field of $K_{1},\ldots,K_{r}$ over $k$. Assume
  • $K_{i}\not\subset K_{j}$ for any $i,j\in\{1,\ldots,r\}$ with $i\neq j$; and
  • $[L:k]$ is a power of $2$.
Then the following are equivalent:
  1. (i)
    $T_{\mathbf{K}/k}$ is stably rational over $k$;
  2. (ii)
    $T_{\mathbf{K}/k}$ is retract rational over $k$;
  3. (iii)
    $\mathbf{K}$ satisfies the condition (iii-a) or (iii-b):
    1. (iii-a)
      $r=1$ and $L$ is cyclic over $k$; or
    2. (iii-b)
      $r\geq 2$, $\operatorname{Gal}(L/k)\cong D_{2^{\nu}}$ for some $\nu\geq 1$, there is $m_{i}\in\mathbb{Z}$ so that $\operatorname{Gal}(L/K_{i})\cong\langle\sigma_{2^{\nu}}^{m_{i}}\tau_{2^{\nu}}\rangle$ for each $i$, and $\{m_{i}\bmod 2\mid 1\leq i\leq r\}=\mathbb{Z}/2\mathbb{Z}$.
Remark 1.4.
In the general setup we allow $\mathbf{K}=\prod_{i=1}^{r}K_{i}^{\,h_{i}}$ with $h_{i}\in\mathbb{Z}_{>0}$, without assuming $K_{i}\not\subset K_{j}$ for each $i\neq j$. However, by Corollary 3.13 (2) we may drop multiplicities and remove factors contained in another $K_{i}$. Accordingly, we state Theorem 1.3 in this reduced form.
The main theorem in the nilpotent case, which will be presented as Theorem 1.5, follows from Theorem 1.2 and Theorem 1.3. As we will explain later, the proof is non-trivial.
We prepare some notions to state Theorem 1.5. For a quasi-trivial torus (or, an induced torus) over $k$, we mean a $k$-torus that is isomorphic to $\operatorname{Res}_{\mathbf{K}/k}\mathbb{G}_{m}$ for some finite étale algebra $\mathbf{K}$ over $k$. The notion of quasi-trivial tori is introduced in [CS77, §2, p. 187]. In addition, a subalgebra of $\mathbf{K}$ refers to a finite product $\prod_{j=1}^{s}K^{\prime}_{i_{j}}$, where $1\leq i_{1}<\cdots<i_{s}\leq r$ are integers and $K^{\prime}_{i_{j}}$ is an intermediate field of $K_{i_{j}}/k$ for each $j\in\{1,\ldots,s\}$.
Theorem 1.5.
Let $k$ be a field, $\mathbf{K}=\prod_{i=1}^{r}K_{i}$ a finite étale $k$-algebra with $r\geq 1$, and $L$ the Galois closure of the composite field of $K_{1},\ldots,K_{r}$ over $k$. Assume
  • $\operatorname{Gal}(L/k)$ is nilpotent.
Then the following are equivalent:
  1. (i)
    $T_{\mathbf{K}/k}$ is stably rational over $k$;
  2. (ii)
    $T_{\mathbf{K}/k}$ is retract rational over $k$;
  3. (iii)
    there exists an isomorphism of $k$-tori
    $T_{\mathbf{K}/k}\times S\cong T_{\mathbf{K}^{\prime}/k}\times S^{\prime},$
    where $S$ and $S^{\prime}$ are quasi-trivial tori over $k$, and $\mathbf{K}^{\prime}=\prod_{i=1}^{r^{\prime}}K^{\prime}_{i}$ is a finite product of intermediate fields $K^{\prime}_{i}$ of $L/k$ that satisfies the condition (iii-a) or (iii-b):
    1. (iii-a)
      $r^{\prime}=1$ and $L^{\prime}/k$ which is cyclic over $k$; or
    2. (iii-b)
      $r^{\prime}=2$, $\operatorname{Gal}(L^{\prime}/k)\cong C_{m}\times D_{2^{\nu}}$ for some $m\in\mathbb{Z}_{>0}\setminus 2\mathbb{Z}$ and $\nu\in\mathbb{Z}_{>0}$, and there is $m_{i}\in\mathbb{Z}$ so that $\operatorname{Gal}(L^{\prime}/K^{\prime}_{i})\cong\langle(1,\sigma_{2^{\nu}}^{2m_{i}+i}\tau_{2^{\nu}})\rangle$ for each $i$.
    Here, $L^{\prime}$ is the Galois closure of the composite field of $K^{\prime}_{1},\ldots,K^{\prime}_{r^{\prime}}$ over $k$.
The condition (iii) implies that $T_{\mathbf{K}/k}$ and $T_{\mathbf{K}^{\prime}/k}$ are stably birationally equivalent over $k$.
Remark 1.6.
We also give an explicit construction of $\mathbf{K}^{\prime}$ in Theorem 1.5 (iii) using group theory. See Lemma 8.3.
Remark 1.7.
Theorem 1.5 in the case $r=1$ is a consequence of the results of Endo–Miyata ([EM75, Theorem 1.5, Theorem 2.3]) and Endo ([End11, Theorem 2.1]). For a norm one torus associated with a non-Galois extension $K/k$ whose Galois closure is nilpotent, it was always not retract rational ([End11, Theorem 2.1]). On the other hand, by extending the scope to multinorm tori, we obtain a new stably rational family as in (b) of Theorem 1.5.
The following is a byproduct of our proof of Theorem 1.3; in particular, we obtain stably rational examples even when the Galois group is not nilpotent.
Theorem 1.8.
Let $k$ be a field, and $\mathbf{K}=K_{1}\times K_{2}$ a finite étale algebra over $k$. Assume that
  • $K_{1}K_{2}/k$ is Galois with Galois group $D_{2m}$ for some $m\in\mathbb{Z}_{>0}$;
  • $K_{1}K_{2}:K_{i}]=2$ and $K_{i}/k$ is non-Galois for each $i\in\{1,2\}$; and
  • $K_{1}$ and $K_{2}$ are not conjugate to each other.
Then the multinorm one torus $T_{\mathbf{K}/k}$ is stably rational over $k$.
Our proofs of the main theorems are based on the study of character groups of the corresponding tori. Let $T$ be a torus over $k$. Then the character group of $T$ is defined as follows:
$X^{*}(T):=\operatorname{Hom}_{k^{\operatorname{sep}}\text{-groups}}(T\otimes_{k}k^{\operatorname{sep}},\mathbb{G}_{m,k^{\operatorname{sep}}}).$
Take a finite Galois extension $L/k$ with Galois group $G$ over which $T$ splits, which is possible in any case. Then $X^{*}(T)$ is a $G$-lattice, that is, a finitely generated free abelian group equipped with an action of $G$. On the other hand, there exist two notions for $G$-lattices: quasi-permutation and quasi-invertible. One can confirm the following:
[…]
$\Updownarrow\hskip 150.0pt\Updownarrow\hskip 60$.0pt
$X^{*}(T)\textrm{ is quasi-permutation}\Rightarrow X^{*}(T)\textrm{ is quasi-invertible. }$
Moreover, we can determine a $G$-lattice $M$ to be quasi-permutation or quasi-invertible by using a flabby resolution of $M$. This efficient technique was introduced by Endo–Miyata [EM75] and Voskresenskii [Vos69], and further developed by Colliot-Thélène–Sansuc [CS77]. The details of this will be discussed in Section 2. In particular, our theorems are reduced to the determination of a $G$-lattices $J_{G/\mathcal{H}}$ to be quasi-permutation or quasi-invertible, in the case where $G$ is a finite nilpotent group. Here, $\mathcal{H}$ is a multiset of subgroups of $G$, and $J_{G/\mathcal{H}}$ is a $G$-lattice which fits into the exact sequence
$0\rightarrow\mathbb{Z}\xrightarrow{(\varepsilon_{G/H}^{\circ})_{H\in\mathcal{H}}}\bigoplus_{H\in\mathcal{H}}\mathbb{Z}[G/H]\rightarrow J_{G/\mathcal{H}}\rightarrow 0.$
See Section 3.1 for details.
We first give a sketch of our proof of Theorem 1.2, which is in the case that $G$ is a $p$-group with $p>2$. Then, one can construct an inductive argument by considering the restriction of $J_{G/\mathcal{H}}$ by a certain maximal subgroup of $G$ (Proposition 3.20). On the other hand, for a proof of Theorem 1.3, which discusses the case that $G$ is a $2$-group, consideration of the above restriction argument alone is insufficient. To overcome this issue, we introduce a generalization of $J_{G/\mathcal{H}}$, which is denoted by $J_{G/\mathcal{H}}^{(\varphi)}$ in this paper (see Section 3.2). We use this notion to study some cases that $\mathcal{H}$ consists of normal subgroups of indices $2$ or $4$. See Section 7.2. This provides the first step of induction, similar to the proof of Theorem 1.2. Finally, Theorem 1.5 follows from Theorems 1.2, 1.3 and some reduction methods on $J_{G/\mathcal{H}}^{(\varphi)}$ that will be given in Section 3.3. Note that one of such methods, which will be presented as Proposition 3.17, is a generalization of [End11, Proposition 1.3].
1.2. Application to multinorm principle
As another motivation for studying the rationality problem for multinorm one tori, we discuss their applications to the multinorm principle.
Here we assume that $k$ is a global field. For a finite étale algebra $\mathbf{K}$ over $k$, let
$\Sha(\mathbf{K}/k):=(\operatorname{N}_{K/k}(\mathbb{A}_{\mathbf{K}}^{\times})\cap k^{\times})/\operatorname{N}_{\mathbf{K}/k}(\mathbf{K}^{\times}).$
Here, $\mathbb{A}_{\mathbf{K}}^{\times}$ is the product of the idèle groups of all the factors of $\mathbf{K}$. We say that the multinorm principle holds for $K/k$ if
$\Sha(\mathbf{K}/k)=1.$
This question has also been the subject of extensive study, for example, [Hür84], [DW14], [BLP19], [Lee22], [HHLY24], and [LOY24].
The study of $\Sha(\mathbf{K}/k)$ in the case that $\mathbf{K}$ is a field is one of the classic problems in algebraic number theory. For example, $\Sha(K/k)=1$ holds for any finite cyclic extension $K/k$. This result is known as Hasse’s norm theorem ([Has31, Satz, p. 64]).
Theorem 1.9.
Let $k$ be a global field, and $\mathbf{K}=K_{1}\times K_{2}$ a finite étale algebra over $k$. We further assume that
  • $K_{1}K_{2}$ is Galois over $k$; and
  • $\operatorname{Gal}(K_{1}K_{2}/k)\cong C_{m}\times D_{2^{\nu}}$ with $m\in\mathbb{Z}_{>0}\setminus 2\mathbb{Z}$ and $\nu\in\mathbb{Z}_{>0}$; and
  • $\operatorname{Gal}(K_{1}K_{2}/K_{i})\cong\{1\}\times\langle\sigma_{2^{\nu}}^{2m_{i}+i}\tau_{2^{\nu}}\rangle$ with $m_{i}\in\mathbb{Z}$ for each $i\in\{1,2\}$.
Then, we have $\Sha(\mathbf{K}/k)=1$.
Theorem 1.9 gives a new example on the validity of the multinorm principle. It follows from Theorem 1.5 together with the standard argument on the multinorm principle. See Section 9 for more details.
Organization of this paper.
In Section 2, we prepare some basic definitions and known results about the rationality of algebraic tori. In particular, we discuss the relationship between algebraic tori and $G$-lattices. In Section 3, we introduce the concept of multinorm tori, and provide a generalization of their corresponding $G$-lattices. Furthermore, we develop the technique of Endo [End11], and construct some reduction methods for an investigation of the rationality problem for arbitrary multinorm one tori. In Section 4, we review some properties of $p$-groups used in this paper. In Section 5, we give a proof of Theorem 1.2. In Section 6, for certain $G$-lattices, we determine whether they are stably permutation or not quasi-invertible using the theory of flabby resolutions. These lattices play a crucial role in Section 7 and Section 8. In Section 7, we give a proof of Theorem 1.3 by dividing into four steps. In Section 8, we give a proof of Theorem 1.5. That is, we give a necessary and sufficient condition for the multinorm one tori to be stably rational and retract rational in the case that split over finite Galois extensions with nilpotent Galois groups. Finally, Section 9 gives a proof of Theorem 1.9, which gives an application of our study on the rationality problem for multinorm one tori.
Acknowledgments.
The authors would like to thank Seidai Yasuda for his helpful comments on this paper. The third-named author was carried out with the support of the JSPS Research Fellowship for Young Scientists and KAKENHI Grant Number JP22KJ0041.
Notations.
Let $G$ be a finite group.
  • For a subgroup $H$ of $G$, we define $N^{G}(H)$ and $N_{G}(H)$ as follows:
    $N^{G}(H):=\bigcap_{g\in G}gHg^{-1},\quad N_{G}(H):=\{g\in G\mid gHg^{-1}=H\}.$
    Note that $N^{G}(H)$ and $N_{G}(H)$ are called the normal core of $H$ and the normalizer of $H$ in $G$, respectively.
  • A $G$-lattice refers to a finitely generated free abelian group equipped with a left action of $G$. For a $G$-lattice $M$, the dual lattice of $M$ is denoted by
    $M^{\circ}:=\operatorname{Hom}_{\mathbb{Z}}(M,\mathbb{Z}).$
    Here we define a left action of $G$ on $M^{\circ}$ as
    $G\times M^{\circ}\rightarrow M^{\circ};\,(g,f)\mapsto[x\mapsto f(g^{-1}x)].$
  • Let $M$ be a $G$-lattice. For a normal subgroup $N$ of $G$, we define $G/N$-lattices $M^{N}$ and $M^{[N]}$ as follows:
[…]
    Note that $M^{[N]}$ is isomorphic to $M_{N}/M_{N,\operatorname{tor}}$, where $M_{N}$ the $N$-coinvariant part of $M$, and $M_{N,\operatorname{tor}}$ is the torsion part of $M_{N}$. Note that $M^{N}$ and $M^{[N]}$ may not coincide in general.
2. Basic facts on the rationality of tori
Let $k$ be a field. For a non-negative integer $n$, we denote by $\mathbb{P}_{k}^{n}$ the projective space of dimension $n$ over $k$. Consider an algebraic variety $X$ over $k$. We say that $X$ is
  • rational over $k$ if it is birationally equivalent to a projective space over $k$;
  • stably rational over $k$ if $X\times_{k}\mathbb{P}_{k}^{m}$ is rational over $k$ for some $m\in\mathbb{Z}_{\geq 0}$;
  • retract rational over $k$ if there exist rational maps $f\colon\mathbb{P}_{k}^{n}\dashrightarrow X$ and $g\colon X\dashrightarrow\mathbb{P}_{k}^{n}$ with $n\in\mathbb{Z}_{\geq 0}$ such that $f\circ g=\operatorname{id}_{X}$;
  • unirational over $k$ if there is a dominant rational map from $\mathbb{P}_{k}^{n}$ to $X$ for some $n\in\mathbb{Z}_{\geq 0}$.
The notion of retract rationality was originally introduced by Saltman ([Sal84]) in the case where $k$ is infinite (see also [Kan12]). It has been generalized for all varieties over arbitrary fields by Merkurjev ([Mer17]). Note that one has implications
rational $\Rightarrow$ stably rational $\Rightarrow$ retract rational $\Rightarrow$ unirational.
In this paper, we concentrate on the case where $X$ is an algebraic torus over $k$. Fix a separable closure $k^{\operatorname{sep}}$ of $G$. Then, we can rephrase the stable rationality and the retract rationality of algebraic tori by means of $G$-lattices, where $G$ is a finite quotient of the Galois group of $k^{\operatorname{sep}}/k$. We follow the same terminology as the text book [Lor05] and the mainly referenced paper [End11].
Definition 2.1 ([End11, §1]).
Let $G$ be a finite group. We say that a $G$-lattice $M$ is
  1. (i)
    permutation if $M$ has a $\mathbb{Z}$-basis permuted by $G$, that is, $M\cong\bigoplus_{i=1}^{m}\mathbb{Z}[G/H_{i}]$ for some subgroups $H_{1},H_{2},\dots,H_{m}$;
  2. (ii)
    quasi-permutation if there is an exact sequence of $G$-lattices
    $0\rightarrow M\rightarrow R\rightarrow F\rightarrow 0,$
    where $R$ and $F$ are permutation;
  3. (iii)
    quasi-invertible if it is a direct summand of a quasi-permutation $G$-lattice.
It is not difficult to confirm that
permutation $\Rightarrow$ quasi-permutation $\Rightarrow$ quasi-invertible.
Definition 2.2 ([Lor05, §§2.3–2.5]).
Let $G$ be a finite group. We say that a $G$-lattice $M$ is
  1. (i)
    stably permutation if $M\oplus R\cong R^{\prime}$ for some permutation $G$-lattices $R$ and $R^{\prime}$;
  2. (ii)
    invertible (or, permutation projective) if it is a direct summand of a permutation $G$-lattice;
  3. (iii)
    coflabby if $H^{1}(H,M)=0$ for any subgroup $H$ of $G$;
  4. (iv)
    flabby if $M^{\circ}$ is coflabby.
It is known that the following hold:
permutation $\Rightarrow$ stably permutation $\Rightarrow$ invertible $\Rightarrow$ flabby and coflabby.
Here the rightmost implication is a consequence of [Len74, (1.1) Proposition].
Let $G$ be a finite group. We say that $G$-lattices $M_{1}$ and $M_{2}$ are similar if there exist permutation $G$-lattices $R_{1}$ and $R_{2}$ such that $M_{1}\oplus R_{1}\cong M_{2}\oplus R_{2}$. We denote by $\mathscr{S}(G)$ the set of similarity classes of $G$-lattices. For a $G$-lattice $M$, we write for $[M]$ the similarity class containing $M$. Then $\mathscr{S}(G)$ is a commutative monoid with respect to the sum
$M_{1}]+[M_{2}]:=[M_{1}\oplus M_{2}].$
By definition, for a $G$-lattice $M$, we have
  • $[M]=0$ in $\mathscr{S}(G)$ if and only if $M$ is stably permutation; and
  • $[M]$ is invertible in $\mathscr{S}(G)$ if and only if $M$ is invertible.
Definition 2.3 ([Lor05, §2.6]).
Let $G$ be a finite group, and $M$ a $G$-lattice.
  1. (i)
    A coflabby resolution of $M$ is an exact sequence of $G$-lattices
    $0\rightarrow U\rightarrow R\rightarrow M\rightarrow 0,$
    where $P$ is permutation and $F$ is coflabby.
  2. (ii)
    A flabby resolution of $M$ is an exact sequence of $G$-lattices
    $0\rightarrow M\rightarrow R\rightarrow F\rightarrow 0,$
    where $P$ is permutation and $F$ is flabby.
There is a coflabby resolution for any $G$-lattice; see [EM75, Lemma 1.1]. This implies the existence of a flabby resolution of every $G$-lattice. Moreover, if
$0\rightarrow M\rightarrow R\rightarrow F\rightarrow 0$
is a flabby resolution of $M$, then the class $[F]$ in $\mathscr{S}(G)$ depends only on $M$. In the sequel, we denote $[F]$ by $M]^{\operatorname{fl}}$. It is known that the map
$\mathscr{S}(G)\rightarrow\mathscr{S}(G);\,[M]\mapsto[M]^{\operatorname{fl}}$
is an endomorphism of monoids, i.e. $M_{1}\oplus M_{2}]^{\operatorname{fl}}=[M_{1}]^{\operatorname{fl}}+[M_{2}]^{\operatorname{fl}}$. In particular, we obtain implications as follows:
stably permutation $\Rightarrow$ quasi-permutation, invertible $\Rightarrow$ quasi-invertible.
Lemma 2.4 ([Lor05, Lemma 2.7.1 (a)]; cf. [Len74, (1.2) Proposition]).
Let $G$ be a finite group, and $F$ an invertible $G$-lattice. Then we have
$F]^{\operatorname{fl}}=-[F].$
The next proposition is not difficult; for the reader’s convenience, we include a proof since a precise reference appears to be lacking.
Proposition 2.5.
Let $G$ be a finite group.
  1. (i)
    A $G$-lattice $M$ is quasi-permutation if and only if $M]^{\operatorname{fl}}=0$.
  2. (ii)
    A $G$-lattice $M$ is quasi-invertible if and only if $M]^{\operatorname{fl}}$ is invertible.
Proof.
(i): It is clear that $M]^{\operatorname{fl}}=0$ if $M$ is quasi-permutation. For the reverse implication, assume $M]^{\operatorname{fl}}=0$. Take a flabby resolution of $M$:
$0\rightarrow M\rightarrow R\xrightarrow{\pi}F\rightarrow 0.$
By assumption, $F$ is stably permutation. Hence, there is a permutation $G$-lattice $R$ such that $F\oplus R^{\prime}$ is permutation. Moreover, the sequence
$0\rightarrow M\xrightarrow{x\mapsto(\iota(x),0)}R\oplus R^{\prime}\xrightarrow{\pi\oplus\operatorname{id}_{R^{\prime}}}F\oplus R^{\prime}\rightarrow 0$
is exact. This implies that $M$ is quasi-permutation as desired.
(ii): We first prove that $M]^{\operatorname{fl}}$ is invertible if $M$ is quasi-invertible. By assumption, there is a $G$-lattice $M^{\prime}$ such that $M\oplus M^{\prime}$ is quasi-permutation. Combining this result with (i), we obtain
$M]^{\operatorname{fl}}+[M^{\prime}]^{\operatorname{fl}}=[M\oplus M^{\prime}]^{\operatorname{fl}}=[0].$
Hence $M]^{\operatorname{fl}}$ is invertible.
On the other hand, assume that $M]^{\operatorname{fl}}$ is invertible. Take a flabby resolution
$0\rightarrow M\rightarrow R\rightarrow F\rightarrow 0$
of $M$, where $F$ is invertible by assumption. Since $F$ is invertible, we have $F]^{\operatorname{fl}}=-[F]$ according to Lemma 2.4. In particular, we obtain an equality
$M\oplus F]^{\operatorname{fl}}=[0].$
This is equivalent to the condition that $M\oplus F$ is quasi-permutation, which follows from (i). This completes the proof. $\blacksquare$
Proposition 2.6.
Let $G$ be a finite group, and $H$ its subgroup. Consider a $G$-lattice $M$. If $M$ is a quasi-permutation (resp. quasi-invertible) $G$-module, then so is as an $H$-lattice.
Lemma 2.7 ([CS77, p. 179, Lemme 2 (i), (ii), (iii)]).
Let $G$ be a finite group, and $N$ its normal subgroup. Consider a $G$-lattice $M$.
  1. (i)
    If $M$ is a permutation $G$-lattice, then $M^{N}$ is a permutation $G/N$-lattice.
  2. (ii)
    If $M$ is a coflabby $G$-lattice, then $M^{N}$ is a coflabby $G/N$-lattice.
  3. (iii)
    Let
    $0\rightarrow U\rightarrow R\rightarrow M\rightarrow 0$
    be a coflabby resolution of $M$ in $G$-lattices. Then
    $0\rightarrow U^{N}\rightarrow R^{N}\rightarrow M^{N}\rightarrow 0$
    is a coflabby resolution of $M^{N}$ in $G/N$-lattices.
Lemma 2.8 ([CS77, p. 179, Lemme 2]).
Let $G$ be a finite group, and $N$ its normal subgroup. Consider a $G/N$-lattice $M$.
  1. (i)
    The $G$-lattice $M$ is permutation (resp. stably permutation; invertible; coflabby; flabby) if and only if it is so as a $G/N$-lattice.
  2. (ii)
    Let
    $0\rightarrow U\rightarrow R\rightarrow M\rightarrow 0$
    be a coflabby resolution of $M$ in $G/N$-lattices. Then it is a coflabby resolution of $M$ in $G$-lattices.
Corollary 2.9.
Let $G$ be a finite group, and $N$ its normal subgroup. Consider a $G$-lattice $M$.
  1. (i)
    If the $G$-lattice $M$ is quasi-permutation (resp. quasi-invertible), then so is for the $G/N$-lattice $M^{[N]}$.
  2. (ii)
    If $N$ acts on $M$ trivially, then $M$ is quasi-permutation (resp. quasi-invertible) $G$-lattice if and only if it is so as an $G/N$-lattice.
Proof.
(i) follows from Lemma 2.7. (ii) is a consequence of Lemma 2.7 (iii) and Lemma 2.8. $\blacksquare$
For a $G$-lattice $M$, we define
$_{\omega}^{2}(G,M):=\operatorname{Ker}\left(H^{2}(G,M)\rightarrow\bigoplus_{g\in G}H^{2}(\langle g\rangle,M)\right).$
Proposition 2.10 ([Lor05, Proposition 2.9.2 (a)]; cf. [CS87, Proposition 9.5 (ii)], [San81, Proposition 9.8]).
Let $G$ be a finite group, and
$0\rightarrow M\rightarrow R\rightarrow F\rightarrow 0$
a flabby resolution of a $G$-lattice $M$. Then, there is an isomorphism
$_{\omega}^{2}(G,M)\cong H^{1}(G,F).$
In particular, if $M$ is quasi-invertible, then we have $_{\omega}^{2}(G,M)=0$.
For a $G$-lattice $M$, we can be summarized as follows:
$\textrm{permutation}\ \ \Rightarrow\ \$$\textrm{stably permutation}\ \ \Rightarrow\ \ \hskip 11.38109pt\textrm{invertible}\hskip 14.22636pt\ \ \Rightarrow\ \ \textrm{flabby\ and\ coflabby}$
[…]
$\textrm{quasi-permutation}\hskip 12.80373pt\Rightarrow\hskip 5.69054pt\textrm{quasi-invertible}\hskip 5.69054pt\Rightarrow\hskip 7.11317pt{}^{2}_{\omega}(G,M)=0$
[…]
$\hskip 28.45274pt\textrm{$[M]^{\operatorname{fl}}=0$}\hskip 31.29802pt\Rightarrow\hskip 2.84526pt\textrm{$[M]^{\operatorname{fl}}$ is invertible}.$
For a torus $T$ over a field $k$, we define the cocharacter module $X_{*}(T)$ and the character module $X^{*}(T)$ as
$X_{*}(T):=\operatorname{Hom}_{k^{\operatorname{sep}}\text{-groups}}(\mathbb{G}_{m,k^{\operatorname{sep}}},T\otimes_{k}k^{\operatorname{sep}}),\quad X^{*}(T):=\operatorname{Hom}_{k^{\operatorname{sep}}\text{-groups}}(T\otimes_{k}k^{\operatorname{sep}},\mathbb{G}_{m,k^{\operatorname{sep}}}).$
These are finite free abelian groups equipped with continuous actions of $\operatorname{Gal}(k^{\operatorname{sep}}/k)$ (with respect to discrete topology).
Proposition 2.11.
Let $k$ be a field, and $T$ a torus over $k$ which splits over a finite Galois extension $L$ of $k$. Define $G:=\operatorname{Gal}(L/k)$.
  1. (i)
    ([EM73, Theorem 1.6], [Lor05, Proposition 9.5.3]) The algebraic torus $T$ is stably rational over $k$ if and only if the $G$-lattice $X^{*}(T)$ is quasi-permutation.
  2. (ii)
    ([Lor05, Proposition 9.5.4], cf. [Sal84, Theorem 3.14]) The algebraic torus $T$ is retract rational over $k$ if and only if the $G$-lattice $X^{*}(T)$ is quasi-invertible.
Let $Y$ be an algebraic variety over $k$. A smooth compactification of $Y$ over $k$ refers to a proper smooth algebraic variety $X$ over $k$ that admits an open immersion $Y\hookrightarrow X$. Note that a smooth compactification of $Y$ over $k$ always exists if $k$ has characteristic 0, which is a consequence of Hironaka ([Hir64]). Moreover, Colliot-Thélène, Harari and Skorobogatov ([CHS05]) gave the existence of smooth compactifications of all tori over arbitrary fields.
Proposition 2.12 ([Vos69, Section 4, p. 1213]).
Let $k$ be a field, and $T$ an algebraic torus over $k$ which splits over a finite Galois extension $L$ over $k$. Take a smooth compactification $X$ of $T$ over $k$, and define $\overline{X}:=X\otimes_{k}k^{\operatorname{sep}}$. Then there is an exact sequence of $\operatorname{Gal}(k^{\operatorname{sep}}/k)$-lattices
$0\rightarrow X^{*}(T)\rightarrow R\rightarrow\operatorname{Pic}(\overline{X})\rightarrow 0,$
where $R$ is permutation and $\operatorname{Pic}(\overline{X})$ is flabby. In particular, we have $X^{*}(T)]^{\operatorname{fl}}=[\operatorname{Pic}(\overline{X})]$.
Combining Proposition 2.12 with Proposition 2.10, we obtain the following.
Corollary 2.13 (cf. [CS87, Proposition 9.5 (ii)], [San81, Proposition 9.8]).
Let $k$ be a field, and $T$ an algebraic torus over $k$ which splits over a finite Galois extension $L$ of $k$. Take a smooth compactification of $X$ over $k$, and put $\overline{X}:=X\otimes_{k}k^{\operatorname{sep}}$. Then there is an isomorphism
[…]
where $G:=\operatorname{Gal}(L/k)$.
3. Multinorm one tori and their character groups
3.1. Multinorm one tori
Let $k$ be a field. Consider a finite étale algebra $\mathbf{K}=\prod_{i=1}^{r}K_{i}$ over $k$, that is, a finite product of finite separable subextensions of $k^{\operatorname{sep}}$. The multinorm one torus associated to $\mathbf{K}/k$ is defined as
$T_{\mathbf{K}/k}=\operatorname{Ker}(\operatorname{N}_{\mathbf{K}/k}\colon\operatorname{Res}_{\mathbf{K}/k}\mathbb{G}_{m}\rightarrow\mathbb{G}_{m}).$
Let $G$ be a finite group, and $H$ a subgroup of $G$. Then one has a surjection
$\varepsilon_{G/H}\colon\mathbb{Z}[G/H]\rightarrow\mathbb{Z};\,\sum_{g\in G/H}a_{g}g\mapsto\sum_{g\in G/H}a_{g},$
which is called the augmentation map. Moreover, the dual of $\varepsilon_{G/H}^{\circ}$ coincides with the homomorphism
$\varepsilon_{G/H}^{\circ}\colon\mathbb{Z}\rightarrow\mathbb{Z}[G/H];\,1\mapsto\sum_{g\in G/H}g.$
On the other hand, consider a finite group $G$ and subgroups $H^{\prime}\subset H$ of $G$. Then the homomorphisms of $H$-lattices
$\varepsilon_{H/H^{\prime}}\colon\mathbb{Z}[H/H^{\prime}]\rightarrow\mathbb{Z},\quad\varepsilon_{H/H^{\prime}}^{\circ}\colon\mathbb{Z}\rightarrow\mathbb{Z}[H/H^{\prime}]$
induce homomorphisms of $G$-lattices
$\operatorname{Ind}_{H}^{G}\varepsilon_{H/H^{\prime}}\colon\mathbb{Z}[G/H^{\prime}]\rightarrow\mathbb{Z}[G/H],\quad\operatorname{Ind}_{H}^{G}\varepsilon_{H/H^{\prime}}^{\circ}\colon\mathbb{Z}[G/H]\rightarrow\mathbb{Z}[G/H^{\prime}].$
For a multiset $\mathcal{H}$ of subgroups of $G$, we define a $G$-module $I_{G/\mathcal{H}}$ by an exact sequence
$0\rightarrow I_{G/\mathcal{H}}\rightarrow\bigoplus_{H\in\mathcal{H}}\mathbb{Z}[G/H]\xrightarrow{(\varepsilon_{G/H})_{H\in\mathcal{H}}}\mathbb{Z}\rightarrow 0.$
Furthermore, we define $J_{G/\mathcal{H}}:=I_{G/\mathcal{H}}^{\circ}$. Then one has an exact sequence
$0\rightarrow\mathbb{Z}\xrightarrow{(\varepsilon_{G/H}^{\circ})_{H\in\mathcal{H}}}\bigoplus_{H\in\mathcal{H}}\mathbb{Z}[G/H]\rightarrow J_{G/\mathcal{H}}\rightarrow 0.$
Proposition 3.1.
Let $k$ be a field, $\mathbf{K}=\prod_{i=1}^{r}K_{i}$ a finite étale algebra over $k$. Take a finite Galois extension $L$ of $k$ containing $K_{1},\ldots,K_{r}$. Put $G:=\operatorname{Gal}(L/k)$ and $\mathcal{H}:=\{\operatorname{Gal}(L/K_{i})\mid i\in\{1,\ldots,r\}\}$. Then there are isomorphisms of $G$-modules
(3.1) $X_{*}(T_{\mathbf{K}/k})\cong I_{G/\mathcal{H}},\quad X^{*}(T_{\mathbf{K}/k})\cong J_{G/\mathcal{H}}.$
In particular, $T_{\mathbf{K}/k}$ is stably (resp. retract) rational over $k$ if and only if the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation (resp. quasi-invertible).
Proof.
The isomorphisms (3.1) follow from the construction of $I_{G/\mathcal{H}}$ and $J_{G/\mathcal{H}}$. The equivalence between $T_{\mathbf{K}/k}$ to be stably rational (resp. retract rational) and $J_{G/\mathcal{H}}$ to be quasi-permutation (resp. quasi-invertible) is a consequence of Proposition 2.11. $\blacksquare$
In this paper, we determine whether the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation or quasi-invertible in order to classify the stably/retract rationality of multinorm one tori. In the sequel of this section, we discuss
  • how to reduce the problem to a smaller $G$-lattice (Corollary 3.13, Proposition 3.16, Proposition 3.17, Corollary 3.19);
  • behavior of $I_{G/\mathcal{H}}$ and $J_{G/\mathcal{H}}$ with respect to the restriction to subgroups of $G$ (Proposition 3.20); and
  • description of $I_{G/\mathcal{H}}^{N}$ and $J_{G/\mathcal{H}}^{[N]}$ for a normal subgroup $N$ of $G$ (Proposition 3.22).
Accordingly, we extend the notions of character groups and cocharacter groups of multinorm one tori, and set up a framework for dealing with them.
3.2. $G$-lattices $I_{G/\mathcal{H}}^{(\varphi)}$ and $J_{G/\mathcal{H}}^{(\varphi)}$
For a multiset $\mathcal{H}$, we use the notation as follows.
  • Denote by $\mathcal{H}^{\operatorname{set}}$ the underlying set of $\mathcal{H}$.
  • For $H\in\mathcal{H}^{\operatorname{set}}$, write $m_{\mathcal{H}}(H)$ for the multiplicity of $H$ in $\mathcal{H}$. Moreover, we set $m_{\mathcal{H}}(H):=0$ if a subgroup $H$ of $G$ does not belong to $\mathcal{H}^{\operatorname{set}}$.
We define $\Delta$ as follows:
$\Delta:=\coprod_{m\in\mathbb{Z}_{>0}}\Delta_{m}.$
Here, $\Delta_{m}:=\{(d_{1},\ldots,d_{m})\in(\mathbb{Z}_{>0})^{m}\mid d_{1}\leq\cdots\leq d_{m}\}$ for each $m\in\mathbb{Z}_{>0}$. For $\mathbf{d}\in\Delta_{m}$ and $i\in\{1,\ldots,m\}$, we denote by $\mathbf{d}_{i}$ the $i$-th factor of $\mathbf{d}$.
Definition 3.2.
Let $\mathcal{H}$ be a multiset. A weight function on $\mathcal{H}$ is defined as a map
$\varphi\colon\mathcal{H}^{\operatorname{set}}\rightarrow\Delta$
such that $\varphi(H)\in\Delta_{m_{\mathcal{H}}(H)}$ for any $H\in\mathcal{H}^{\operatorname{set}}$.
Definition 3.3.
Let $G$ be a finite group, $\mathcal{H}$ a multiset of its subgroups, and $\varphi$ a weight function on $\mathcal{H}$.
  1. (i)
    We define a weight function $d_{\varphi}$ of $\mathcal{H}^{\operatorname{set}}$ as follows:
    $d_{\varphi}\colon\mathcal{H}^{\operatorname{set}}\rightarrow\Delta_{1};\,H\mapsto\gcd(\varphi(H)_{1},\ldots,\varphi(H)_{m_{\mathcal{H}}(H)}).$
  2. (ii)
    We say that $\varphi$ is normalized if $\gcd(d_{\varphi}(H)\mid H\in\mathcal{H}^{\operatorname{set}})=1$.
The following can be confirmed from the definition:
Lemma 3.4.
Let $G$ be a finite group, $\mathcal{H}$ a multiset of its subgroups, and $\varphi$ a weight function on $\mathcal{H}$. Put
$\varphi^{\operatorname{nor}}\colon\mathcal{H}^{\operatorname{set}}\rightarrow\Delta;\,H\mapsto(d^{-1}\varphi(H)_{1},\ldots,d^{-1}\varphi(H)_{m_{\mathcal{H}}(H)}),$
where $d:=\gcd(d_{\varphi}(H)\mid H\in\mathcal{H}^{\operatorname{set}})$. Then $\varphi^{\operatorname{nor}}$ is a normalized weight function on $\mathcal{H}$.
Definition 3.5.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of its subgroups. Consider a weight function $\varphi$ of $\mathcal{H}$. We define a $G$-lattice $I_{G/\mathcal{H}}^{(\varphi)}$ by the exact sequence
(3.2) $0\rightarrow I_{G/\mathcal{H}}^{(\varphi)}\rightarrow\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\xrightarrow{(\varphi(H)_{1}\cdot\varepsilon_{G/H},\,\ldots,\,\varphi(H)_{m_{\mathcal{H}}(H)}\cdot\varepsilon_{G/H})_{H\in\mathcal{H}^{\operatorname{set}}}}\mathbb{Z}.$
Furthermore, set $J_{G/\mathcal{H}}^{(\varphi)}:=(I_{G/\mathcal{H}}^{(\varphi)})^{\circ}$.
If $\varphi$ is normalized, then the rightmost homomorphism of (3.2) is surjective. Moreover, we have an exact sequence of $G$-lattices
$0\rightarrow\mathbb{Z}\xrightarrow{\left(\varphi(H)\varepsilon_{G/H}^{\circ}\right)_{H\in\mathcal{H}^{\operatorname{set}}}}\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\rightarrow J_{G/\mathcal{H}}^{(\varphi)}\rightarrow 0.$
Remark 3.6.
  1. (i)
    If $\varphi(H)=\underbrace{(1,\ldots,1)}_{m_{\mathcal{H}}(H)}$ for all $H\in\mathcal{H}^{\operatorname{set}}$, then the $G$-lattices $I_{G/\mathcal{H}}^{(\varphi)}$ and $J_{G/\mathcal{H}}^{(\varphi)}$ coincide with $I_{G/\mathcal{H}}$ and $J_{G/\mathcal{H}}$ respectively.
  2. (ii)
    Assume that $G$ is an elementary $p$-abelian group, where $p$ is a prime number, and $\mathcal{H}$ consists of subgroups of index $p$ and the whole group $G$. Let
    $\varphi\colon\mathcal{H}^{\operatorname{set}}\rightarrow\Delta;H\mapsto\begin{cases}\underbrace{(1,\ldots,1)}_{m_{\mathcal{H}}(H)}&\text{if }H\neq G;\\ \underbrace{(p,\ldots,p)}_{m_{\mathcal{H}}(G)}&\text{if }H=G.\end{cases}$
    Then the $G$-lattice $I_{G/\mathcal{H}}^{(\varphi)}$ is the same as $L=\operatorname{Ker}\Phi$ in [End01, §2, l9–14].
Lemma 3.7.
Let $G$ be a finite group, $\mathcal{H}$ a multiset of its subgroups, and $\varphi$ a weight function on $\mathcal{H}$. Then one has $I_{G/\mathcal{H}}^{(\varphi^{\operatorname{nor}})}=I_{G/\mathcal{H}}^{(\varphi)}$ and $J_{G/\mathcal{H}}^{(\varphi^{\operatorname{nor}})}=J_{G/\mathcal{H}}^{(\varphi)}$.
Proof.
It suffices to prove $I_{G/\mathcal{H}}^{(\varphi)}=I_{G/\mathcal{H}}^{(\varphi^{\operatorname{nor}})}$. This follows from the fact that the multiplication by a non-zero integer on $\mathbb{Z}$ is injective. $\blacksquare$
3.3. Reduction to smaller $G$-lattices
We give two types of reduction to smaller $G$-lattices. To accomplish it, we first prepare some lemmas.
Lemma 3.8.
Let $A$ be a (non-necessarily commutative) ring with unit. Consider a commutative diagram of left $A$-modules
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{0}}$${M_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{1}}$${M_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N,}$
where the horizontal sequence is exact and the images of $h_{0}$ and $h_{1}$ coincide. Assume that there is a left splitting $f^{\prime}\colon M_{1}\rightarrow M_{0}$ of the exact sequence satisfying $h_{0}\circ f^{\prime}=h_{1}$. Then there is a split exact sequence
$0\rightarrow\operatorname{Ker}(h_{0})\rightarrow\operatorname{Ker}(h_{1})\rightarrow M_{2}\rightarrow 0.$
Proof.
This follows from the snake lemma. $\blacksquare$
Lemma 3.9.
Let $G$ be a finite group, and $H$ its subgroup. Consider a homomorphism of $G$-lattices
$f:=(c_{i}\varepsilon_{G/H})_{i}\colon\mathbb{Z}[G/H]^{\oplus m}\rightarrow\mathbb{Z},$
where $m\in\mathbb{Z}_{>0}$ and $c_{1},\ldots,c_{m}$ are integers with great common divisor $d$. Then there is an automorphism $\lambda$ of $G$-lattices $\mathbb{Z}[G/H]^{\oplus m}$ such that $f\circ\lambda$ coincides with the composite
$\mathbb{Z}[G/H]^{\oplus m}\xrightarrow{\operatorname{pr}_{1}}\mathbb{Z}[G/H]\xrightarrow{d\varepsilon_{G/H}}\mathbb{Z}.$
Proof.
By the Frobenius reciprocity, there is an isomorphism
$\operatorname{Hom}_{\mathbb{Z}[G]}(\mathbb{Z}[G/H]^{\oplus m},\mathbb{Z})\cong\operatorname{Hom}_{\mathbb{Z}[H]}(\mathbb{Z}^{\oplus m},\mathbb{Z})=\operatorname{Hom}_{\mathbb{Z}}(\mathbb{Z}^{\oplus m},\mathbb{Z}).$
Hence the assertion follows from the theory of invariant factors for finitely generated abelian groups. $\blacksquare$
The following is the first type of reduction to smaller $G$-lattices.
Lemma 3.10.
Let $G$ be a finite group, $\mathcal{H}$ a multiset of its subgroups, and $\varphi$ a weight function on $\mathcal{H}$. Assume that there exist $H_{0},H^{\prime}_{0}\in\mathcal{H}$ and $i_{0},i^{\prime}_{0}\in\{1,\ldots,m_{\mathcal{H}}(H)\}$ such that
  • $H_{0}\subset H^{\prime}_{0}$; and
  • ${\varphi}(H_{0})_{i_{0}}\in{\varphi}(H^{\prime}_{0})_{i^{\prime}_{0}}\mathbb{Z}$.
We denote by $\mathcal{H}^{\prime}$ the multiset of subgroups of $G$ that satisfies the following for every subgroup $H$ of $G$:
$m_{\mathcal{H}^{\prime}}(H^{\prime})=\begin{cases}m_{\mathcal{H}}(H^{\prime})&\text{if }H^{\prime}\in\mathcal{H}^{\operatorname{set}}\setminus\{H_{0}\};\\ m_{\mathcal{H}}(H_{0})-1&\text{if }H^{\prime}=H_{0}.\end{cases}$
Furthermore, we define a weight function $\varphi^{\prime}$ of $\mathcal{H}^{\prime}$ as
$\varphi^{\prime}(H^{\prime})=\begin{cases}\varphi(H^{\prime})&\text{if }H^{\prime}\in\mathcal{H}^{\operatorname{set}}\setminus\{H_{0}\};\\ (\varphi(H_{0})_{i})_{i\in\{1,\ldots,m_{\mathcal{H}}(H)\}\setminus\{i_{0}\}}&\text{if }H^{\prime}=H_{0}.\end{cases}$
Then there exist isomorphisms of $G$-lattices
$I_{G/\mathcal{H}}^{(\varphi)}\cong I_{G/\mathcal{H}^{\prime}}^{({\varphi^{\prime}})}\oplus\mathbb{Z}[G/H_{0}],\quad J_{G/\mathcal{H}}^{(\varphi)}\cong J_{G/\mathcal{H}^{\prime}}^{({\varphi^{\prime}})}\oplus\mathbb{Z}[G/H_{0}].$
Remark 3.11.
If $\varphi(H)=(\underbrace{1,\ldots,1)}_{m_{\mathcal{H}}(H)}$ for all $H\in\mathcal{H}^{\operatorname{set}}$, then Lemma 3.10 is essentially the same as [End11, Proposition 1.3].
Proof.
It suffices to give an isomorphism
(3.3) $I_{G/\mathcal{H}}^{(\varphi)}\cong I_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}\oplus\mathbb{Z}[G/H],$
which easily implies an isomorphism $J_{G/\mathcal{H}}^{(\varphi)}\cong J_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}\oplus\mathbb{Z}[G/H]$. Fix $H^{\prime}_{0}\in\mathcal{H}^{\operatorname{set}}$ containing $H_{0}$, $i_{0}\in\{1,\ldots,m_{\mathcal{H}}(H_{0})\}$ and $i^{\prime}_{0}\in\{1,\ldots,m_{\mathcal{H}}(H^{\prime}_{0})\}$ such that $\varphi(H_{0})_{i_{0}}\in\varphi(H^{\prime}_{0})_{i^{\prime}_{0}}\mathbb{Z}$. Then one has a commutative diagram
[…]
where the horizontal sequence is the canonical split exact sequence. Then the images of the vertical homomorphisms coincide. Now we define a homomorphism $\Phi$ as the direct sum of the identity maps on $\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}$ for all $H\in\mathcal{H}^{\operatorname{set}}\setminus\{H_{0}\}$ and the map
$\mathbb{Z}[G/H_{0}]^{\oplus m_{\mathcal{H}}(H_{0})}\oplus\mathbb{Z}[G/H^{\prime}_{0}]\rightarrow\mathbb{Z}[G/H_{0}]^{\oplus m_{\mathcal{H}}(H_{0})-1}\oplus\mathbb{Z}[G/H^{\prime}_{0}]$
defined by
$((x_{i})_{i},y)\mapsto\left((x_{i})_{i\neq i_{0}},\frac{{\varphi}(H_{0})_{i_{0}}}{{\varphi}(H^{\prime}_{0})_{i_{0}}}\operatorname{Ind}_{H^{\prime}_{0}}^{G}(\varepsilon_{H^{\prime}_{0}/H_{0}})(x)+y\right).$
Then the map $\Phi$ gives a left splitting of the exact sequence in (3.4). Furthermore, by definition, the diagram
${\bigoplus_{H\in\mathcal{H}}\mathbb{Z}[G/H]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi\hskip 5.0pt}$$\scriptstyle{({\varphi}(H)\varepsilon_{G/H})_{H}}$${\bigoplus_{H^{\prime}\in\mathcal{H}^{\prime}}\mathbb{Z}[G/H^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{({\varphi^{\prime}}(H^{\prime})\varepsilon_{G/H^{\prime}})_{H^{\prime}}}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\mathbb{Z}}$
is commutative. Therefore Lemma 3.8 implies the existence of (3.3). This completes the proof of Lemma 3.10. $\blacksquare$
Definition 3.12.
Let $G$ be a finite group, and $\mathcal{H}$ a set of its subgroups (that is, all elements in $\mathcal{H}$ have multiplicity $1$). We say that $\mathcal{H}$ is reduced if $H\not\subset H^{\prime}$ for any $H,H^{\prime}\in\mathcal{H}$ with $H\neq H^{\prime}$ as subgroups of $G$.
For a multiset $\mathcal{H}$ of subgroups of a finite group $G$, we denote by $\mathcal{H}^{\operatorname{red}}$ the subset of $\mathcal{H}^{\operatorname{set}}$ consisting of all elements of $\mathcal{H}^{\operatorname{set}}$ that are maximal with respect to inclusion. Note that it is reduced in the sense of Definition 3.12.
Corollary 3.13.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of its subgroups.
  1. (i)
    Let $\varphi$ be a normalized weight function on $\mathcal{H}$. Then there exist isomorphisms of $G$-lattices
    $I_{G/\mathcal{H}}^{(\varphi)}$ $\cong I_{G/\mathcal{H}^{\operatorname{set}}}^{(d_{\varphi})}\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)-1}\right);$
    $J_{G/\mathcal{H}}^{(\varphi)}$ $\cong J_{G/\mathcal{H}^{\operatorname{set}}}^{(d_{\varphi})}\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)-1}\right).$
    In particular, the $G$-lattice $J_{G/\mathcal{H}}^{(\varphi)}$ is quasi-permutation (resp. quasi-invertible) if and only if $J_{G/\mathcal{H}^{\operatorname{set}}}^{(d_{\varphi})}$ is so.
  2. (ii)
    There exist isomorphisms of $G$-lattices
    $I_{G/\mathcal{H}}$ $\cong I_{G/\mathcal{H}^{\operatorname{red}}}\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{red}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)-1}\right)\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}\setminus\mathcal{H}^{\operatorname{red}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\right);$
    $J_{G/\mathcal{H}}$ $\cong J_{G/\mathcal{H}^{\operatorname{red}}}\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{red}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)-1}\right)\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}\setminus\mathcal{H}^{\operatorname{red}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\right).$
    In particular, the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation (resp. quasi-invertible) if and only if $J_{G/\mathcal{H}^{\operatorname{red}}}$ is so.
Proof.
(i): It suffices to construct an isomorphism on $I_{G/\mathcal{H}}^{(\varphi)}$. By Lemma 3.9, there is an isomorphism
$I_{G/\mathcal{H}}^{(\varphi)}\cong I_{G/\mathcal{H}}^{(\widetilde{d}_{\varphi})},$
where $\widetilde{d}_{\varphi}$ is defined as
$\widetilde{d}_{\varphi}(H):=(0,\ldots,0,d_{\varphi}(H))\in\Delta_{m_{\mathcal{H}}(H)}$
for every $H\in\mathcal{H}^{\operatorname{set}}$. In this case, we have $\widetilde{d}_{\varphi}(H)_{i}\in\widetilde{d}_{\varphi}(H)_{m_{\mathcal{H}}(H)}\mathbb{Z}$ for any $i\in\{1,\ldots,m_{\mathcal{H}}(\mathcal{H})-1\}$. Then the assertion follows from Lemma 3.10.
(ii): By (i), we may assume that $\mathcal{H}$ is a set. Let $\varphi$ be the weight function on $\mathcal{H}$ which takes the value $1$. Then, we have $I_{G/\mathcal{H}}^{(\varphi)}=I_{G/\mathcal{H}}$. Moreover, for every $H\in\mathcal{H}^{\operatorname{set}}\setminus\mathcal{H}^{\operatorname{red}}$, there is $H^{\prime}\in\mathcal{H}^{\operatorname{red}}$ such that $H\subset H^{\prime}$ and $\varphi(H)\in\varphi(H^{\prime})\mathbb{Z}$. Hence the assertion is a consequence of Lemma 3.10. $\blacksquare$
Proposition 3.14.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of its subgroups. Take $H_{0}\in\mathcal{H}$ and $g\in G$. Consider a multiset $\mathcal{H}^{\prime}$ of subgroups of $G$ which satisfies the following for any subgroup $H$ of $G$:
$m_{\mathcal{H}^{\prime}}(H)=\begin{cases}m_{\mathcal{H}}(gH_{0}g^{-1})+1&\text{if }H=gH_{0}g^{-1};\\ m_{\mathcal{H}}(H_{0})-1&\text{if }H=H_{0};\\ m_{\mathcal{H}}(H)&\text{otherwise}.\end{cases}$
Then there exist isomorphisms of $G$-lattices
$I_{G/\mathcal{H}}\cong I_{G/\mathcal{H}^{\prime}},\quad J_{G/\mathcal{H}}\cong J_{G/\mathcal{H}^{\prime}}.$
Proof.
It suffices to prove the left isomorphism. Consider a homomorphism of $G$-lattices
$\mathbb{Z}[G/H_{0}]\rightarrow\mathbb{Z}[G/gH_{0}g^{-1}];\,\sum_{g^{\prime}\in G/H_{0}}a_{g^{\prime}}g^{\prime}\mapsto\sum_{g^{\prime}\in G/H_{0}}a_{g^{\prime}}g^{\prime}g^{-1},$
which is an isomorphism. Then the direct summand of this map and the identity map on $\mathbb{Z}[G/H]$ for all $H\in\mathcal{H}^{\operatorname{set}}$ induce a commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${I_{G/\mathcal{H}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}}\mathbb{Z}[G/H]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\hskip 35.0pt(\varepsilon_{G/H})_{H}}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${I_{G/\mathcal{H}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{H^{\prime}\in\mathcal{H}^{\prime}}\mathbb{Z}[G/H^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hskip 35.0pt(\varepsilon_{G/H^{\prime}})_{H^{\prime}}}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Hence the assertion follows from snake lemma. $\blacksquare$
Definition 3.15.
Let $G$ be a finite group. We say that a set of subgroups $\mathcal{H}$ of $G$ is strongly reduced if $H\not\subset gH^{\prime}g^{-1}$ for any $H,H^{\prime}\in\mathcal{H}$ with $H\neq H^{\prime}$ as subgroups of $G$ and any $g\in G$.
For a multiset $\mathcal{H}$ of subgroups of a finite group $G$, we denote by $\mathcal{H}^{\operatorname{srd}}$ a subset of $\mathcal{H}^{\operatorname{set}}$ that is strongly reduced and maximal with respect to inclusion. Note that the subset $\mathcal{H}^{\operatorname{srd}}$ is not uniquely determined. However, we have $\mathcal{H}^{\operatorname{srd}}\subset\mathcal{H}^{\operatorname{red}}$ by definition.
Proposition 3.16.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of its subgroups. Then there exist isomorphisms of $G$-lattices
$I_{G/\mathcal{H}}\cong I_{G/\mathcal{H}^{\operatorname{srd}}}\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{srd}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)-1}\right)\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}\setminus\mathcal{H}^{\operatorname{srd}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\right),$
$J_{G/\mathcal{H}}\cong J_{G/\mathcal{H}^{\operatorname{srd}}}\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{srd}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)-1}\right)\oplus\left(\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}\setminus\mathcal{H}^{\operatorname{srd}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\right).$
Proof.
Let $\mathcal{H}^{\prime}$ be the multiset of subgroups of $G$ that satisfies $(\mathcal{H}^{\prime})^{\operatorname{set}}=\mathcal{H}^{\operatorname{srd}}$ and
$m_{\mathcal{H}^{\prime}}(H)=\sum_{g\in G/N_{G}(H)}m_{\mathcal{H}}(gHg^{-1})$
for every $H\in\mathcal{H}^{\operatorname{srd}}$. Then, Proposition 3.14 implies that there exist isomorphisms of $G$-lattices
$I_{G/\mathcal{H}}\cong I_{G/\mathcal{H}^{\prime}},\quad J_{G/\mathcal{H}^{\prime}}\cong J_{G/\mathcal{H}}.$
Hence the assertion follows from Corollary 3.13 (ii). $\blacksquare$
The following is the second type of reduction to smaller $G$-lattices.
Proposition 3.17.
Let $G$ be a finite group, $\mathcal{H}$ a multiset of its subgroups, and $\varphi$ be a weight function on $\mathcal{H}$. Assume that there exist $H_{0},H^{\prime}_{0}\in\mathcal{H}$ and $i_{0}\in\{1,\ldots,m_{\mathcal{H}}(H)\}$ such that
  • $H_{0}^{\prime}\subset H_{0}$; and
  • $\varphi(H_{0})_{i_{0}}\in(H_{0}:H^{\prime}_{0})d_{\varphi}(H^{\prime}_{0})\mathbb{Z}$.
We denote by $\mathcal{H}^{\prime}$ the multiset of subgroups of $G$ that satisfies the following for every subgroup $H$ of $G$:
$m_{\mathcal{H}^{\prime}}(H)=\begin{cases}m_{\mathcal{H}}(H_{0})-1&\text{if }H=H_{0};\\ m_{\mathcal{H}}(H)&\text{if }H\neq H_{0}.\end{cases}$
Furthermore, we define a weight function $\varphi^{\prime}$ of $\mathcal{H}^{\prime}$ as
$\varphi^{\prime}(H):=\begin{cases}(\varphi(H_{0})_{1},\ldots,\varphi(H_{0})_{i-1},\varphi(H_{0})_{i+1},\ldots,\varphi(H_{0})_{i})&\text{if }H=H_{0}\in\mathcal{H}^{\operatorname{set}};\\ \varphi(H)&\text{if }H\neq H_{0},\end{cases}$
which is normalized. Then there exist isomorphisms of $G$-lattices
(3.5) $I_{G/\mathcal{H}}^{(\varphi)}\cong I_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}\oplus\mathbb{Z}[G/H_{0}],\quad J_{G/\mathcal{H}}^{(\varphi)}\cong J_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}\oplus\mathbb{Z}[G/H_{0}].$
In particular, the $G$-lattice $J_{G/\mathcal{H}}^{(\varphi)}$ is quasi-permutation (resp. quasi-invertible) if and only if $J_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}$ is so.
Remark 3.18.
Assume that
  • $G\cong(C_{p})^{\nu}$ for some prime number $p$ and $\nu\in\mathbb{Z}_{>0}$;
  • $\mathcal{H}$ consists of $G$ and some subgroups of index $p$ in $G$; and
  • $\varphi(H)=\begin{cases}(\underbrace{1,\ldots,1}_{m_{\mathcal{H}}(H)})&\text{if }H\neq G;\\ (\underbrace{p,\ldots,p}_{m_{\mathcal{H}}(H)})&\text{if }H=G.\end{cases}$
Then Theorem 3.17 implies [End01, p. 29, Lemma].
Proof.
It suffices to prove the isomorphism on $I_{G/\mathcal{H}}^{(\varphi)}$. Consider the commutative diagram
(3.6) ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{H\in(\mathcal{H}^{\prime})^{\operatorname{set}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}^{\prime}}(H)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\varphi^{\prime}(H)\cdot\varepsilon_{G/H})_{H\in(\mathcal{H}^{\prime})^{\operatorname{set}}}}$${\bigoplus_{H\in\mathcal{H}^{\operatorname{set}}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\varphi(H)\cdot\varepsilon_{G/H})_{H\in\mathcal{H}^{\operatorname{set}}}}$${\mathbb{Z}[G/H_{0}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z},}$
where the horizontal sequence is the canonical split exact sequence. By the definitions of $\mathcal{H}^{\prime}$ and $\varphi^{\prime}$, the images of the vertical maps coincide. Now, we define the map $\Psi$ as the direct sum of the identity maps on $\mathbb{Z}[G/H_{0}]^{\oplus m_{\mathcal{H}}(H_{0})-1}$ and $\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}^{\prime}}(H)}$ for all $H\in(\mathcal{H}^{\prime})^{\operatorname{set}}$, and the map
$\mathbb{Z}[G/H_{0}]^{\oplus m_{\mathcal{H}}(H)}\rightarrow\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}^{\prime}}(H)}$
defined as
$(x_{i})_{i}\mapsto(x_{i})_{i\neq i_{0}}+\frac{\varphi(H_{0})_{i_{0}}}{(H_{0}:H^{\prime}_{0})d_{\varphi}(H_{0})}\operatorname{Ind}_{H_{0}}^{G}\varepsilon_{H_{0}/H^{\prime}_{0}}(x_{i_{0}}).$
Then it gives a left splitting of the exact sequence in (3.6). Moreover, the diagram
${\bigoplus_{H\in\mathcal{H}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}}(H)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi\hskip 10.0pt}$$\scriptstyle{(\varphi(H)\cdot\varepsilon_{G/H})_{H\in\mathcal{H}^{\operatorname{set}}}}$${\bigoplus_{H\in\mathcal{H}^{\prime}}\mathbb{Z}[G/H]^{\oplus m_{\mathcal{H}^{\prime}}(H)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\varphi(H)\cdot\varepsilon_{G/H})_{H\in(H^{\prime})^{\operatorname{set}}}}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\mathbb{Z}}$
is commutative. Hence we obtain the left isomorphism in (3.5) as desired. $\blacksquare$
Corollary 3.19.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of its subgroups. Consider a normalized weight function $\varphi$ on $\mathcal{H}$, and define a set of subgroup of $G$ as follows:
$\mathcal{H}_{\varphi}[1]:=\{H\in\mathcal{H}^{\operatorname{set}}\mid d_{\varphi}(H)=1\}.$
Assume that
  • (a)
    $\mathcal{H}_{\varphi}[1]\neq\emptyset$; and
  • (b)
    for each $H\in\mathcal{H}^{\operatorname{set}}\setminus\mathcal{H}_{\varphi}[1]$, there is $H^{\prime}\in\mathcal{H}_{\varphi}[1]$ such that $d_{\varphi}(H)\in(H:H\cap H^{\prime})\mathbb{Z}$.
Then the following hold:
$I_{G/\mathcal{H}}^{(\varphi)}]=[I_{G/\mathcal{H}_{\varphi}[1]}],\quad[J_{G/\mathcal{H}}^{(\varphi)}]=[J_{G/\mathcal{H}_{\varphi}[1]}].$
In particular, the $G$-lattice $J_{G/\mathcal{H}}^{(\varphi)}$ is quasi-permutation (resp. quasi-invertible) if and only if $J_{G/\mathcal{H}_{\varphi}[1]}$ is so.
Proof.
It suffices to prove the isomorphism on $I_{G/\mathcal{H}}^{(\varphi)}$. By Corollary 3.13 (i), we may assume $\mathcal{H}=\mathcal{H}^{\operatorname{set}}$. In particular, we have $\varphi=d_{\varphi}$. Write $\mathcal{H}\setminus\mathcal{H}_{\varphi}[1]=\{H_{1},\ldots,H_{s}\}$. For each $i\in\{1,\ldots,s\}$, take $H_{i}^{\dagger}\in\mathcal{H}_{\varphi}[1]$ so that
(3.7) $d_{\varphi}(H_{i})\in(H_{i}:H_{i}\cap H_{i}^{\dagger})\mathbb{Z}.$
Note that this is possible by (b). Consider a set $\mathcal{H}^{\dagger}:=\{H_{i}\cap H_{i}^{\dagger}\mid i\in\{1,\ldots,s\}\}$, and we define a multiset $\widetilde{\mathcal{H}}$ of subgroups of $G$ as the disjoint union of $\mathcal{H}$ and $\mathcal{H}^{\dagger}$. Moreover, let $\widetilde{\varphi}$ be the normalized weight function on $\widetilde{\mathcal{H}}$ defined as
$\widetilde{\varphi}(H^{\prime}):=\begin{cases}(1,\varphi(H^{\prime}))&\text{if }H^{\prime}\in\mathcal{H}\cap\mathcal{H}^{\dagger};\\ 1&\text{if $H^{\prime}\in\mathcal{H}^{\dagger}\setminus\mathcal{H}$;}\\ \varphi(H^{\prime})&\text{if }H^{\prime}\in\mathcal{H}\setminus\mathcal{H}^{\dagger}.\end{cases}$
Since $\widetilde{\varphi}(H^{\prime})_{1}=1$ for any $H^{\prime}\in\mathcal{H}^{\dagger}$, Lemma 3.10 gives an equality
$I_{G/\widetilde{\mathcal{H}}}^{(\widetilde{\varphi})}]=[I_{G/\mathcal{H}}^{(\varphi)}].$
Moreover, since $\widetilde{\mathcal{H}}^{\operatorname{set}}=\mathcal{H}\cup\mathcal{H}^{\dagger}$, Corollary 3.13 (i) implies
$I_{G/\widetilde{\mathcal{H}}}^{(\widetilde{\varphi})}]=[I_{G/(\mathcal{H}\cup\mathcal{H}^{\dagger})}^{(d_{\widetilde{\varphi}})}$].
On the other hand, take $H\in(\mathcal{H}\cup\mathcal{H}^{\dagger})\setminus\mathcal{H}^{\prime}$. Then we have $H\in\mathcal{H}\setminus\mathcal{H}_{\varphi}[1]$, and hence $H=H_{i}$ for some $i$. Moreover, one has
$d_{\widetilde{\varphi}}(H_{i})=\varphi(H_{i})\in(H_{i}:H_{i}\cap H_{i}^{\dagger})\mathbb{Z}=(H_{i}:H_{i}\cap H_{i}^{\dagger})d_{\widetilde{\varphi}}(H_{i}\cap H_{i}^{\dagger})\mathbb{Z}$
by (3.7). Therefore, we can apply Proposition 3.17 to $\mathcal{H}\cup\mathcal{H}^{\dagger}$, $d_{\widetilde{\varphi}}$ and the inclusion $H_{i}\cap H_{i}^{\dagger}\subset H_{i}=H$. Consequently, we obtain an equality
$I_{G/(\mathcal{H}\cup\mathcal{H}^{\dagger})}^{(d_{\widetilde{\varphi}})}]=[I_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}].$
Here, $\mathcal{H}^{\prime}:=\mathcal{H}_{\varphi}[1]\cup\mathcal{H}^{\dagger}$, and $\varphi^{\prime}$ is the restriction to $\mathcal{H}^{\prime}$ of $d_{\widetilde{\varphi}}$. Then we have $\varphi^{\prime}(H^{\prime})=1$ for any $H^{\prime}\in\mathcal{H}^{\prime}$, and hence we obtain an equality
$I_{G/\mathcal{H}^{\prime}}^{(\varphi^{\prime})}=I_{G/\mathcal{H}^{\prime}}.$
Now, recall that any element $H^{\prime}$ of $\mathcal{H}^{\dagger}\setminus\mathcal{H}_{\varphi}[1]$ satisfies $H^{\prime}=H_{i}\cap H_{i}^{\dagger}$ for some $i$. Since $H_{i}^{\dagger}$ is an element of $\mathcal{H}_{\varphi}[1]$, we can apply Lemma 3.10 to $\mathcal{H}^{\prime}$, $\varphi^{\prime}$ and the inclusion $H^{\prime}\subset H_{i}^{\dagger}$. Repeating this argument for all $H^{\prime}\in\mathcal{H}^{\dagger}\setminus\mathcal{H}_{\varphi}[1]$, we get
$I_{G/\mathcal{H}^{\prime}}]=[I_{G/\mathcal{H}_{\varphi}[1]}].$
Consequently, we obtain the desired assertion. $\blacksquare$
3.4. Reduction to lattices over smaller groups
We first describe $I_{G/\mathcal{H}}^{(\varphi)}$ and $J_{G/\mathcal{H}}^{(\varphi)}$ as $P$-lattices, where $P$ is a subgroup of $G$.
Proposition 3.20.
Let $G$ be a finite group, and $P$ a subgroup of $G$. Consider a multiset of subgroups $\mathcal{H}$ of $G$ and a normalized weight function $\varphi$ on $\mathcal{H}$. Then there are isomorphisms of $P$-modules
$I_{G/\mathcal{H}}^{(\varphi)}\cong I_{P/\mathcal{H}_{P}}^{(\varphi_{P})},\quad J_{G/\mathcal{H}}^{(\varphi)}\cong J_{P/\mathcal{H}_{P}}^{(\varphi_{P})}.$
Here $\varphi_{P}$ is defined as follows:
[…]
  • $\mathcal{H}_{P}$ the multiset of subgroups of $G$ consisting $P\cap gHg^{-1}$ for all $H\in\mathcal{H}$ and $g\in C_{H}$; and
  • $\varphi_{P}$ is the normalized weight function on $\mathcal{H}_{P}$ which sends $H^{\prime}\in\mathcal{H}_{P}$ to the element of $\Delta$ defined by $\varphi(H)$ for all $H\in\mathcal{H}$ with $H^{\prime}=P\cap gHg^{-1}$ for some $g\in C_{H}$.
In particular, if $J_{G/\mathcal{H}}^{(\varphi)}$ is quasi-permutation (resp. quasi-invertible), then $J_{P/\mathcal{H}_{P}}^{(\varphi_{P})}$ is so.
Proof.
This is a consequence of Mackey’s decomposition. See [Ser77, Section 7.3, Proposition 22]. $\blacksquare$
Next, we describe $(I_{G/\mathcal{H}}^{(\varphi)})^{N}$ and $(J_{G/\mathcal{H}}^{(\varphi)})^{[N]}$ for a normal subgroup $N$ of $G$.
Lemma 3.21.
Let $G$ be a finite group, and $H$ its subgroup.
  1. (i)
    For any subgroup $H^{\prime}$ of $G$ containing $H$, the diagram
    ${\mathbb{Z}[G/H^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon_{G/H^{\prime}}}$${\mathbb{Z}[G/H]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon_{G/H}}$$\scriptstyle{\operatorname{Ind}_{H^{\prime}}^{G}(\varepsilon_{H^{\prime}/H})}$${\mathbb{Z}[G/H^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon_{G/H^{\prime}}}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(H^{\prime}:H)}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\mathbb{Z}}$
    is commutative.
  2. (ii)
    Let $N$ be a normal subgroup of $G$. Then the image of the canonical injection
    $\mathbb{Z}[G/HN]\hookrightarrow\mathbb{Z}[G/H]$
    coincides with $\mathbb{Z}[G/H]^{N}$.
Proof.
This follows from the definitions of the augmentation maps and their induced maps. $\blacksquare$
Proposition 3.22.
Let $G$ be a finite group, and $N$ its normal subgroup. Take a multiset of subgroups $\mathcal{H}$ of $G$ and a normalized weight function $\varphi$ of $\mathcal{H}$. Then there are isomorphisms
$(I_{G/\mathcal{H}}^{(\varphi)})^{N}\cong I_{(G/N)/\mathcal{H}^{N}}^{(\overline{\varphi}_{G/N}^{\operatorname{nor}})},\quad(J_{G/\mathcal{H}}^{(\varphi)})^{[N]}\cong J_{(G/N)/\mathcal{H}^{N}}^{(\overline{\varphi}_{G/N}^{\operatorname{nor}})}.$
Here $\mathcal{H}^{N}$ and $\overline{\varphi}_{G/N}^{\operatorname{nor}}$ are defined as follows:
  • $\mathcal{H}^{N}$ is the multiset of subgroups of $G/N$ consisting of $HN/N$ for all $H\in\mathcal{H}$ (in particular, the multiplicity of $\overline{H}$ in $\mathcal{H}^{N}$ is the sum of $m_{\mathcal{H}}(H)$ for all $H\in\mathcal{H}$ with $HN/N=\overline{H}$); and
  • the weight function $\overline{\varphi}_{G/N}$ on $\mathcal{H}^{N}$ maps $\overline{H}\in\mathcal{H}^{N}$ to the element of $\Delta$ defined by $(HN:H)\varphi(H)$ for all $H\in\mathcal{H}$ with $HN/N=\overline{H}$; and
  • $\overline{\varphi}_{G/N}^{\operatorname{nor}}$ is as in Lemma 3.4.
In particular, if $J_{G/\mathcal{H}}^{(\varphi)}$ is quasi-permutation (resp. quasi-invertible), then $J_{G/\mathcal{H}^{N}}^{(\overline{\varphi}_{G/N}^{\operatorname{nor}})}$ is so.
Proof.
By Lemma 3.21, we obtain an isomorphism of $G/N$-lattice $(I_{G/\mathcal{H}}^{(\varphi)})^{N}\cong I_{(G/N)/\mathcal{H}^{N}}^{(\overline{\varphi}_{G/N})}$. On the other hand, we have $I_{G/N}^{(\overline{\varphi}_{G/N})}=I_{G/N}^{(\overline{\varphi}_{G/N}^{\operatorname{nor}})}$ by Lemma 3.7. Hence we obtain $(I_{G/\mathcal{H}}^{(\varphi)})^{N}\cong I_{G/N}^{(\overline{\varphi}_{G/N}^{\operatorname{nor}})}$ as desired. The isomorphism on $(J_{G/\mathcal{H}}^{(\varphi)})^{[N]}$ is a consequence of that on $(I_{G/\mathcal{H}}^{(\varphi)})^{N}$. $\blacksquare$
For a multiset $\mathcal{H}$ of subgroups of $G$, we denote by $N^{G}(\mathcal{H})$ the maximum normal subgroup of $G$ which is contained in $H$ for all $H\in\mathcal{H}$. Moreover, we simply denote $N^{G}(\mathcal{H})$ by $N^{G}(H)$ if $\mathcal{H}$ consists of a single subgroup $H$.
Corollary 3.23.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of subgroups of $G$. Consider a normal subgroup $N$ of $G$ that is contained in $N^{G}(\mathcal{H})$. Then there exist isomorphisms of $G$-lattices
$I_{G/\mathcal{H}}\cong I_{(G/N)/\mathcal{H}^{N}},\quad J_{G/\mathcal{H}}\cong J_{(G/N)/\mathcal{H}^{N}}.$
Here we regard $I_{(G/N)/\mathcal{H}^{N}}$ and $J_{(G/N)/\mathcal{H}^{N}}$ as $G$-lattices by the natural surjection $G\twoheadrightarrow G/N$.
Proof.
This follows from Proposition 3.22 since the actions of $N$ on $I_{G/\mathcal{H}}$ and $J_{G/\mathcal{H}}$ are trivial. $\blacksquare$
4. $p$-groups
For a finite group $G$, we write for $\Phi(G)$ the Frattini subgroup of $G$, that is, the intersection of all maximal subgroups of $G$. Here maximal subgroups mean proper subgroups which are maximal with respect to inclusion.
Proposition 4.1 ([Hal59, Theorem 4.3.2]).
Let $G$ be a $p$-group, where $p$ is a prime number. Then all maximal subgroups of $G$ are normal of index $p$. In particular, the Frattini subgroup $\Phi(G)$ contains the derived subgroup of $G$, and $G/\Phi(G)$ is an elementary $p$-abelian group.
Corollary 4.2.
Let $p$ be a prime number, $G$ a $p$-group, and $H$ its subgroup.
  1. (i)
    The subgroups $\Phi(G)$ and $H$ do not generate $G$.
  2. (ii)
    We further assume $(N_{G}(H):H)=p$. If a subgroup $P$ of $G$ contains $H$ properly, then we have $N_{G}(H)\subset P$.
Proof.
(i): Let $P$ be a maximal subgroup of $G$ containing $H$. Then it contains $\Phi(G)$ by definition, and hence $\Phi(G)H\subset P\subsetneq G$ as desired.
(ii): By Proposition 4.1, there exists a subgroup $P^{\prime}$ of $P$ of order $p\cdot\#H$ that contains $H$. On the other hand, the assumption $(N_{G}(H):H)=p$ implies that $N_{G}(H)$ is the unique subgroup of $G$ of order $p\cdot\#H$ that contains $H$. Hence we obtain $P^{\prime}=N_{G}(H)$, which concludes the desired assertion. $\blacksquare$
We use a basic result of $p$-groups as follows:
Lemma 4.3.
Let $p$ be a prime number, $G$ a $p$-group, and $N$ a normal subgroup of $G$. Then we have $Z(G)\cap N\neq\{1\}$.
In the sequel of this paper, for a positive integer $n$, we denote by $D_{n}$ the dihedral group of order $2n$, that is,
$D_{n}=\langle\sigma_{n},\tau_{n}\mid\sigma_{n}^{n}=\tau_{n}^{2}=1,\tau_{n}\sigma_{n}\tau_{n}=\sigma_{n}^{-1}\rangle.$
Let $p$ be a prime number. Recall that a finite group $G$ of order $p^{n}$ is said to be of maximal class if its nilpotency class is n1n-1.
Proposition 4.4 ([Ber08, Corollary 1.7]).
Let $G$ be a $2$-group of maximal class. Then $G$ is isomorphic to one of the following:
  1. (i)
    the dihedral group $D_{2^{\nu}}$ of order $2^{\nu+1}$ for $\nu\geq 2$;
  2. (ii)
    the semi-dihedral group $SD_{2^{\nu+1}}$ of order $2^{\nu+1}$ for $\nu\geq 3$;
  3. (iii)
    the generalized quaternion group $Q_{2^{\nu}}$ of order $2^{\nu}$ for $\nu\geq 3$, that is,
    $Q_{2^{\nu}}=\langle i_{\nu},j_{\nu}\mid i_{\nu}^{2^{\nu-1}}=1,j_{\nu}^{2}=i_{\nu}^{2^{\nu-2}},j_{\nu}i_{\nu}j_{\nu}^{-1}=i_{\nu}^{-1}\rangle.$
Proposition 4.5.
Let $G$ be a $2$-group of order a multiple of $8$ which is not of maximal class.
  1. (i)
    There exists an abelian normal subgroup $E$ of $G$ of order $8$.
  2. (ii)
    Under the notation in (i), we further assume $E\cong C_{4}\times C_{2}$. Then $\Phi(E)$ is contained in the center of $G$.
Proof.
This is explained in [End11, p. 91, Proof of Step 4]. However, we give a proof for reader’s convenience.
(i): By [Ber08, Lemma 1.4], one can take a non-cyclic normal subgroup $E^{\prime}$ of $G$ of order $4$ since $G$ is not of maximal class. Furthermore, [Ber08, Proposition 1.8] implies that the centralizer of $E^{\prime}$ in $G$ does not coincide with $E^{\prime}$. Now, let $E$ be the subgroup of $G$ generated by $E^{\prime}$ and an element whose image in $G/E^{\prime}$ is central of order $2$. Then $E$ is normal in $G$ and has order $8$.
(ii): This follows from the fact that $\Phi(E)$ is a characteristic subgroup of $E$. $\blacksquare$
The following two lemmas can be obtained by direct computation.
Lemma 4.6.
Let $\nu\geq 2$ be an integer.
  1. (i)
    For any $m\in\mathbb{Z}$, we have $N_{D_{2^{\nu}}}(\langle\sigma_{2^{\nu}}^{m}\tau_{2^{\nu}}\rangle)=\langle\sigma_{2^{\nu}}^{2^{\nu-1}},\sigma_{2^{\nu}}^{m}\tau_{2^{\nu}}\rangle$.
  2. (ii)
    For two integers $m$ and $m^{\prime}$, $\langle\sigma_{2^{\nu}}^{m}\tau_{2^{\nu}}\rangle$ and $\langle\sigma_{2^{\nu}}^{m^{\prime}}\tau_{2^{\nu}}\rangle$ are conjugate in $D_{2^{\nu}}$ if and only if $m-m^{\prime}$ is even.
  3. (iii)
    Every non-normal subgroup of $D_{2^{\nu}}$ is of the form $\langle\sigma_{2^{\nu}}^{m}\tau_{2^{\nu}}\rangle$ for some integer $m$.
Lemma 4.7.
Let $\nu\geq 3$ be an integer.
  1. (i)
    All non-normal subgroups of order $2$ in $SD_{2^{\nu+1}}$ are conjugate to each other.
  2. (ii)
    There exists a unique subgroup of order $2$ in $Q_{2^{\nu}}$, which is the center of $Q_{2^{\nu}}$.
5. Proof of Theorem 1.2
First, we specify previous results given by Endo–Miyata ([EM75]) and Endo ([End01], [End11]).
Proposition 5.1 ([EM75, Theorem 2.3]).
For a finite group $G$, the following are equivalent:
  1. (i)
    the $G$-lattice $J_{G}$ is quasi-permutation;
  2. (ii)
    there is an isomorphism $G\cong C_{m}\rtimes C_{2^{\nu}}$, where $m$ is an odd integer, $\nu\in\mathbb{Z}_{\geq 0}$, and the action of $C_{2^{\nu}}$ on $C_{m}$ factors through a homomorphism $C_{2^{\nu}}\rightarrow C_{2}$.
Proposition 5.2.
Let $p$ be a prime number, $G$ a $p$-group, and $H$ a subgroup of $G$. Then the following are equivalent:
  1. (i)
    $J_{G/H}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/H}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $G/N^{G}(H)$ is cyclic.
Proof.
Case 1. $H$ is normal in $G$. It suffices to prove (ii) $\Rightarrow$ (iii) $\Rightarrow$ (i). We may assume $H=\{1\}$. We first assume that $J_{G/H}$ is quasi-invertible. Then, [EM75, Theorem 1.5] implies the cyclicity of all the Sylow subgroups of $G$. This implies that $G$ is cyclic, because it is a $p$-group. Hence, (iii) is valid. On the other hand, if $G$ is cyclic, then (i) follows from Proposition 5.1.
Case 2. $H$ is not normal in $G$. We may assume $N^{G}(H)=\{1\}$. It suffices to prove that $J_{G/H}$ is not quasi-invertible. However, it is the same as [End11, Theorem 2.1]. $\blacksquare$
Proposition 5.3 ([End01, Theorem 1]).
Let $p$ be a prime number, and $G$ an elementary $p$-abelian group. Take a reduced set $\mathcal{H}$ of subgroups of $G$. If $(G:H)=p$ for all $H\in\mathcal{H}$, then the following are equivalent:
  1. (i)
    $J_{G/\mathcal{H}}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/\mathcal{H}}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $\mathcal{H}=1$ or $p=\#\mathcal{H}=2$.
Definition 5.4.
For a multiset $\mathcal{H}$ of subgroups of a finite group $G$, put
$\mu(\mathcal{H}):=\min\{(G:H)\in\mathbb{Z}_{>0}\mid H\in\mathcal{H}\},\quad M(\mathcal{H}):=\max\{(G:H)\in\mathbb{Z}_{>0}\mid H\in\mathcal{H}\}.$
Lemma 5.5.
Let $G$ be a finite group, and $\mathcal{H}$ a reduced set of its subgroups. If $\mathcal{H}\geq 2$, then we have $(G:N^{G}(\mathcal{H}))>M(\mathcal{H})$.
Proof.
Take $H_{0}\in\mathcal{H}$ with $(G:H_{0})=M(\mathcal{H})$. If $(G:N^{G}(H_{0}))>M(\mathcal{H})$, then the assertion is clear. Otherwise, $H_{0}$ is normal in $G$. Take $H\in\mathcal{H}\setminus\{H_{0}\}$, then $N^{G}(H)$ does not contain $H_{0}$ since $\mathcal{H}$ is reduced. Hence $(G:N^{G}(\{H,H_{0}\}))>M(\mathcal{H})$. This implies the desired assertion since $N^{G}(\mathcal{H})$ is contained in $N^{G}(\{H,H_{0}\})$. $\blacksquare$
Definition 5.6.
For a multiset $\mathcal{H}$ of subgroups of a finite group $G$, set
$\mathcal{H}^{\operatorname{nor}}:=\{H\in\mathcal{H}\mid H\triangleleft G\}.$
Lemma 5.7.
Let $p$ be a prime number, $G$ a $p$-group, and $\mathcal{H}$ a reduced set of its subgroups. If $G/N^{G}(\mathcal{H})$ is cyclic, then $\mathcal{H}=1$ and $\mathcal{H}^{\operatorname{nor}}=\mathcal{H}$.
Proof.
Take $H,H^{\prime}\in\mathcal{H}$. By definition, $H$ and $H^{\prime}$ contain $N^{G}(\mathcal{H})$. Since $G/N^{G}(\mathcal{H})$ is cyclic, we have $H\subset H^{\prime}$ or $H^{\prime}\subset H$. This implies $H=H^{\prime}$ since $\mathcal{H}$ is reduced. Hence, we obtain $\mathcal{H}=1$. The equality $\mathcal{H}^{\operatorname{nor}}=\mathcal{H}$ follows from the fact that all subgroups of $G$ containing its derived subgroup are normal in $G$. $\blacksquare$
Definition 5.8.
Let $G$ be a finite group, and $P$ its subgroup. For a multiset $\mathcal{H}$ of subgroups of $G$, set
$\mathcal{H}_{\subset P}:=\{H\in\mathcal{H}\mid H\subset P\}.$
Lemma 5.9.
Let $p$ be a prime number, $\nu\geq 3$ a positive integer, and $G$ a finite group of order $p^{\nu}$. Consider a reduced set $\mathcal{H}$ of subgroups of $G$ satisfying $\mathcal{H}\geq 2$ and $\mu(\mathcal{H})\geq p^{2}$. Take a maximal subgroup $P$ of $G$. Assume that there is $H_{1}\in\mathcal{H}$ with $(G:H_{1})=\mu(\mathcal{H})$ such that all elements of $\mathcal{H}_{P}^{\operatorname{red}}$ are conjugate to $H_{1}$. If $H_{0}\in\mathcal{H}^{\operatorname{nor}}\setminus\{H_{1}\}$ has index $\mu(\mathcal{H})$ in $G$, then it is contained in $N_{G}(H_{1})$. Moreover, we can take a maximal subgroup $P^{\prime}$ of $G$ so that $\mathcal{H}_{\subset P^{\prime}}$ contains $H_{0}$ and $H_{1}$ (in particular, $\mathcal{H}_{P}^{\operatorname{red}}$ contains $H_{0}$).
Proof.
We first prove $H_{0}\subset N_{G}(H_{1})$. By assumption, $P\cap H_{0}$ is contained in $gH_{1}g^{-1}$ for some $g\in G$. This is equivalent to $P\cap H_{0}\subset H_{1}$ since $P$ and $H_{0}$ are normal in $G$ Hence one has
$H_{0}\cap H_{1}=P\cap H_{0}\cap H_{1}=P\cap H_{1}.$
Combining this equality with the normality of $H_{0}$ in $G$, we obtain
$(H_{0}H_{1}:H_{1})=(H_{0}H_{1}:H_{0})=(H_{1}:H_{0}\cap H_{1})=p.$
On the other hand, $H_{0}H_{1}$ is a $p$-group since $G$ is so. Consequently we have
$H_{0}\subset H_{0}H_{1}\subset N_{G}(H_{1})$
by Proposition 4.1.
Secondly, we construct a maximal subgroup $P^{\prime}$ of $G$ that satisfies $\{H_{0},H_{1}\}\subset\mathcal{H}_{\subset P}$. Since $\mu(\mathcal{H})\geq p^{2}$, one has
$(G:H_{0}H_{1})=p^{-1}(G:H_{1})=p^{-1}\mu(\mathcal{H})\geq p.$
Hence we may take $P^{\prime}$ so that $H_{0}H_{1}$ is contained. $\blacksquare$
Theorem 5.10.
Let $p$ be an odd prime number, and $G$ a $p$-group. Consider a reduced set $\mathcal{H}$ of subgroups of $G$. Then the following are equivalent:
  1. (i)
    $J_{G/\mathcal{H}}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/\mathcal{H}}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $\mathcal{H}=1$ and $G/N^{G}(\mathcal{H})$ is cyclic.
Proof.
(i) $\Rightarrow$ (ii) is clear. (iii) $\Rightarrow$ (i) follows from Proposition 5.2. In the following, we prove (ii) $\Rightarrow$ (iii), which is archived by giving a proof of the contraposition. We may assume that $\mathcal{H}$ contains at least two elements. In particular, $G$ is not cyclic according to Lemma 5.7. It suffices to prove that $J_{G/H}$ is not quasi-invertible if $\mathcal{H}\geq 2$. Write $G=p^{\nu}$. We give a proof of the above assertion by induction on $\nu$. If $\nu=2$, the assertion follows from Proposition 5.3. Now suppose $\nu\geq 3$, and the assertion holds for all $\nu-1$. If $N^{G}(\mathcal{H})\neq\{1\}$, take a subgroup $N$ of order $p$ in $Z(G)\cap N^{G}(\mathcal{H})$. Note that $Z(G)\cap N^{G}(\mathcal{H})$ is non-trivial by Lemma 4.3. Then it suffices to prove the assertion for the $G/N$-lattice $J_{(G/N)/\mathcal{H}^{N}}$. Since $(G/N)=p^{\nu-1}$, the assertion follows from the induction hypothesis. Hence, we may further assume $N^{G}(\mathcal{H})=\{1\}$. If $M(\mathcal{H})=p$, then $G$ is elementary $p$-abelian since $N^{G}(\mathcal{H})=\{1\}$. Therefore, the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible by Proposition 5.3. Therefore, we can impose $M(\mathcal{H})\geq p^{2}$ in the sequel. By the induction hypothesis, it suffices to prove that there is a maximal subgroup $P$ of $G$ that satisfies $\mathcal{H}_{P}^{\operatorname{red}}\geq 2$.
Case 1. $\mathcal{H}=\mathcal{H}^{\operatorname{nor}}$.
Recall the notations in Definition 5.4, that is,
$M(\mathcal{H}):=\max\{(G:H)\in\mathbb{Z}_{>0}\mid(G:H)\in\mathcal{H}\},\quad\mu(\mathcal{H}):=\min\{(G:H)\in\mathbb{Z}_{>0}\mid(G:H)\in\mathcal{H}\}.$
Case 1-a. $\mu(\mathcal{H})<M(\mathcal{H})$. Take $H_{0}\in\mathcal{H}$ with $(G:H_{0})=M(\mathcal{H})$. Pick a maximal subgroup $P$ of $G$ containing $H_{0}$. Then, for any $H\in\mathcal{H}$ with $(G:H)<M(\mathcal{H})$, one has $H_{0}\not\subset P\cap H$ and $H_{0}\not\supset P\cap H$. In particular, $\mathcal{H}_{P}^{\operatorname{red}}$ contains at least two elements. This completes the proof in this case.
Case 1-b. $\mu(\mathcal{H})=M(\mathcal{H})$. In this case, the assertion follows from Lemma 5.9 since $\mathcal{H}\geq 2$ and $\mu(\mathcal{H})\geq p^{2}$.
Case 2. $\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}$ is non-empty.
Note that one has $M(\mathcal{H})\geq p^{2}$ by Proposition 4.1. Put
$\sigma:=\mu(\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}).$
Take $H\in\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}$ with $(G:H)=\sigma$, and pick a maximal subgroup $P$ of $G$ containing $N_{G}(H)$. Fix a complete representative $C$ of $P\backslash G/H\cong G/P$ in $G$. Let $H^{\prime}\in\mathcal{H}$. If $H^{\prime}\in\mathcal{H}^{\operatorname{nor}}$, then $gHg^{-1}\not\subset P\cap H^{\prime}$ since $H^{\prime}$ is normal in $G$. On the other hand, if $H^{\prime}\in\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}$, then the inclusion $gHg^{-1}\not\subset P\cap g^{\prime}H^{\prime}(g^{\prime})^{-1}$ holds for any $g^{\prime}\in G$ since $(G:H)=\sigma$. Therefore, $\mathcal{H}_{P}^{\operatorname{red}}$ contains $gHg^{-1}$ for any $g\in C$. In particular, $\mathcal{H}_{P}^{\operatorname{red}}$ contains at least $p$ elements. Hence the proof is complete. $\blacksquare$
Remark 5.11.
In the proof of Theorem 5.10, Case 2 does not use the assumption $\mathcal{H}\geq 2$. In particular, we also obtain an alternative proof of Proposition 5.2 in the case where $H$ is not normal in $G$.
Proof of Theorem 1.2.
Put $G:=\operatorname{Gal}(L/k)$ and
$\mathcal{H}:=\{\operatorname{Gal}(L/K_{1}),\ldots,\operatorname{Gal}(L/K_{r})\}.$
Then Proposition 3.1 gives an isomorphism $X^{*}(T_{\mathbf{K}/k})\cong J_{G/\mathcal{H}}$. Hence, by Proposition 3.1, it suffices to prove that the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible. In this case, $\mathcal{H}$ is a reduced set. Therefore, the assertion follows from Theorem 5.10. $\blacksquare$
6. Flabby resolutions of particular lattices
6.1. Quasi-permutation lattices
Here we give some examples of $G$-lattices $J_{G/\mathcal{H}}^{(\varphi)}$ that are quasi-permutation. For $n\in\mathbb{Z}_{>0}$, we denote by $D_{n}$ the dihedral group of order $2n$, that is,
$D_{n}:=\langle\sigma_{n},\tau_{n}\mid\sigma_{n}^{n}=\tau_{n}^{2}=1,\tau_{n}\sigma_{n}\tau_{n}=\sigma_{n}^{-1}\rangle.$
Theorem 6.1.
Let $m$ be a positive integer. Consider a strongly reduced set
$\mathcal{H}:=\{\langle\tau_{2m}\rangle,\langle\sigma_{2m}\tau_{2m}\rangle\}$
of subgroups of $D_{2m}$. Then there is an exact sequence of $D_{2m}$-lattices
$0\rightarrow J_{D_{2m}/\mathcal{H}}\rightarrow\mathbb{Z}[D_{2m}]\oplus\mathbb{Z}\rightarrow\mathbb{Z}[D_{2m}/\langle\sigma\rangle]\rightarrow 0.$
In particular, $J_{D_{2m}/\mathcal{H}}$ is quasi-permutation.
Theorem 1.8 follows from Theorem 6.1 by the same argument as Theorem 1.2.
Proof.
Consider two three homomorphisms of $D_{n}$-lattices as follows:
$\iota_{m}\colon\mathbb{Z}[D_{2m}/\langle\sigma_{2m}\rangle]\rightarrow\mathbb{Z}[D_{2m}]\oplus\mathbb{Z};\,1\mapsto(1+\cdots+\sigma_{2m}^{2m-1},-1);$
$\omega_{m}\colon\mathbb{Z}[D_{2m}]\rightarrow I_{D_{2m}/\mathcal{H}};\,1\mapsto(1,-1);$
$\psi_{m}\colon\mathbb{Z}\rightarrow I_{D_{2m}/\mathcal{H}};\,1\mapsto(1+\cdots+\sigma_{2m}^{2m-1},-(1+\cdots+\sigma_{2m}^{2m-1})).$
It suffices to prove that the sequence of $D_{n}$-lattices
(6.1) $0\rightarrow\mathbb{Z}[D_{2m}/\langle\sigma_{2m}\rangle]\xrightarrow{\iota_{m}}\mathbb{Z}[D_{2m}]\oplus\mathbb{Z}\xrightarrow{(\omega_{m},\psi_{m})}I_{D_{2m}/\mathcal{H}}\rightarrow 0$
is exact. Inclusion $\operatorname{Im}(\iota_{m})\subset\operatorname{Ker}(\omega_{m},\psi_{m})$ can be confirmed by direct computation. For reverse inclusion, pick an element $x=(x_{1},x_{2})$ from $\operatorname{Ker}(\omega_{m},\psi_{m})$. Then we have
$x-x_{2}\iota_{m}(1)=(x_{1}-x_{2}(1+\cdots+\sigma_{2m}^{2m-1}),0).$
This implies that the first factor of $x-x_{2}\iota_{m}(1)$ is contained in $\operatorname{Ker}(\omega_{m})$. However, since $\operatorname{Ker}(\omega_{m})$ is generated by $(1-\tau_{2m})(1+\cdots+\sigma_{2m}^{2m-1})$ as an abelian group, we see that $x-x_{2}\iota_{m}(1)$ is a multiple of $\iota_{m}(1-\tau_{2m})$ by an integer. Consequently, one has $x\in\operatorname{Im}(\iota_{m})$ as desired. $\blacksquare$
Remark 6.2.
If $m=1$, then the exact sequence (6.1) is given by [End01, p. 30] (note that the map $\nu$ in [End01, p.30, l.18] needs to be modified as $\nu(1)=(1+\tau,-1-\sigma)$).
Lemma 6.3.
Let $G$ be a finite group, and
$0\rightarrow F\rightarrow R\xrightarrow{\Phi}M\rightarrow 0$
a coflabby resolution of a $G$-lattice $M$. Consider a homomorphism of $G$-lattices $\psi\colon R^{\prime}\rightarrow M$, where $R^{\prime}$ is a permutation $G$-lattice, and we denote by $\Phi^{\prime}$ the sum of $\Phi$ and $\psi$. Then there is an exact sequence
(6.2) $0\rightarrow F\oplus R^{\prime}\xrightarrow{\iota}R\oplus R^{\prime}\xrightarrow{\Phi^{\prime}}M\rightarrow 0$
which satisfies a commutative diagram
[…]
In particular, (6.2) is also a coflabby resolution of $M$.
Proof.
Consider the commutative diagram
[…]
where the horizontal sequences are exact. Applying the snake lemma to this diagram, we get an exact sequence
(6.3) $0\rightarrow F\rightarrow F^{\prime}\rightarrow R^{\prime}\rightarrow 0.$
Here $F^{\prime}$ is the kernel of $\Phi^{\prime}$. However, one has $\operatorname{Ext}_{\mathbb{Z}[G]}^{1}(R^{\prime},F)=0$ by [CS77, Lemme 1]. Hence (6.3) splits, and the proof is complete. $\blacksquare$
Lemma 6.4.
Consider a commutative diagram of finite free abelian groups
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}$${M_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{2}}$${M_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{3}}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M^{\prime}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M^{\prime}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M^{\prime}_{3},}$
where the horizontal sequences are exact. We further assume that
  • $\operatorname{rank}_{\mathbb{Z}}(M_{1})=\operatorname{rank}_{\mathbb{Z}}(M^{\prime}_{1})$;
  • $f_{2}$ and $f_{3}$ are injective; and
  • the cokernel of $f_{2}$ is torsion-free.
Then the homomorphism $f_{1}$ is an isomorphism.
Proof.
By the snake lemma, one has an exact sequence
$0\rightarrow\operatorname{Ker}(f_{1})\rightarrow\operatorname{Ker}(f_{2})\rightarrow\operatorname{Ker}(f_{3})\rightarrow\operatorname{Coker}(f_{1})\rightarrow\operatorname{Coker}(f_{2}).$
Since $\operatorname{Ker}(f_{2})$ is trivial by assumption, we have $\operatorname{Ker}(f_{1})=0$. On the other hand, since $\operatorname{Ker}(f_{3})=0$ and $\operatorname{Coker}(f_{2})$ is torsion-free, we obtain that $\operatorname{Coker}(f_{1})$ has no torsion. However, the equality $\operatorname{rank}_{\mathbb{Z}}(M_{1})=\operatorname{rank}_{\mathbb{Z}}(M^{\prime}_{1})$ implies $\operatorname{rank}_{\mathbb{Z}}(\operatorname{Coker}(f_{1}))=0$, and hence $\operatorname{Coker}(f_{1})=0$. This completes the proof. $\blacksquare$
Lemma 6.5.
Let $G$ be a finite group, and $N$ its normal subgroup. Consider a commutative diagram of $G$-lattices as follows:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}$${M_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{2}}$${M_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{3}}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M^{\prime}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M^{\prime}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M^{\prime}_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
where the horizontal sequences are exact. We further assume that
  • (a)
    $H^{1}(N,\operatorname{Ker}(f_{1}))=0$; and
  • (b)
    NN acts trivially on $M_{3}$.
Then $\operatorname{Ker}(f_{2})$ is generated by $\operatorname{Ker}(f_{1})$ and $\operatorname{Ker}(f_{2})^{N}$.
Proof.
We denote by $M_{3}^{\dagger}$ the image of $\operatorname{Ker}(f_{2})$ under the surjection $M_{2}\twoheadrightarrow M_{3}$. Then one has an exact sequence
(6.4) $0\rightarrow\operatorname{Ker}(f_{1})\rightarrow\operatorname{Ker}(f_{2})\rightarrow M_{3}^{\dagger}\rightarrow 0.$
Since $M^{\dagger}_{3}$ is contained in $M_{3}$, the assumption (b) implies that the action of $N$ on $M_{3}^{\dagger}$ is trivial. Taking $N$-fixed parts of (6.4), we obtain an exact sequence
$0\rightarrow\operatorname{Ker}(f_{1})^{N}\rightarrow\operatorname{Ker}(f_{2})^{N}\rightarrow M_{3}^{\dagger}\rightarrow 0.$
Here we use (a) for the surjectivity. This implies the desired assertion. $\blacksquare$
Lemma 6.6.
Let $G$ be a finite group, $H$ its subgroup, and $N_{1}$ and $N_{2}$ normal subgroups of $G$. Consider homomorphisms of $G$-lattices as follows:
$f_{1}\colon M_{1}\oplus\mathbb{Z}[G/H]^{N_{1}}\rightarrow M^{\prime}_{1}\oplus\mathbb{Z}[G/H];$
$f_{2}\colon M_{2}\oplus\mathbb{Z}[G/H]^{N_{2}}\rightarrow M^{\prime}_{1}\oplus\mathbb{Z}[G/H]\oplus M^{\prime}_{2}.$
Denote by $M^{\prime}\subset M^{\prime}_{1}\oplus\mathbb{Z}[G/H]\oplus M^{\prime}_{2}$ the image of the sum of $f_{1}$ and $f_{2}$. We further assume that
  • (1)
    $\gcd((HN_{1}N_{2}:HN_{1}),(HN_{1}N_{2}:N_{2}))=1$;
  • (2)
    $H\cap N_{1}N_{2}=\{1\}$;
  • (3)
    there exist commutative diagrams
    ${\mathbb{Z}[G/H]^{N_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x\mapsto(0,x)\hskip 15.0pt}$$\scriptstyle{x\mapsto x}$${M_{1}\oplus\mathbb{Z}[G/H]^{N_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}$${\mathbb{Z}[G/H]}$${M^{\prime}_{1}\oplus\mathbb{Z}[G/H],\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{pr}_{2}\hskip 10.0pt}$${\mathbb{Z}[G/H]^{N_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x\mapsto(0,x)\hskip 15.0pt}$$\scriptstyle{x\mapsto x}$${M_{2}\oplus\mathbb{Z}[G/H]^{N_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{2}}$${\mathbb{Z}[G/H]}$${M^{\prime}_{1}\oplus\mathbb{Z}[G/H]\oplus M^{\prime}_{2};\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{pr}_{2}\hskip 25.0pt}$
  • (4)
    $M^{\prime}\cap(M^{\prime}_{1}\oplus\{0\}\oplus M^{\prime}_{2})^{N_{1}N_{2}}\subset f_{1}(M_{1}\oplus\{0\})+f_{2}(M_{2}\oplus\{0\})\subset M^{\prime}_{1}\oplus\{0\}\oplus M^{\prime}_{2}$; and
  • (5)
    $\operatorname{rank}_{\mathbb{Z}}\operatorname{Ker}(f)=(G:HN_{1}N_{2})$.
Then there is an isomorphism
$M^{\prime}\oplus\mathbb{Z}[G/HN_{1}N_{2}]\cong M_{1}\oplus M_{2}\oplus\mathbb{Z}[G/HN_{1}]\oplus\mathbb{Z}[G/HN_{2}].$
Proof.
For each $i\in\{1,2\}$, the homomorphism $\operatorname{Ind}_{H}^{G}\varepsilon_{HN_{i}/H}^{\circ}$ induces an isomorphism
$\varepsilon_{i}\colon\mathbb{Z}[G/HN_{i}]\xrightarrow{\cong}\mathbb{Z}[G/H]^{N_{i}}.$
Take two integers $c_{1}$ and $c_{2}$ which satisfy $c_{1}(HN_{1}N_{2}:N_{1})-c_{2}(HN_{1}N_{2}:N_{2})=1$. Note that it is possible by (2). We define $\pi$ as the composite
$(M_{1}\oplus\mathbb{Z}[G/H]^{N_{1}})\oplus(M_{2}\oplus\mathbb{Z}[G/H]^{N_{2}})\xrightarrow{(\varepsilon_{1}^{-1}\circ\operatorname{pr}_{2})\oplus(\varepsilon_{2}^{-1}\circ\operatorname{pr}_{2})}$ $\,\,\mathbb{Z}[G/HN_{1}]\oplus\mathbb{Z}[G/HN_{2}]$
$\xrightarrow{(c_{1}\operatorname{Ind}_{HN_{1}N_{2}}^{G}\varepsilon_{HN_{1}N_{2}/HN_{1}},\,c_{2}\operatorname{Ind}_{HN_{1}N_{2}}^{G}\varepsilon_{HN_{1}N_{2}/HN_{2}})}$ $\,\,\mathbb{Z}[G/HN_{1}N_{2}]$
We denote by $f$ the sum of $f_{1}$ and $f_{2}$, and put $E:=\operatorname{Ker}(f)$.
Claim.
The restriction of $\pi$ to $E$ is an isomorphism.
We first prove the surjectivity. Put
$y:=\operatorname{Ind}_{H}^{G}\varepsilon_{HN_{1}N_{2}/H}^{\circ}(1)\in\mathbb{Z}[G/H]^{N_{1}N_{2}}.$
Then we have the following:
$\pi((0,y),(0,-y))=c_{1}(HN_{1}N_{2}:HN_{1})-c_{2}(HN_{1}N_{2}:HN_{2})=1,$
$f((0,y),(0,-y))\in M^{\prime}\cap(M^{\prime}_{1}\oplus\{0\}\oplus M^{\prime}_{2})^{N_{1}N_{2}}.$
Here we use (2) and (3) for the lower assertion. By (4), there exist $x_{1}\in M_{1}$ and $x_{2}\in M_{2}$ so that $f((x_{1},0),(x_{2},0))=f((0,y),(0,-y))$. Then $((-x_{1},y),(-x_{2},-y))$ lies in $E$ and
$\pi((-x_{1},y),(-x_{2},-y))=\pi((0,y),(0,-y))=1.$
Now, (5) implies that the kernel of the restriction of $\pi$ to $E$ has rank 0. Hence it must be trivial, and the proof of Claim is complete. $\blacksquare$
By Claim, we get a commutative diagram
[…]
where the horizontal sequence is exact. In particular, this exact sequence splits, and therefore we obtain an isomorphism
$M^{\prime}\oplus\mathbb{Z}[G/HN_{1}N_{2}]\cong M_{1}\oplus M_{2}\oplus\mathbb{Z}[G/H]^{N_{1}}\oplus\mathbb{Z}[G/H]^{N_{2}}.$
This implies the desired assertion. $\blacksquare$
The following generalizes Theorem 6.1 in the case that $n$ is a $2$-power.
Theorem 6.7.
Let $m$ be an odd positive, and $\nu\in\mathbb{Z}_{>0}$. Consider the finite group
$G_{m,\nu}:$ $=C_{m}\times D_{2^{\nu}}$
$=\langle\rho_{m},\sigma_{2^{\nu}},\tau_{2^{\nu}}\mid\rho_{m}^{m}=\sigma_{2^{\nu}}^{2^{\nu}}=\tau_{2^{\nu}}^{2}=1,\rho_{m}\sigma_{2^{\nu}}=\sigma_{2^{\nu}}\rho_{m},\rho_{m}\tau_{2^{\nu}}=\tau_{2^{\nu}}\rho_{m},\tau_{2^{\nu}}\sigma_{2^{\nu}}\tau_{2^{\nu}}^{-1}=\sigma_{2^{\nu}}^{-1}\rangle,$
and define $\mathcal{H}:=\{\langle\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle\}$. Then the $G_{m,\nu}$-lattice $J_{G_{m,\nu}/\mathcal{H}}$ is quasi-permutation.
Proof.
In this proof, put $I_{m,\nu}:=I_{G_{m,\nu}/\mathcal{H}_{m,\nu}}$. By Theorem 6.1, we may assume $m>1$. Consider homomorphisms of $G$-lattices as follows:
$\psi_{m,\nu,0}\colon R_{m,\nu,0}:=\mathbb{Z}[G_{m,\nu}/\langle\rho_{m}\rangle]\rightarrow I_{m,\nu};\,1\mapsto\left(\sum_{i=0}^{m-1}\rho_{m}^{i},-\sum_{i=0}^{m-1}\rho_{m}^{i}\right);$
$\psi_{m,\nu,1}\colon R_{m,\nu,1}:=\mathbb{Z}\rightarrow I_{m,\nu};\,1\mapsto\left(\left(\sum_{i=0}^{m-1}\rho_{m}^{i}\right)\left(\sum_{j=0}^{2^{\nu}-1}\sigma_{2^{\nu}}^{j}\right),-\left(\sum_{i=0}^{m-1}\rho_{m}^{i}\right)\left(\sum_{j=0}^{2^{\nu}-1}\sigma_{2^{\nu}}^{j}\right)\right);$
$\psi_{m,\nu,2}\colon R_{m,\nu,2}:=\mathbb{Z}[G_{m,\nu}/\langle\tau_{2^{\nu}}\rangle]\rightarrow I_{m,\nu};\,1\mapsto(1+\rho_{m},-(1+\sigma_{2^{\nu}}^{-1}));$
$\omega_{m,\nu,i}\colon R^{\prime}_{m,\nu,i}:=\mathbb{Z}[G_{m,\nu}/\langle\sigma_{2^{\nu}}^{2^{\nu-i}n^{\prime}},\sigma\tau\rangle]\rightarrow I_{m,\nu};\,1\mapsto\left(0,\left(\sum_{j=0}^{2^{i}-1}\sigma_{2^{\nu}}^{2^{\nu-i}j}\right)(1-\rho_{m}\sigma_{2^{\nu}}^{2^{\nu-i-1}})\right).$
Here $i\in\{0,\ldots,\nu-1\}$. We denote by $\Phi_{m,\nu,\ast}\colon R_{m,\nu,\ast}\rightarrow I_{m,\nu}$ the sum of $\psi_{m,\nu,i}$ for all $i\in\{0,1,2\}$ and $\omega_{m,\nu,0}$. Then there is an isomorphism $\operatorname{Coker}(\Phi_{m,\nu,\ast})\cong(\mathbb{Z}/m\mathbb{Z})^{\oplus 2^{\nu-1}-1}$, which follows from the following:
$\left(1,\sum_{j=0}^{(m-3/2)}\rho_{m}^{1+2j}(1+\sigma_{2^{\nu}}^{-1})-\sum_{i=0}^{m-1}\rho_{m}^{i}\right)=\psi_{m,\nu,0}(1)-\sum_{j=0}^{(m-3)/2}\psi_{m,\nu,2}(\rho_{m}^{1+2j});$
$(0,1-\rho_{m})=\omega_{m,\nu,0}((1+\rho_{m}^{2}+\cdots+\rho_{m}^{m-1})(1+\rho_{m}\sigma_{2^{\nu}}^{2^{\nu-1}}));$
$(0,1-\sigma_{2^{\nu}}^{2^{\nu-1}})=\omega_{m,\nu,0}(1+\rho_{m}\sigma_{2^{\nu}}^{2^{\nu-1}}+\cdots+(\rho_{m}\sigma_{2^{\nu}}^{2^{\nu-1}})^{m-1});$
$m(0,1-\sigma_{2^{\nu}}^{-1})=\psi_{m,\nu,0}(\tau_{2^{\nu}}-1)+(1-\sigma_{2^{\nu}}^{-1})\sum_{i=1}^{m-1}(0,1-\rho_{m}^{i}).$
On the other hand, we write for $\Psi_{m,\nu,\bullet}\colon R^{\prime}_{m,\nu,\bullet}\rightarrow I_{m,\nu}$ for the sum of $\omega_{m,\nu,i}$ for all $i\in\{1,\ldots,n-1\}$. Moreover, $\Phi_{m,\nu}\colon R_{m,\nu}\rightarrow I_{m,\nu}$ denotes the sum of $\Phi_{m,\nu,\ast}$ and $\Phi_{m,\nu,\bullet}$. Then, for $n\geq 2$, we have
$2^{i}(0,1-\rho\sigma_{2^{\nu}}^{2^{\nu-i-1}})=\omega_{m,\nu,i}(1)+\sum_{j=1}^{2^{\nu}-1}(0,1-\rho\sigma^{2^{\nu-i}j})$
for any $i\in\{1,\ldots,\nu-1\}$. Hence, by induction we obtain that the sum $R_{m,\nu}\rightarrow I_{m,\nu}$ is surjective. In particular, on has an exact sequence of $G$-lattices
$0\rightarrow U_{m,\nu}\rightarrow R_{m,\nu}\rightarrow I_{m,\nu}\rightarrow 0.$
In the following, we write $C_{m}$ and $Z_{\nu}$ for the subgroups of $G_{m,\nu}$ generated by $\rho_{m}$ and $\sigma_{2^{\nu}}^{2^{\nu-1}}$ respectively. In addition, we define $G_{m,\nu}$-sublattices of $R_{m,\nu}$ as follows:
$R_{m,\nu,\ast}^{-}:=R_{m,\nu,0}\oplus R_{m,\nu,1}\oplus R_{m,\nu,2},\quad R_{m,\nu}^{-}:=R_{m,\nu,\ast}^{-}\oplus R^{\prime}_{m,\nu,\bullet},$
Claim.
  1. (i)
    There exist isomorphisms of $G_{m,\nu}$-lattices
    (6.5) $U_{m,\nu}\cap R_{m,\nu,\ast}^{-}$ $\cong\mathbb{Z}[G_{m,\nu}/\langle\rho_{m},\sigma_{2^{\nu}}\rangle]\oplus R_{m,\nu,2}^{C_{m}},$
    (6.6) $U_{m,\nu}\cap R_{m,\nu,\ast}$ $\cong(U_{m,\nu}\cap R_{m,\nu,\ast}^{-})\oplus(R^{\prime}_{m,\nu,0})^{C_{m}}.$
    Moreover, the following holds:
    ${(R^{\prime}_{m,\nu,0})^{C_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x\mapsto(0,x)\hskip 50.0pt}$${(U_{m,\nu}\cap R_{m,\nu,\ast}^{-})\oplus(R^{\prime}_{m,\nu,0})^{C_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{\prime}_{m,n,0}}$${R_{m,\nu,\ast}^{-}\oplus R^{\prime}_{m,n,0}.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{pr}_{2}}$
    Here the left vertical map is the natural inclusion, and the right vertical map is defined by (6.6).
  2. (ii)
    We have $U_{m,\nu}=(U_{m,\nu}\cap R_{n,\ast})+U_{m,\nu}^{Z_{\nu}}$.
Proof of Claim.
(i): By the definition of $\Phi_{m,\nu,\ast}$, we have a commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U_{m,\nu}\cap R_{m,\nu,\ast}^{C_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R_{m,\nu,\ast}^{C_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{m,\nu,\ast}\hskip 12.0pt}$${I_{G_{m,\nu}/\mathcal{H}_{m,\nu}}^{C_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U_{m,\nu}\cap R_{m,\nu,\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R_{m,\nu,\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{m,\nu,\ast}\hskip 12.0pt}$${I_{G_{m,\nu}/\mathcal{H}_{m,\nu}},}$
where the horizontal sequences are exact. Note that the central and the rightmost vertical maps are injective, and the cokernel of the central vertical map is torsion-free. Moreover, since the the cokernel of $\Phi_{m,\nu,\ast}$ is torsion, we have
$\operatorname{rank}_{\mathbb{Z}}(U_{m,\nu}\cap R_{m,\nu,\ast})$ $=\operatorname{rank}_{\mathbb{Z}}(R_{m,\nu,\ast})-\operatorname{rank}_{\mathbb{Z}}(I_{m,\nu})$
$=(2^{\nu+1}(m+1)+1)-(2^{\nu+1}m-1)$
$=2^{\nu+1}+2.$
Therefore, the ranks of $U_{m,\nu}\cap R_{m,\nu,\ast}$ and $U_{m,\nu}\cap R_{m,\nu,\ast}^{C_{m}}$ coincide. Hence Lemma 6.4 implies
$U_{m,\nu}\cap R_{m,\nu,\ast}^{C_{m}}=U_{m,\nu}\cap R_{m,\nu,\ast}.$
On the other hand, Theorem 6.1 gives an exact sequence of $G_{m,\nu}$-lattices
$0\rightarrow\mathbb{Z}[G_{m,\nu}/\langle\rho_{m},\sigma_{2^{\nu}}\rangle]\rightarrow R_{m,\nu,0}\oplus R_{m,\nu,1}\rightarrow I_{G_{m,\nu}/\mathcal{H}_{m,\nu}}^{C_{m}}\rightarrow 0.$
Hence the assertion follows from Lemma 6.3.
(ii): By (i), $U_{m,\nu}\cap R_{m,\nu,\ast}=\operatorname{Ker}(\Phi_{m,\nu,\ast})$ is a permutation $G_{m,\nu}$-lattice. In particular, we have $H^{1}(Z_{\nu},U_{m,\nu}\cap R_{m,\nu,\ast})=0$. Moreover, one has $(R^{\prime}_{m,\nu,\bullet})^{Z_{\nu}}=R^{\prime}_{m,\nu,\bullet}$ by definition. Hence the assertion follows follows from Lemma 6.5 for the commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R_{m,\nu,\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{m,\nu,\ast}}$${R_{m,\nu}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\prime}_{m,\nu}}$$\scriptstyle{\Phi_{m,\nu}}$${R^{\prime}_{m,\nu,\bullet}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${I_{m,\nu}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${I_{m,\nu}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
where $\pi^{\prime}_{m,\nu}$ is the canonical projection $R_{m,\nu}\twoheadrightarrow R^{\prime}_{m,\nu,\bullet}$. $\blacksquare$
In the following, we prove that there is an isomorphism of $G_{m,\nu}$-lattices
(6.7) $U_{m,\nu}\oplus S_{m,\nu}\cong(U_{m,\nu}\cap R_{m,\nu,\ast})\oplus S^{\prime}_{m,\nu}\oplus S_{m,\nu},$
where
$S_{m,\nu}:=\bigoplus_{i=1}^{\nu-1}\mathbb{Z}[G_{m,\nu}/\langle\rho_{m},\sigma_{2^{\nu}}^{2^{\nu-i}},\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle],\quad S^{\prime}_{m,\nu}:=\bigoplus_{i=1}^{\nu-1}\mathbb{Z}[G_{m,\nu}/\langle\sigma_{2^{\nu}}^{2^{\nu-i}},\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle].$
This clearly implies the desired assertion. We prove (6.7) by induction on $n$. If $n=1$, then the $G_{m,\nu}$-lattice $R_{m,\nu,\bullet}$ is zero. Hence the assertion follows from Claim (i). Next, let $n\geq 2$, and suppose that the assertion holds for n1n-1. By definition, there is a canonical isomorphism $I_{m,\nu-1}\cong I_{m,\nu}^{Z_{\nu}}$. Moreover, the restriction of $\Phi_{m,\nu}$ to $R_{m,\nu}^{-}$ induces a coflabby resolution of $I_{m,\nu}^{Z_{\nu}}$ as follows:
$0\rightarrow U_{m,\nu}\cap(R_{m,\nu}^{-})^{Z_{\nu}}\rightarrow(R_{m,\nu}^{-})^{Z_{\nu}}\rightarrow I_{m,\nu}^{Z_{\nu}}\rightarrow 0.$
Consequently, Lemma 6.3 gives an isomorphism
$U_{m,\nu}^{Z_{\nu}}\cong(U_{m,\nu}\cap(R_{m,\nu}^{-})^{Z_{\nu}})\oplus(R^{\prime}_{m,\nu,0})^{Z_{\nu}}$
that satisfy the following:
(6.8) ${(R^{\prime}_{m,\nu,0})^{Z_{\nu}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x\mapsto x}$${U_{m,\nu}^{Z_{\nu}}\oplus S_{m,\nu-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{\prime}_{m,\nu,0}}$${R_{m,\nu}^{-}\oplus R^{\prime}_{m,\nu,0}\oplus S_{m,\nu-1}.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{pr}_{2}\hskip 40.0pt}$
On the other hand, by the induction hypothesis, we obtain an isomorphism of $G_{m,\nu}$-lattices
$(U_{m,\nu}\cap(R_{m,\nu}^{-})^{Z_{\nu}})\oplus S_{m,\nu-1}\cong(U_{m,\nu}\cap R_{m,\nu,\diamond}^{Z_{\nu}})\oplus S^{\prime}_{m,\nu-1}\oplus S_{m,\nu-1},$
where $R_{m,\nu,\diamond}:=R_{m,\nu,0}\oplus R_{m,\nu,1}\oplus R_{m,\nu,2}\oplus R^{\prime}_{m,\nu,1}$. In addition, by Claim (i), one has an isomorphism
$U_{m,\nu}\cap R_{m,\nu,\diamond}^{Z_{\nu}}\cong(U_{m,\nu}\cap(R_{m,\nu,\ast}^{-})^{Z_{\nu}})\oplus(R^{\prime}_{m,\nu,1})^{C_{m}}.$
In summary, we obtain an isomorphism of $G_{m,\nu}$-lattices
(6.9) $(U_{m,\nu}\cap(R_{m,\nu}^{-})^{Z_{\nu}})\oplus S_{m,\nu-1}\cong(U_{m,\nu}\cap(R_{m,\nu,\ast}^{-})^{Z_{\nu}})\oplus U_{m,\nu-1}^{\dagger}.$
where
$U_{m,\nu-1}^{\dagger}:=(R^{\prime}_{m,\nu,1})^{C_{m}}\oplus S^{\prime}_{m,\nu-1}\oplus S_{m,\nu-1}=S^{\prime}_{m,\nu-1}\oplus S_{m,\nu}.$
Now, consider two homomorphisms as follows:
$\Xi_{m,\nu,1}\colon U_{m,\nu}\cap R_{m,\nu,\ast}=(U_{m,\nu}\cap R_{m,\nu,\ast}^{-})\oplus(R^{\prime}_{m,\nu,0})^{C_{m}}\rightarrow R_{m,\nu,\ast}=R_{m,\nu,\ast}^{-}\oplus R^{\prime}_{m,\nu,0};$
$\Xi_{m,\nu,2}\colon U_{m,\nu-1}^{\dagger}\oplus(R^{\prime}_{m,\nu,0})^{Z_{\nu}}\rightarrow R_{m,\nu}\oplus S_{m,\nu-1}=R_{m,\nu,\ast}^{-}\oplus R^{\prime}_{m,\nu,0}\oplus(R^{\prime}_{m,\nu,\bullet}\oplus S_{m,\nu-1}),$
where $\Xi_{m,\nu,1}$ and $\Xi_{m,\nu,2}$ are induced by $U_{m,\nu}\cap R_{m,\nu,\ast}\subset R_{m,\nu,\ast}$ and (6.9) respectively. Denote by $\Xi_{m,\nu}$ the sum of $\Xi_{m,\nu,1}$ and $\Xi_{m,\nu,2}$. Then we have $\operatorname{Im}(\Xi_{m,\nu})=U_{m,\nu}\oplus S_{m,\nu-1}$, which is a consequence of Claim (ii) and (6.9). In particular, one has an exact sequence
(6.10) $0\rightarrow E_{m,\nu}\rightarrow(U_{m,\nu}\cap R_{m,\nu,\ast})\oplus U_{m,\nu-1}^{\dagger}\xrightarrow{\Xi_{m,\nu}}U_{m,\nu}\oplus S_{m,\nu-1}\rightarrow 0.$
Now, we confirm that $\Xi_{m,\nu,1}$ and $\Xi_{m,\nu,2}$ satisfy the five assumptions in Lemma 6.6. The conditions (1) and (2) are not difficult by using the property that $m$ is odd. Moreover, (3) is a consequence of Claim (i) and (6.8). On the other hand, (6.9) implies the following:
$\Xi_{m,\nu,1}(U_{m,\nu}\cap(R_{m,\nu,\ast}^{-})^{Z_{\nu}})+\Xi_{m,\nu,2}(U_{m,\nu-1}^{\dagger})$ $=(U_{m,\nu}\cap(R_{m,\nu}^{-})^{Z_{\nu}})\oplus S_{m,\nu-1}$
$=(U_{m,\nu}\oplus S_{m,\nu-1})\cap(R_{m,\nu}^{-}\oplus S_{m,\nu-1})^{Z_{\nu}}.$
Note that the second equality follows from the fact that $Z_{\nu}$ acts trivially on $S_{m,\nu-1}$. Hence (4) is valid. Finally, (6.10) implies $\operatorname{rank}_{\mathbb{Z}}(\operatorname{Ker}(\Xi_{m,\nu}))=2^{\nu-1}m$, that is, (5) holds true. Therefore, we can apply Lemma 6.6, and hence one has an isomorphism
$U_{m,\nu}\oplus S_{m,\nu-1}\oplus\mathbb{Z}[G_{m,\nu}/\langle\rho_{m},\sigma_{2^{\nu}}^{2^{\nu-1}},\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle]\cong(U_{m,\nu}\cap R_{m,\nu,\ast})\oplus U_{m,\nu-1}^{\dagger}\oplus(R^{\prime}_{m,\nu,0})^{Z_{\nu}}.$
Then the equality
$S_{m,\nu-1}\oplus\mathbb{Z}[G_{m,\nu}/\langle\rho_{m},\sigma_{2^{\nu}}^{2^{\nu-1}},\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle]=S_{m,\nu}$
and an isomorphism
$U_{m,\nu-1}^{\dagger}\oplus(R^{\prime}_{m,\nu,0})^{Z_{\nu}}=S^{\prime}_{m,\nu-1}\oplus S_{m,\nu}\oplus(R^{\prime}_{m,\nu,0})^{Z_{\nu}}\cong S^{\prime}_{m,\nu}\oplus S_{m,\nu}$
imply (6.7). Hence the proof is complete. $\blacksquare$
6.2. Non-quasi-invertible lattices
we prove that some $J_{G/\mathcal{H}}$ are not quasi-invertible. The proofs are based on [End01, §§3–4] and [End11, Lemma 2.2]. For a $G$-lattice $M$, put $M_{2}:=M\otimes_{\mathbb{Z}}\mathbb{Z}_{2}$ where $\mathbb{Z}_{2}$ is the ring of $2$-adic integers.
The following will be used in the next proposition.
Lemma 6.8.
Let $G:=(C_{2})^{2}=\langle\sigma,\tau\mid\sigma^{2}=\tau^{2}=1,\sigma\tau=\tau\sigma\rangle$, $\mathcal{H}:=\{\{1\},G\}$, and $\varphi$ the weight function on $\mathcal{H}$ defined as $\varphi(\{1\})=1$ and $\varphi(G)=2$. Then there is an exact sequence of $G$-lattices
$0\rightarrow\mathbb{Z}^{\oplus 2}\rightarrow\mathbb{Z}[G/\langle\sigma\rangle]\oplus\mathbb{Z}[G/\langle\tau\rangle]\oplus\mathbb{Z}[G/\langle\sigma\tau\rangle]\rightarrow I_{G/\mathcal{H}}^{(\varphi)}\rightarrow 0.$
Proof.
Consider homomorphisms as follows:
$f_{1}\colon\mathbb{Z}[G/\langle\sigma\rangle]\rightarrow I_{G/\mathcal{H}}^{(\varphi)};$  $\,1\mapsto(1+\sigma,-1);$
$f_{2}\colon\mathbb{Z}[G/\langle\tau\rangle]\rightarrow I_{G/\mathcal{H}}^{(\varphi)};$  $\,1\mapsto(1+\tau,-1);$
$f_{3}\colon\mathbb{Z}[G/\langle\sigma\tau\rangle]\rightarrow I_{G/\mathcal{H}}^{(\varphi)};$  $\,1\mapsto(1+\sigma\tau,-1).$
We set
$f\colon\mathbb{Z}[G/\langle\sigma\rangle]\oplus\mathbb{Z}[G/\langle\tau\rangle]\oplus\mathbb{Z}[G/\langle\sigma\tau\rangle]\rightarrow I_{G/\mathcal{H}}^{(\varphi)};\,(x_{1},x_{2},x_{3})\mapsto f_{1}(x_{1})+f_{2}(x_{2})+f_{3}(x_{3}).$
By definition, the kernel of the sum of $f_{1}$ is generated by $(1+\tau,-(1+\sigma),0)$ and $(0,1+\sigma,-(1+\sigma))$. These elements are fixed under $G$, and hence the assertion holds. $\blacksquare$
Proposition 6.9.
Let $G:=(C_{2})^{3}$ and $\mathcal{H}:=\{C_{2}\times\{1\}\times\{1\},\{1\}\times C_{2}\times C_{2}\}$. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
Write
$G=\langle\rho,\sigma,\tau\mid\rho^{2}=\sigma^{2}=\tau^{2}=1,\rho\sigma=\sigma\rho,\sigma\tau=\tau\sigma,\tau\rho=\rho\tau\rangle.$
We may assume $\mathcal{H}:=\{H_{1},H_{2}\}$, where $H_{1}=\langle\rho\rangle$ and $H_{2}=\langle\sigma,\tau\rangle$. Let $I=I_{G/\mathcal{H}}$ and $J=J_{G/\mathcal{H}}$. We regard $I$ as a $G$-submodule of $\mathbb{Z}[G/H_{1}]\oplus\mathbb{Z}[G/H_{2}]$, which induces an exact sequence
(6.11) $0\rightarrow I\rightarrow\mathbb{Z}[G/H_{1}]\oplus\mathbb{Z}[G/H_{2}]\xrightarrow{(\varepsilon_{G/H_{1}},\,\varepsilon_{G/H_{2}})}\mathbb{Z}\rightarrow 0.$
Put $\overline{G}:=G/H_{1}\cong(C_{2})^{2}$. By definition, there is an isomorphism of $\overline{G}$-lattices
$I^{H_{1}}\cong I_{\overline{G}/\overline{\mathcal{H}}}^{(\varphi)}.$
Here $\overline{\mathcal{H}}:=\{\{1\},\overline{G}\}$ and $\varphi$ is the weight function on $\overline{\mathcal{H}}$ defined as $\varphi(\{1\})=1$ and $\varphi(\overline{G})=2$. Hence, by Lemma 6.8, we obtain a coflabby resolution
$0\rightarrow\mathbb{Z}^{\oplus 2}\rightarrow\mathbb{Z}[G/\langle\rho,\sigma\rangle]\oplus\mathbb{Z}[G/\langle\rho,\tau\rangle]\oplus\mathbb{Z}[G/\langle\rho,\sigma\tau\rangle]\xrightarrow{f^{\prime}}I^{\langle\rho\rangle}\rightarrow 0.$
On the other hand, we have the following for any $g\in\{\sigma,\tau,\sigma\tau\}$:
$I^{\langle g\rangle}=I^{\langle\rho,g\rangle}+\langle(0,1-\rho)\rangle_{\mathbb{Z}},\quad I^{\langle g\rangle}=I^{\langle\rho,g\rangle}.$
Moreover, $I$ is generated by (1,1)(1,-1) as a $G$-lattice. Consider homomorphisms of $G$-lattices
$f_{1}\colon\mathbb{Z}[G]\rightarrow I$ ; $;\,1\mapsto(1,-1);$
$f_{2}\colon\mathbb{Z}[G/\langle\sigma,\tau\rangle]\rightarrow I$ ; $;\,1\mapsto(0,1-\rho).$
Then $f^{\prime}$, $f_{1}$ and $f_{2}$ induce a coflabby resolution of $I$:
$0\rightarrow U\rightarrow R\xrightarrow{f}I\rightarrow 0.$
Claim.
The exponent of $\widehat{H}^{0}(G,F)$ is a divisor of $4$.
Proof of Claim.
Consider an exact sequence
$\widehat{H}^{-1}(G,I)\rightarrow\widehat{H}^{0}(G,U)\rightarrow\widehat{H}^{0}(G,R).$
Then we have $\widehat{H}^{-1}(G,I)=0$, which follows from the surjectivity of the horizontal homomorphisms of the commutative diagram
${\widehat{H}^{-2}(G,\mathbb{Z})\oplus\widehat{H}^{-2}(G,\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hskip 35.0pt(\operatorname{Cor}_{G/H_{1}},\,\operatorname{Cor}_{G/H_{2}})}\scriptstyle{\cong}$${\widehat{H}^{-2}(G,\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H_{1}\oplus H_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G.}$
Moreover, there is an isomorphism $\widehat{H}^{0}(G,P)\cong(\mathbb{Z}/4\mathbb{Z})^{\oplus 4}$. This implies that the exponent of $\widehat{H}^{0}(G,U)$ is a divisor of $4$. $\blacksquare$
Now suppose that $J$ is quasi-invertible, that is, $U^{\circ}$ is invertible. Then $F$ is also invertible. Moreover, by the same argument as [End11, Lemma 2.2], $U_{2}$ is a permutation $\mathbb{Z}_{2}[G]$-lattice. Hence $F_{2}$ contains $\mathbb{Z}_{2}$ as a direct summand of $\mathbb{Z}_{2}[G]$-modules since $\operatorname{rank}_{\mathbb{Z}}(U)=11$. In particular, $\widehat{H}^{0}(G,U)$ has exponent $8$, which contradicts Claim. Therefore, the $G$-lattice $J$ is not quasi-invertible as desired. $\blacksquare$
Proposition 6.10.
Let $G:=C_{4}\times C_{2}$ and $\mathcal{H}:=\{C_{4}\times\{1\},\{1\}\times C_{2}\}$. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
Write
$G=\langle\sigma,\tau\mid\sigma^{4}=\tau^{2}=1,\sigma\tau=\tau\sigma\rangle,$
and put $H_{1}:=\langle\sigma\rangle$ and $H_{2}:=\langle\tau\rangle$. We may assume $\mathcal{H}=\{H_{1},H_{2}\}$. Let $I=I_{G/\mathcal{H}}$ and $J=J_{G/\mathcal{H}}$. We regard $I$ as a sublattice of $\mathbb{Z}[G/\langle\sigma\rangle]\oplus\mathbb{Z}[G/\langle\tau\rangle]$. Then one has an exact sequence
(6.12) $0\rightarrow I\rightarrow\mathbb{Z}[G/H_{1}]\oplus\mathbb{Z}[G/H_{2}]\xrightarrow{(\varepsilon_{G/H_{1}},\,\varepsilon_{G/H_{2}})}\mathbb{Z}\rightarrow 0.$
We regard $I$ as a sublattice of $\mathbb{Z}[G/H_{1}]\oplus\mathbb{Z}[G/H_{2}]$. Write
$G=\langle\sigma,\tau\mid\sigma^{4}=\tau^{2}=1,\sigma\tau=\tau\sigma\rangle.$
Then we have
$I=\langle(1-\tau,0),(0,1-\sigma),(0,\sigma(1-\sigma)),(0,\sigma^{2}(1-\sigma)),(1,-1)\rangle_{\mathbb{Z}}.$
Moreover, the following hold:
$I^{\langle\sigma^{2}\rangle}=\langle(1-\tau,0),(1+\tau,-(1+\sigma^{2})),(1+\tau,-\sigma(1+\sigma^{2}))\rangle_{\mathbb{Z}},$
$I^{\langle\tau\rangle}=\langle(0,1-\sigma),(0,\sigma(1-\sigma)),(0,\sigma^{2}(1-\sigma)),(1+\tau,-(1+\sigma^{2}))\rangle_{\mathbb{Z}},$
$I^{\langle\sigma^{2}\tau\rangle}=I^{\langle\sigma^{2},\tau\rangle}=\langle(1+\tau,-(1+\sigma^{2})),(1+\tau,-\sigma(1+\sigma^{2})))\rangle_{\mathbb{Z}}=\mathbb{Z}[G/\langle\sigma^{2},\tau\rangle](1+\tau,-(1+\sigma^{2})),$
$I^{\langle\sigma\rangle}=\langle(1-\tau,0),(2(1+\tau),-(1+\sigma+\sigma^{2}+\sigma^{3}))\rangle_{\mathbb{Z}},$
$I^{\langle\sigma\tau\rangle}=I^{G}=\langle(2(1+\tau),-(1+\sigma+\sigma^{2}+\sigma^{3}))\rangle_{\mathbb{Z}},$
Now, consider homomorphisms of $G$-lattices as follows:
$f_{1}\colon\mathbb{Z}[G]\rightarrow I$ ; $;\,1\mapsto(1,-1),$
$f_{2}\colon\mathbb{Z}[G/\langle\tau\rangle]\rightarrow I$ ; $;\,1\mapsto(0,1-\sigma),$
$f_{3}\colon\mathbb{Z}[G/\langle\sigma^{2},\tau\rangle]\rightarrow I$ ; $;\,1\mapsto(1+\tau,-(1+\sigma^{2})),$
$f_{4}\colon\mathbb{Z}[G/\langle\sigma\rangle]\rightarrow I$ ; $;\,1\mapsto(1-\tau,0).$
Then the sum of these maps induce a coflabby resolution
$0\rightarrow U\rightarrow R\xrightarrow{f}I\rightarrow 0$
of $I$ which admits an isomorphism
(6.13) $U^{\langle\sigma^{2}\tau\rangle}\cong\mathbb{Z}[G/\langle\sigma^{2}\tau\rangle]\oplus\mathbb{Z}[G/\langle\sigma^{2},\tau\rangle]\oplus\mathbb{Z}.$
Now, suppose that $J$ is quasi-invertible. Since $\operatorname{rank}_{\mathbb{Z}}(U)=11$ and $\operatorname{rank}_{\mathbb{Z}}(F^{G})=3$, there exist a subgroup $H$ of $G$ of order $4$ and an isomorphism
$U_{2}\cong\mathbb{Z}_{2}[G]\oplus\mathbb{Z}_{2}[G/H]\oplus\mathbb{Z}_{2}.$
In particular, we obtain an isomorphism
$\widehat{H}^{0}(G,U)\cong\mathbb{Z}/4\mathbb{Z}\oplus\mathbb{Z}/8\mathbb{Z}.$
However, since $\widehat{H}^{0}(G,R)\cong\mathbb{Z}/2\mathbb{Z}\oplus(\mathbb{Z}/4\mathbb{Z})^{\oplus 2}$, the same argument as Claim in Proposition 6.9 implies that the exponent of $\widehat{H}^{0}(G,U)$ is a divisor of $4$. Hence, we obtain a contradiction, and the proof is complete. $\blacksquare$
7. Proof of Theorem 1.3
We give a proof of Theorem 1.3 by dividing into four steps.
7.1. First step: The case for groups of order 88
Proposition 7.1.
Let $G$ be a $p$-group, where $p$ is a prime number. Consider a reduced set $\mathcal{H}$ of subgroups of $G$ which contains a normal subgroup $H_{0}$ of $G$ of index $\mu(\mathcal{H})$ such that $G/H_{0}$ is not cyclic. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
Let $\varphi$ be the weight function on $\mathcal{H}$ that takes the value $1$ for all $H\in\mathcal{H}$. By Proposition 3.22, one has an isomorphism
$J_{G/\mathcal{H}}^{[H_{0}]}\cong J_{G/\mathcal{H}^{H_{0}}}^{(\overline{\varphi}_{G/H_{0}})}.$
Note that $\overline{\varphi}_{G/H_{0}}$ is normalized since $\overline{\varphi}_{G/H_{0}}(\{1\})=1$. Moreover, since $(G:H_{0})=\mu(H)$, all the factors of $\overline{\varphi}_{G/H_{0}}(\overline{H})$ are divisible by $(\pi^{-1}(\overline{H}):H_{0})$ for any $\overline{H}\in\mathcal{H}^{H_{0}}$. Here $\pi$ denotes the natural surjection from $G$ onto $G/H_{0}$. Therefore Corollary 3.19 implies the existence of permutation $G$-lattices $R_{1}$ and $R_{2}$ and an isomorphism of $G$-lattices
$J_{(G/H_{0})/\mathcal{H}^{H_{0}}}^{(\overline{\varphi}_{G/H_{0}})}\oplus R_{1}\cong J_{G/H_{0}}\oplus R_{2}.$
However, the $G$-lattice $J_{G/H_{0}}$ is not quasi-invertible, which is a consequence of Proposition 5.2. Hence the assertion follows from Proposition 2.6 (ii). $\blacksquare$
Lemma 7.2.
Let $p$ be a prime number, and $G$ a $p$-group. Consider a reduced set of subgroups $\mathcal{H}$ of $G$. If $\mathcal{H}$ contains all maximal subgroups that are not cyclic, then $N^{G}(\mathcal{H})$ contains $\Phi(G)$.
Proof.
If all maximal subgroups of $G$ are cyclic, then the assertion follows from the definition of $\Phi(G)$. From now on, assume that $G$ admits its cyclic maximal subgroup. Pick $H\in\mathcal{H}$ which is not contained in any non-cyclic maximal subgroup of $G$. It suffices to prove that $H$ is maximal in $G$. Take a maximal subgroup $P$ of $G$ containing $H$. Then the assumption on $\mathcal{H}$ implies that $P$ is cyclic. On the other hand, Proposition 4.1 implies that $\Phi(G)$ contains the unique subgroup of $P$ of index $p$. Hence we must have $H=P$, which completes the proof. $\blacksquare$
Proposition 7.3.
Let $G:=(C_{2})^{3}$, and $\mathcal{H}$ be a reduced set of its subgroups with $\mathcal{H}\geq 2$ and $N^{G}(\mathcal{H})=\{1\}$. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
For $i\in\{2,4\}$, put
$r_{i}:=\#\{H\in\mathcal{H}\mid(G:H)=i\}.$
By Propositions 5.3 and 7.1, we may assume that $r_{2}$ and $r_{4}$ are grater than 0.
Case 1. $r_{2}=r_{4}=1$.
In this case, the assertion follows from Proposition 6.9.
Case 2. $r_{4}\geq 2$.
Pick $H_{0},H_{1},H_{2}\in\mathcal{H}$ which satisfy $(G:H_{0})=2$ and $(G:H_{1})=(G:H_{2})=4$. Put $P:=H_{1}H_{2}$, which is isomorphic to $(C_{2})^{2}$. Then we have $(P:P\cap H_{0})=2$ and $(P\cap H_{0})\cap H_{1}=(P\cap H_{0})\cap H_{2}=\{1\}$. Hence we have $\mathcal{H}_{P}^{\operatorname{red}}=\{P\cap H_{0},H_{1},H_{2}\}$. On the other hand, Proposition 5.3 implies that the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{red}}}$ is not quasi-invertible. Therefore, the assertion follows from Proposition 2.6.
Case 3. $r_{2}\geq 2$.
Write $G=\langle\rho,\sigma,\tau\mid\rho^{2}=\sigma^{2}=\tau^{2}=1,\rho\sigma=\sigma\rho,\sigma\tau=\tau\sigma,\tau\rho=\rho\tau\rangle$. Take $H_{1},H_{2},H_{3}\in\mathcal{H}$ with $(G:H_{1})=(G:H_{2})=2$ and $(G:H_{3})=4$. We may assume
$H_{1}=\langle\rho,\sigma\rangle,\quad H_{2}=\langle\sigma,\tau\rangle,\quad H_{3}=\langle\rho\tau\rangle.$
Now put $P:=\langle\rho\sigma,\sigma\tau\rangle$, which is isomorphic to $(C_{2})^{2}$. Note that $P$ is not contained in $\mathcal{H}$ since $H_{3}\subset P$ and $\mathcal{H}$ is reduced. Then we have
$P\cap H_{1}=\langle\rho\sigma\rangle,\quad P\cap H_{2}=\langle\sigma\tau\rangle,\quad P\cap H_{3}=H_{3}=\langle\rho\tau\rangle.$
Therefore, the assertion follows from the same argument as Case 2. $\blacksquare$
Proposition 7.4.
Let $G:=C_{4}\times C_{2}$, and $\mathcal{H}$ a reduced set of its subgroups with $\mathcal{H}\geq 2$ and $N^{G}(\mathcal{H})=\{1\}$. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
By Propositions 5.3 and 7.1, we may assume that $\mathcal{H}$ contains a subgroup $H_{0}$ of index $4$ such that $G/H_{0}$ is cyclic. Write $G=\langle\sigma,\tau\mid\sigma^{4}=\tau^{2}=1,\sigma\tau=\tau\sigma\rangle$, and set
$N:=\langle\sigma^{2},\tau\rangle.$
Since $N$ is the unique non-cyclic maximal subgroup of $G$, it must not be contained in $\mathcal{H}$ by Lemma 7.2. Consequently we may consider the case where $\mathcal{H}$ satisfies at least one of (i)–(iv) as follows:
  1. (i)
    $\mathcal{H}=\{\langle\sigma\rangle,\langle\tau\rangle\}$;
  2. (ii)
    $\mathcal{H}=\{\langle\sigma\rangle,\langle\sigma\tau\rangle,\langle\tau\rangle\}$;
  3. (iii)
    $\mathcal{H}=\{\langle\tau\rangle,\langle\sigma^{2}\tau\rangle\}$;
  4. (iv)
    $\{\langle\tau\rangle,\langle\sigma^{2}\tau\rangle\}\subsetneq\mathcal{H}$.
In the following, we prove the assertion case-by-case.
Case (i): In this case, the assertion follows from Proposition 6.10.
Case (ii): It suffices to prove that there is an isomorphism
(7.1) $_{\omega}^{2}(G,J_{G/\mathcal{H}})\cong\mathbb{Z}/2\mathbb{Z}.$
In fact, if it holds, then Proposition 2.10 gives the desired assertion. Put
$H_{0}:=\langle\sigma\rangle,\quad H_{1}:=\langle\sigma\tau\rangle,\quad H_{2}:=\langle\tau\rangle,$
By the global class field theory, we can take a finite abelian extension $L/\mathbb{Q}$ with group $G$. Let $\mathbf{K}:=\prod_{i=0}^{2}K_{i}$, where $K_{i}$ is the $H_{i}$-fixed subfield of $L$ for each $i\in\{0,1,2\}$. Then, combining Proposition 3.1 with [BP20, Lemma 3.3], we get an isomorphism
[…]
Here the right-hand side is defined as follows:
$_{\omega}^{2}(\mathbb{Q},X^{*}(T_{\mathbf{K}/\mathbb{Q}})):=\{x\in H^{2}(\mathbb{Q},X^{*}(T_{\mathbf{K}/\mathbb{Q}}))\mid\operatorname{Res}_{\mathbb{Q}_{v}/\mathbb{Q}}(x)=0\text{ for almost all place $v$ of $\mathbb{Q}$}\}.$
Therefore, we can apply the theory of [Lee22]. First, note that the numbering of elements of $\mathcal{H}$ satisfies the assumption in [Lee22, p. 8, l. 6]. Put $\mathcal{I}:=\{1,2\}$. Since $H_{0}H_{1}=H_{0}H_{2}=G$, we have $U_{0}=\mathcal{I}$. Furthermore, one has
$n_{l}(\mathcal{I}):=\#(\mathcal{I}/\sim_{l})=\begin{cases}1&\text{if }l=0,\\ 2&\text{if }l>0.\end{cases}$
In particular, we obtain
  • $L(\mathcal{I})=0$ (see [Lee22, p. 17, Section 5, l. 17–18]); and
  • $n_{l+1}(c)=1$ for every $l>0$ and $c\in\mathcal{I}/\sim_{l}$.
On the other hand, since $H_{0}$ contains $H_{1}\cap N=\Phi(G)$, the integer $f_{\mathcal{I}}^{\omega}$ in [Lee22, p. 18, l.10–12] must be $1$. Here we use the fact that $N$ is the unique subgroup of index $2$ containing $H_{2}$. Therefore [Lee22, Corollary 6.3 (1)] gives an isomorphism
$_{\omega}^{2}(\mathbb{Q},X^{*}(T_{K/\mathbb{Q}}))\cong(\mathbb{Z}/2\mathbb{Z}^{f_{\mathcal{I}}^{\omega}})^{\oplus n_{1}(\mathcal{I})-1}=\mathbb{Z}/2\mathbb{Z},$
which concludes the proof of (7.1).
Case (iii): This is the same as [End11, Lemma 2.2].
Case (iv): By assumption, there is $H\in\mathcal{H}$ which has index $2$ in $G$. We may assume $H=\langle\sigma\rangle$. Then $\mathcal{H}_{N}^{\operatorname{red}}$ contains $P\cap H$, $H_{1}$ and $H_{2}$. Since they are distinct from each other, the assertion follows from the same argument as Case 1 in the proof of Proposition 7.3. $\blacksquare$
Proposition 7.5.
Let $G:=D_{4}$, and $\mathcal{H}$ be a reduced set of its subgroups satisfying $\mathcal{H}\geq 2$ and $N^{G}(\mathcal{H})=\{1\}$. We further assume that $\mathcal{H}^{\operatorname{nor}}$ is non-empty. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
By Proposition 5.3, we may assume that $\mathcal{H}$ contains a subgroup of index $4$. Furthermore, we may further assume that $\mathcal{H}$ does not contain $\langle\sigma_{4}^{2}\rangle$, which is a consequence of Proposition 7.1. Hence, it suffices to consider the case where $\mathcal{H}$ contains $\langle\tau_{4}\rangle$. combining this with the non-emptiness of $\mathcal{H}^{\operatorname{nor}}$, we obtain that $\mathcal{H}$ contains a subgroup $H_{0}$ of index $2$. Then $H_{0}$ coincides with $\langle\sigma_{4}\rangle$ or $\langle\sigma_{4}^{2},\sigma_{4}\tau_{4}\rangle$ since $\mathcal{H}$ is reduced. Now put $P:=\langle\sigma_{4}^{2},\tau_{4}\rangle$, which is isomorphic to $(C_{2})^{2}$. Then $\mathcal{H}_{P}^{\operatorname{red}}$ contains $\langle\sigma_{4}^{2}\rangle$, $\langle\tau_{4}\rangle$ and $\langle\sigma_{4}^{2}\tau_{4}\rangle$. Therefore, the same argument as Case 1 in the proof of Proposition 7.3 gives the desired assertion. $\blacksquare$
In summary, we obtain the following.
Theorem 7.6.
Let $G$ be a $2$-group of order $8$, and $\mathcal{H}$ a reduced set of its subgroups satisfying $\mathcal{H}\geq 2$ and $N^{G}(\mathcal{H})=\{1\}$. We further assume that $\mathcal{H}^{\operatorname{nor}}$ is non-empty. Then the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible.
Proof.
By the assumption on $\mathcal{H}$ and the classification of groups of order $8$, we obtain that $G$ is isomorphic to $(C_{2})^{3}$, $C_{4}\times C_{2}$ or $D_{4}$. Hence, the assertion follows from Propositions 7.3, 7.4 and 7.5. $\blacksquare$
7.2. Second step: The case consisting of normal subgroups with indices $\leq 4$
Here we generalize Theorem 7.6, which will be needed in the final step.
Proposition 7.7.
Let $G$ be a $2$-group, and $\mathcal{H}$ a reduced set of its subgroups that satisfy $\mu(\mathcal{H})=2$, $M(\mathcal{H})=4$ and $\mathcal{H}=\mathcal{H}^{\operatorname{nor}}$. Then the $G$-lattice $J_{G/H}$ is not quasi-invertible.
Proof.
Take $H_{1},H_{2}\in\mathcal{H}$ satisfying $(G:H_{1})=2$ and $(G:H_{2})=4$. Put $N:=H_{1}\cap H_{2}$, which has index $8$ in $G$. Write $\varphi$ for the weight function on $\mathcal{H}$ that takes the value $1$ for every $H\in\mathcal{H}$. Then Proposition 3.22 gives an isomorphism
$J_{G/\mathcal{H}}^{N}\cong J_{(G/N)/\mathcal{H}^{N}}^{(\overline{\varphi}_{G/N})}.$
Note that $\overline{\varphi}_{G/N}$ is normalized since $\overline{\varphi}_{G/N}(H_{1}/N)=\overline{\varphi}_{G/N}(H_{2}/N)=1$. Let $\mathcal{H}_{\overline{\varphi}_{G/N}}^{N}[1]$ be as in Corollary 3.19, that is,
$\mathcal{H}_{\overline{\varphi}_{G/N}}^{N}[1]:=\{H^{\prime}\in\mathcal{H}^{N}\mid d_{\overline{\varphi}_{G/N}}(H^{\prime})=1\}.$
Denote by $H^{\prime}_{i}$ the image of $H_{i}$ in $G/N$ for each $i\in\{1,2\}$. If $H^{\prime}\notin\mathcal{H}^{N}\setminus\mathcal{H}_{\overline{\varphi}_{G/N}}$ satisfies $(G/N:H^{\prime})=2$, then the assumption $M(\mathcal{H})=4$ implies an equality
$\overline{\varphi}_{G/N}(H^{\prime})=(\underbrace{2,\ldots,2}_{m_{\mathcal{H}^{N}}(H^{\prime})}).$
Hence we have $d_{\overline{\varphi}_{G/N}}(H^{\prime})=(H^{\prime}:H^{\prime}_{1}\cap H^{\prime})=2$. On the other hand, $\overline{\varphi}_{G/N}(G/N)\in\Delta_{m_{\mathcal{H}^{N}}(G/N)}$ is a sequence consisting of $2$ and $4$. Consequently, one has $d_{\overline{\varphi}_{G/N}}(G)=(G/N:H^{\prime}_{1})=2$. Therefore, we can apply Corollary 3.19. In particular we obtain an equality
$J_{(G/N)/\mathcal{H}^{N}}^{(\overline{\varphi}_{G/N})}]=[J_{(G/N)/\{H^{\prime}_{1},H^{\prime}_{2}\}}].$
However, the $G$-lattice $J_{(G/N)/\{H^{\prime}_{1},H^{\prime}_{2}\}}$ is not quasi-invertible by Theorem 7.6. Combining this result with Proposition 2.6 (ii), we obtain the desired assertion. $\blacksquare$
Lemma 7.8.
Let $p$ be a prime number, and $G$ a $p$-group. Consider a reduced set $\mathcal{H}$ of subgroups of $G$ satisfying $\mathcal{H}\geq 2$, $\mathcal{H}^{\operatorname{nor}}\neq\emptyset$ and $\mu(\mathcal{H})=M(\mathcal{H})\geq p^{2}$. Assume that a maximal subgroup $P$ of $G$ that satisfies $\mathcal{H}_{\subset P}=\emptyset$ and $\mathcal{H}_{P}^{\operatorname{red}}=1$. Then one has
$M(\mathcal{H})=p^{-1}\#G.$
Proof.
Pick $H\in\mathcal{H}^{\operatorname{nor}}$. Then the assumption on $P$ implies $P\cap H=P\cap H^{\prime}$ for any $H^{\prime}\in\mathcal{H}$. In particular, one has $P\cap N^{G}(\mathcal{H})=P\cap H$. Hence we have
$G=(G:N^{G}(\mathcal{H}))\leq(G:P\cap N^{G}(\mathcal{H}))=p(P:P\cap H)=p(G:H)=pM(\mathcal{H}).$
Combining the above inequality with $\mathcal{H}\geq 2$, we obtain $G=pM(\mathcal{H})$ as desired. $\blacksquare$
Lemma 7.9.
Let $p$ be a prime number, and $G$ a $p$-group. Fix $m\in\mathbb{Z}_{>0}$. Consider a multiset $\mathcal{H}$ of subgroups of $G$ satisfying $N^{G}(\mathcal{H})=\{1\}$. Suppose
  1. (i)
    $H\triangleleft G$ and $G/H\cong C_{p^{m}}$ for every $H\in\mathcal{H}$; and
  2. (ii)
    every maximal subgroup of $G$ contains an element of $\mathcal{H}$.
Then the group $G$ is isomorphic to a product of finite copies of $C_{p^{m}}$.
Proof.
Assume that the conclusion fails. Then there is an element $g\in G$ whose order is smaller than $p^{m}$ so that $g\neq h^{p}$ for all $h\in G$. Take a maximal subgroup $P$ of $G$ which does not contain $g$ and $H\in\mathcal{H}_{\subset P}$. Then the image $\overline{g}$ of $g$ in $G/H$ must be a generator. However, we have $\overline{g}^{p^{m-1}}=1$ since $g^{p^{m-1}}=1$. This contradicts the cyclicity of $G/H_{0}$, and hence the proof is complete. $\blacksquare$
Theorem 7.10.
Let $G$ be a $2$-group, and $\mathcal{H}$ a reduced set of its subgroups satisfying $\mathcal{H}\geq 2$, $M(\mathcal{H})=4$ and $\mathcal{H}^{\operatorname{nor}}=\mathcal{H}$. Then the $G$-lattice $J_{G/H}$ is not quasi-invertible.
Proof.
By Proposition 7.7, we may assume $\mu(\mathcal{H})=4$.
Step 1. Here we prove the assertion in the case where all maximal subgroups of $G$ contain a member of $\mathcal{H}$. By Proposition 7.1, we may further assume that $G/H$ is cyclic for any $H\in\mathcal{H}$. Then Lemma 7.9 gives an isomorphism $G\cong(C_{4})^{m}$ for some $m\in\mathbb{Z}_{>0}$. Take a maximal subgroup $P$ of $G$, which is possible by assumption. Then it is clear that $\mu(\mathcal{H}_{P}^{\operatorname{red}})=2$.
In the following, we shall give $H^{\prime}\in\mathcal{H}$ satisfying $P\cap H\in\mathcal{H}_{P}^{\operatorname{red}}$ and $(P:P\cap H^{\prime})=4$. Fix elements $g_{1},\ldots,g_{m}$ of $G$ satisfying
$G=\langle g_{1},\ldots,g_{m}\mid g_{1}^{4}=\cdots=g_{n}^{4}=1,g_{i}g_{j}=g_{j}g_{i}\,(i\neq j)\rangle$
and
$P=\langle g_{1}^{2},g_{2},\ldots,g_{m}\rangle.$
Pick $H\in\mathcal{H}_{P}$. Then there are $a_{2},\ldots,a_{m}\in\{0,2\}$ so that
$H=\langle g_{1}^{a_{2}}g_{2},\ldots,g_{1}^{a_{n}}g_{m}\rangle.$
Now let
$P^{\prime}:=\langle g_{1}^{2},g_{1}g_{2},g_{3},\ldots,g_{m}\rangle,$
and pick $H^{\prime}\in\mathcal{H}_{\subset P^{\prime}}$. We prove that this $H^{\prime}$ satisfies the desired properties. There exist $b_{2},\ldots,b_{m}\in\{0,2\}$ so that
$H^{\prime}=\langle g_{1}^{b_{2}+1}g_{2},g_{1}^{b_{3}}g_{3}\ldots,g_{1}^{b_{n}}g_{m}\rangle.$
In particular, one has an equality
$P\cap H^{\prime}=\langle g_{1}^{2}g_{2}^{2},g_{1}^{b_{3}}g_{3}\ldots,g_{1}^{b_{n}}g_{m}\rangle.$
Hence we obtain $P/(P\cap H^{\prime})\cong G/H^{\prime}\cong C_{4}$. Furthermore, this isomorphism implies that
$H_{0}:=\langle g_{1}^{2},g_{2}^{2},g_{3}\ldots,g_{m}\rangle$
is the unique subgroup of $P$ of index $2$ containing $P\cap H^{\prime}$. Hence, if $P\cap H_{3}$ does not lie in $\mathcal{H}_{P}^{\operatorname{red}}$, then the assumption $\mu(\mathcal{H})=4$ implies that $H_{0}$ is an element of $\mathcal{H}$. However, it is a contradiction since we assume the cyclicity of $G/H$ for all $H\in\mathcal{H}$. Therefore we obtain $P\cap H^{\prime}\in\mathcal{H}_{P}^{\operatorname{red}}$ as desired.
As above, we know that $\mu(\mathcal{H}_{P}^{\operatorname{red}})=2$ and $M(\mathcal{H}_{P}^{\operatorname{red}})=4$. Then it follows from Proposition 7.7 that the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{red}}}$ is not quasi-invertible. Hence the assertion follows from Proposition 3.20 and Proposition 2.6 (i).
Step 2. Let $G=2^{\nu}$, where $\nu\geq 3$ is an integer. Here we give a proof of the assertion in general by induction on $\nu$. If $\nu=3$, then the claim is contained in Theorem 7.6. In the following, suppose $\nu\geq 4$ and that the assertion holds for $\nu-1$. By Step 1, we may further assume that $\mathcal{H}_{\subset P}$ is empty for some maximal subgroup $P$ of $G$. Then we have $M(\mathcal{H}_{P}^{\operatorname{red}})=4$. Moreover, the inequality
$G=2^{\nu}\geq 2^{4}>8=2M(\mathcal{H})$
and Lemma 7.8 imply $\mathcal{H}_{P}^{\operatorname{red}}\geq 2$. By the induction hypothesis, the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{red}}}$ is not quasi-invertible. Therefore the assertion follows from Proposition 3.20 and Proposition 2.6 (i). $\blacksquare$
7.3. Third step: The case admitting normal factors
Theorem 7.11.
Let $G$ be a $2$-group, and $\mathcal{H}$ a reduced set of subgroups of $G$. We further assume that $\mathcal{H}^{\operatorname{nor}}$ is non-empty. Then the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation if and only if it is quasi-invertible. Moreover, the above two conditions hold if and only if
  1. (i)
    $J_{G/\mathcal{H}}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/\mathcal{H}}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $G$ and $\mathcal{H}$ satisfy one of the following:
    • (iii-1)
      $\mathcal{H}=1$ and $G/N^{G}(\mathcal{H})$ is cyclic; or
    • (iii-2)
      $\mathcal{H}=2$ and $G/N^{G}(\mathcal{H})\cong(C_{2})^{2}$.
Proof.
We may assume that $\mathcal{H}$ contains at least two elements. In particular, $G$ is not cyclic by Lemma 5.7. Write $G=2^{\nu}$. The assertion holds for $\nu\geq 2$, which is a consequence of Proposition 5.3. In the sequel of the proof, assume $\nu\geq 3$. It suffices to prove that $J_{G/H}$ is not quasi-invertible if $(G:N^{G}(\mathcal{H}))\geq 8$ or $\mathcal{H}\geq 3$.
We give a proof of the above assertion by induction on $\nu$. If $\nu=3$, the claim follows from Theorem 7.6. Now suppose $\nu\geq 4$ and the assertion holds for $\nu-1$. If $N^{G}(\mathcal{H})\neq\{1\}$, then the same argument as Theorem 5.10 implies that the assertion follows from the induction hypothesis. Hence we may further assume $N^{G}(\mathcal{H})=\{1\}$. By the induction hypothesis, it suffices to prove that there is a maximal subgroup $P$ of $G$ which satisfies $(P:N^{P}(\mathcal{H}_{P}^{\operatorname{red}}))\geq 8$ or $\mathcal{H}_{P}^{\operatorname{red}}\geq 3$.
Case 1. $\mathcal{H}=\mathcal{H}^{\operatorname{nor}}$.
By Theorem 7.10, we may assume $M(\mathcal{H})\geq 8$. Take $H_{0}\in\mathcal{H}$ with $(G:H_{0})=M(\mathcal{H})$. By the same argument as Case 1 in the proof of Theorem 5.10, there is a maximal subgroup $P$ of $G$ satisfying $H_{0}\subset P$ and $\mathcal{H}_{P}^{\operatorname{red}}\geq 2$. Then we have $M(\mathcal{H}_{P}^{\operatorname{red}})\geq M(\mathcal{H})/2\geq 4$. Hence the inequality $(P:N^{P}(\mathcal{H}_{P}^{\operatorname{red}}))\geq 8$ follows from Lemma 5.5.
Case 2. $\mathcal{H}^{\operatorname{nor}}$ and $\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}$ are non-empty.
Note that one has $M(\mathcal{H})\geq p^{2}$ by Proposition 4.1. Consider positive integers as follows:
(7.2) $\Sigma:=\min\{(G:H)\in\mathbb{Z}_{>0}\mid H\in\mathcal{H}^{\operatorname{nor}}\},\quad\sigma:=\min\{(G:H)\in\mathbb{Z}_{>0}\mid H\in\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}\}.$
Take $H_{0}\in\mathcal{H}^{\operatorname{nor}}$ with $(G:H_{0})=\Sigma$ and $H_{1}\in\mathcal{H}\setminus\mathcal{H}^{\operatorname{nor}}$ with $(G:H_{1})=\sigma$.
Case 2-a. $\Sigma<\sigma$. Pick a maximal subgroup $P$ of $G$ containing $N_{G}(H_{1})$. Then we have $gH_{1}g^{-1}\not\subset P\cap H_{0}\subsetneq H_{0}$ and $gH_{1}g^{-1}\not\supset P\cap H_{0}$ for any $g\in G$. Hence $\mathcal{H}_{P}^{\operatorname{red}}$ has at least $3$ elements.
Case 2-b. $\Sigma>\sigma$. Take a maximal subgroup $P$ of $G$ containing $H_{0}$. Then one has $H_{0}\not\subset P\cap H_{1}\subsetneq H_{1}$ and $H_{0}\not\supset P\cap H_{1}$. In particular, $P\cap H_{1}$ is contained in $\mathcal{H}_{P}^{\operatorname{red}}$ since $(G:H_{1})=\sigma=\mu(\mathcal{H})$. Therefore we obtain the desired inequality
$(P:N^{P}(\mathcal{H}_{P}^{\operatorname{red}}))\geq(P:H_{0}\cap(P\cap H_{1}))\geq 2(P:H_{0})=2\Sigma\geq 8.$
Case 2-c. $\Sigma=\sigma$. Let $P$ be a maximal subgroup of $G$ containing $N_{G}(H_{1})$. It suffices to prove that $\mathcal{H}_{P}^{\operatorname{red}}$ is not a subset of $\{gH_{1}g^{-1}\mid g\in G\}$. This implies $\mathcal{H}_{P}^{\operatorname{red}}\geq 3$ since $H_{1}$ is not normal in $G$. Assume not, that is, $\mathcal{H}_{P}^{\operatorname{red}}$ is contained in $\{gH_{1}g^{-1}\mid g\in G\}$. Since $H_{0}$ is normal of index $\Sigma=\sigma=\mu(\mathcal{H})$, it is contained in $N_{G}(H_{1})$ by Lemma 5.9. In particular, $H_{0}$ is an element of $\mathcal{H}_{P}^{\operatorname{red}}$, which contradicts to the assumption. Hence the proof is complete. $\blacksquare$
7.4. Final step: General case
Lemma 7.12.
[…]
Proof.
By assumption, we have $g_{2}=xg_{1}h$ for some $x\in E$ and $h\in H$. This implies an equality
$E\cap g_{2}Hg_{2}^{-1}=x(E\cap g_{1}Hg_{1}^{-1})x^{-1}.$
Now, the right-hand side equals $E\cap g_{1}Hg_{1}^{-1}$ since $E$ is abelian. Hence the proof is complete. $\blacksquare$
Proposition 7.13.
Let $G$ be a group of order $2^{\nu}$, where $\nu\geq 3$. Consider a strongly reduced set $\mathcal{H}$ of subgroups of $G$ satisfying $\mathcal{H}\geq 2$, $\mathcal{H}^{\operatorname{nor}}=\emptyset$ and $\mu(\mathcal{H})=\#G/2$. Then the following are equivalent:
  1. (i)
    $J_{G/\mathcal{H}}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/\mathcal{H}}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $G\cong D_{2^{\nu}}$ and $\mathcal{H}=\{\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle\}$ for some integers $m$ and $m^{\prime}$.
Proof.
If $G\cong D_{2^{\nu}}$, then the assumption on $\mathcal{H}$ implies $\mathcal{H}=\{\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle\}$ for some integers $m$ and $m^{\prime}$. Hence the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation by Theorem 6.1. Otherwise, by Proposition 4.4 and Lemma 4.7, the group $G$ is not of maximal class. This implies that we can take a non-cyclic abelian normal subgroup $E$ of $G$ of order $8$, which is a consequence of Proposition 4.5.
Case 1. $\mu(\mathcal{H}_{E}^{\operatorname{red}})=8$.
In this case, we have $\mathcal{H}_{E}^{\operatorname{red}}=\{\{1\}\}$ and $E/N^{E}(\mathcal{H}_{E}^{\operatorname{red}})$ is not cyclic. Then Proposition 5.2 gives the desired assertion.
Case 2. $\mu(\mathcal{H}_{E}^{\operatorname{red}})=4$ and $E\cong(C_{2})^{3}$.
By assumption, there is an isomorphism
$E/N^{E}(\mathcal{H}_{E}^{\operatorname{red}})\cong\begin{cases}(C_{2})^{2}&\text{if }\#\mathcal{H}_{E}^{\operatorname{red}}=1;\\ (C_{2})^{3}&\text{if }\#\mathcal{H}_{E}^{\operatorname{red}}\geq 2.\end{cases}$
Then the assertion follows from Theorem 7.11.
Case 3. $\mu(\mathcal{H}_{E}^{\operatorname{red}})=4$ and $E\cong C_{4}\times C_{2}$.
Write $E=\langle\sigma,\tau\mid\sigma^{4}=\tau^{2}=1,\sigma\tau=\tau\sigma\rangle$. Then, by Lemma 4.7, $\Phi(G)$ is contained in the center of $G$. Pick $H\in\mathcal{H}_{E}^{\operatorname{red}}$, then $\mu(\mathcal{H})=2$ implies $H\subset E$ and $H=g_{0}H_{0}g_{0}^{-1}$ for some $H_{0}\in\mathcal{H}$ and $g\in G$. Then, $H$ coincides with $\langle\tau\rangle$ or $\langle\sigma^{2}\tau\rangle$ since it is not normal in $G$. We may assume $H=\langle\tau\rangle$. On the other hand, the elements $\tau$ and $\sigma^{2}\tau$ are conjugate in $G$. In particular, there is $g\in G$ such that $gHg^{-1}=\langle\sigma^{2}\tau\rangle$. Then one has $Eg_{0}H_{0}\neq Egg_{0}H_{0}$ in $E\backslash H/H_{0}$, which is a consequence of Lemma 7.12. Therefore, $\mathcal{H}_{E}^{\operatorname{red}}$ contains $\langle\tau\rangle$ and $\langle\sigma^{2}\tau\rangle$. In particular, we have $\mathcal{H}_{N}^{\operatorname{red}}\geq 2$ and $N^{E}(\mathcal{H}_{E}^{\operatorname{red}})=\{1\}$. Now, the assertion follows from Theorem 7.11. $\blacksquare$
Lemma 7.14.
Let $G$ be a $2$-group, and $\mathcal{H}$ a strongly reduced set of its subgroups. Assume
  • $N^{G}(\mathcal{H})=\{1\}$;
  • $\mu(\mathcal{H})=M(\mathcal{H})$; and
  • $(H:N^{G}(H))=(N_{G}(H):H)=2$ and $N^{G}(H)\neq\{1\}$ for any $H\in\mathcal{H}$.
Then there is a maximal subgroup $P$ of $G$ such that $\mathcal{H}_{P}^{\operatorname{srd}}\geq 3$.
Proof.
By assumption, there exist $H,H^{\prime}\in\mathcal{H}$ such that $N^{G}(H)\neq N^{G}(H^{\prime})$.
Case 1. $N^{G}(H)H^{\prime}=G$.
Let $P$ be a maximal subgroup of $G$ containing $N_{G}(H)$. Fix $g\in G\setminus P$, then $\mathcal{H}_{P}^{\operatorname{srd}}$ contains $H$ and $gHg^{-1}$. On the other hand, we have
$(G:N^{G}(H)\cdot(P\cap H^{\prime}))\leq 2$
since $N^{G}(H)H^{\prime}=G$. Combining this inequality with the inclusion $N^{G}(H)\cdot(P\cap H^{\prime})\subset P$, we obtain an equality
$N^{G}(H)\cdot(P\cap H^{\prime})=P.$
In particular, $P\cap H^{\prime}$ is not contained in $H$ or $gHg^{-1}$. If $P\cap H^{\prime}\in\mathcal{H}_{P}^{\operatorname{red}}$, then $\mathcal{H}_{P}^{\operatorname{red}}\geq 3$ is clear. Otherwise, there exist $g_{0}\in G$ and $H_{0}\in\mathcal{H}$ so that $P\cap H^{\prime}\subsetneq P\cap g_{0}Hg_{0}^{-1}$. Since $(P\cap H^{\prime})=\#H^{\prime}/2=\#H_{0}/2$, we have $P\cap g_{0}Hg_{0}^{-1}=g_{0}H_{0}g_{0}^{-1}$, that is, $g_{0}H_{0}g_{0}^{-1}\subset P$. In particular, we obtain $H_{0}$ and $gH_{0}g^{-1}$ are contained in $P$. Therefore, one has $\mathcal{H}_{P}^{\operatorname{srd}}\geq 4$, which completes the proof in this case.
Case 2. $N^{G}(H)H^{\prime}\neq G$.
Let $P$ be a maximal subgroup of $G$ containing $N^{G}(H)H^{\prime}$. Since $(N_{G}(H^{\prime}):H^{\prime})=2$, Corollary 4.2 implies $N_{G}(H^{\prime})\subset P$. Fix $g\in G\setminus P$. If $H\subset P$, then the same argument as above yields $N_{G}(H^{\prime})\subset P$. Then, one has $H,H^{\prime},gHg^{-1},gH^{\prime}g^{-1}\in\mathcal{H}_{P}^{\operatorname{srd}}$, in particular $\mathcal{H}_{P}^{\operatorname{srd}}\geq 4$. If $H\not\subset P$, then we have $P\cap H=N^{G}(H)$ since $P$ contains $N^{G}(H)$. Therefore, the same argument as Case 1 implies the desired assertion. $\blacksquare$
Lemma 7.15.
Let $G$ be a $2$-group, and $\mathcal{H}$ a strongly reduced set of subgroups of $G$. Then the following are equivalent:
  1. (i)
    $G/N^{G}(\mathcal{H})\cong D_{2^{\nu}}$ and $\mathcal{H}=\{\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle\}$ for some integer $\nu\geq 2$ and $m,m^{\prime}\in\mathbb{Z}$;
  2. (ii)
    $\mathcal{H}$ satisfies all the conditions as follows:
    • $\mathcal{H}=2$;
    • $(G:N^{G}(\mathcal{H}))=2\mu(\mathcal{H})$;
    • $G/N^{G}(\mathcal{H})$ is a $2$-group of maximal class; and
    • $(H:N^{G}(H))=(N_{G}(H):H)=2$ for all $H\in\mathcal{H}$.
Proof.
(i) $\Rightarrow$ (ii) is clear by Lemma 4.6 (i). In the following, we prove the reverse implication. We may assume $N^{G}(\mathcal{H})=\{1\}$. Then $G$ is of maximal class. Hence Proposition 4.4 yields one of the following:
  • (a)
    $G\cong D_{2^{\nu}}$ for some $\nu\geq 2$;
  • (b)
    $G\cong SD_{2^{\nu+1}}$ for some $\nu\geq 3$; or
  • (c)
    $G\cong Q_{2^{\nu}}$ for some $\nu\geq 3$.
Moreover, since $\mathcal{H}=2$ and $(H:N^{G}(H))=2$ for all $H\in\mathcal{H}$, only (a) is valid by Lemma 4.7. Hence the assertion follows from Lemma 4.6 (ii), (iii). $\blacksquare$
Theorem 7.16.
Let $G$ be a $2$-group, and $\mathcal{H}$ a strongly reduced set of its subgroups. Then, the following are equivalent:
  1. (i)
    $J_{G/\mathcal{H}}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/\mathcal{H}}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $G$ and $\mathcal{H}$ satisfy one of the following:
    • (iii-I)
      $\mathcal{H}=1$ and $G/N^{G}(\mathcal{H})$ is cyclic; or
    • (iii-II)
      $G/N^{G}(\mathcal{H})\cong D_{2^{\nu}}$ and $\mathcal{H}=\{\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle\}$ for some $\nu\in\mathbb{Z}_{>0}$ and $m,m^{\prime}\in\mathbb{Z}$.
Proof.
If $\mathcal{H}=1$, then the assertion follows from Proposition 5.2. Hence, we may assume $\mathcal{H}\geq 2$.
(i) $\Rightarrow$ (ii): This is clear.
(iii) $\Rightarrow$ (i): If (iii-I) is valid, then Proposition 5.2 gives the desired assertion. On the other hand, if (iii-II) is satisfied, then the assertion follows from Theorem 6.1.
(ii) $\Rightarrow$ (iii): It suffices to prove that the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible if (iii-II) fails. Write $G=2^{\nu}$, which satisfies $\nu\geq 2$. We prove the assertion by induction on $\nu$. If $\nu\in\{2,3\}$, then the failure of (iii-II) implies that $\mathcal{H}^{\operatorname{nor}}$ is non-empty. Hence, the assertion follows from Theorem 7.11. In the following, suppose $\nu\geq 4$ and that the assertion holds for $\nu-1$. We may assume that $\mathcal{H}^{\operatorname{nor}}$ is empty, which is a consequence of Theorem 7.11. Moreover, it suffices to discuss the case $N^{G}(\mathcal{H})=\{1\}$. Furthermore, we may assume $\mathcal{H}^{\operatorname{nor}}=\emptyset$, which follows from Theorem 7.11. In addition, we only need a consideration for $\mu(\mathcal{H})<\#G/2$, which is a consequence of Proposition 7.13.
Case 1. $\mu(\mathcal{H})<M(\mathcal{H})$.
Take $H_{0}\in\mathcal{H}$ with $(G:H_{0})=M(\mathcal{H})$, and pick a maximal subgroup of $P$ of $G$ containing $N_{G}(H_{0})$. Fix $g\in G\setminus P$. Then $P\cap H$ does not contain $H_{0}$ or $gH_{0}g^{-1}$ for any $H\in\mathcal{H}$ with $(G:H)<M(\mathcal{H})$. This implies $\mathcal{H}_{P}^{\operatorname{srd}}\geq 3$, and hence the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{srd}}}$ is not quasi-invertible by the induction hypothesis and Lemma 7.15. Therefore the assertion follows from Proposition 3.20.
Case 2. $(N_{G}(H):H)\geq 4$ for some $H\in\mathcal{H}$.
Let $P$ be a maximal subgroup of $G$ containing $N_{G}(H)$. Pick $g\in G\setminus P$. Then $H$ and $gHg^{-1}$ lie in $\mathcal{H}^{\operatorname{srd}}$. Moreover, we have $N_{P}(H)=N_{G}(H)$ since $P$ contains $N_{G}(H)$. This implies $(N_{P}(H):H)\geq 4$, and hence the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{srd}}}$ is not quasi-invertible by the induction hypothesis and Lemma 7.15. Hence the assertion follows from Proposition 3.20.
Case 3. $(H:N^{G}(H))\geq 4$ for some $H\in\mathcal{H}$.
By assumption, we have $(G:H)\geq 8$. Take a maximal subgroup $P$ of $G$ containing $\Phi(G)H$. Then it also contains $N_{G}(H)$ by Corollary 4.2 (ii). Moreover, we have $N_{G}(H)\neq P$ since $(G:H)\geq 8$. Now, pick $g\in G\setminus P$, then $H$ and $gHg^{-1}$ are non-normal subgroups of $P$ that are not conjugate to each other. This is a consequence of the inclusion $N_{G}(H)\subsetneq P$ and the inequality $(G:H)\geq 8$. Moreover, they are contained in $\mathcal{H}_{P}^{\operatorname{red}}$. In particular, we have $\mathcal{H}_{P}^{\operatorname{red}}\geq 2$. If $\mathcal{H}_{P}^{\operatorname{srd}}\geq 3$, then the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{srd}}}$ is not quasi-invertible by the induction hypothesis. In the following, suppose $\mathcal{H}_{P}^{\operatorname{srd}}=2$, which implies $\mathcal{H}_{P}^{\operatorname{srd}}=\{H,gHg^{-1}\}$ for some $g\in G\setminus P$. Moreover, the equality $N^{G}(H)=N^{P}(H)\cap N^{P}(gHg^{-1})$ implies that
  • (a)
    $N^{P}(H)=N^{G}(H)$; or
  • (b)
    $N^{P}(H)\neq N^{G}(H)$ and $N^{P}(H)\neq N^{P}(gHg^{-1})$.
In both cases, the inequality $(G:N^{P}(\mathcal{H}_{P}^{\operatorname{srd}}))\geq 4\mu(\mathcal{H}_{P}^{\operatorname{srd}})$ follows. Now, the $P$-lattice $J_{P/\mathcal{H}_{P}^{\operatorname{srd}}}$ is not quasi-invertible by the induction hypothesis and Lemma 7.15. Consequently, the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible by Proposition 3.20.
Case 4. $\mu(\mathcal{H})=M(\mathcal{H})$ and $(H:N^{G}(H))=(N_{G}(H):H)=2$ for any $H\in\mathcal{H}$.
Since we assume $\mu(\mathcal{H})<\#G/2$, we have $N^{G}(H)\neq\{1\}$ for all $H\in\mathcal{H}$. Hence, Lemma 7.14 yields that there is a maximal subgroup $P$ of $G$ satisfying $\mathcal{H}_{P}^{\operatorname{srd}}\geq 3$. Then the $P$-lattice $J_{G/\mathcal{H}_{P}^{\operatorname{srd}}}$ is not quasi-invertible, which is a consequence of the induction hypothesis and Lemma 7.15. Now, Proposition 3.20 gives the desired assertion. $\blacksquare$
Proof of Theorem 1.3.
Put $G:=\operatorname{Gal}(L/k)$, and
$\mathcal{H}:=\{\operatorname{Gal}(L/K_{1}),\ldots,\operatorname{Gal}(L/K_{r})\}.$
Then, we have $N^{G}(\mathcal{H})=\{1\}$. Moreover, Proposition 3.1 gives an isomorphism $X^{*}(T_{\mathbf{K}/k})\cong J_{G/\mathcal{H}}$. Hence, Proposition 2.11 implies that $T_{\mathbf{K}/k}$ is stably (resp. retract) rational over $k$ if and only if $J_{G/\mathcal{H}}$ is a quasi-permutation (resp. quasi-invertible) $G$-lattice. Moreover, by Proposition 3.16, the $G$-lattice $J_{G/\mathcal{H}}]$ is quasi-permutation (resp. quasi-invertible) if and only if $J_{G/\mathcal{H}^{\operatorname{srd}}}]$ is so. On the other hand, we have the following:
  • (iii-1) holds if and only if $\mathcal{H}=\{\{1\}\}$ and $G$ is cyclic; and
  • (iii-2) holds if and only if $G\cong D_{2^{\nu}}$ with $\nu\in\mathbb{Z}_{\geq 0}$ and $\mathcal{H}^{\operatorname{srd}}=\{\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle\}$ for some $m,m^{\prime}\in\mathbb{Z}$.
Consequently, the assertion follows from Theorem 7.16. $\blacksquare$
8. Proof of Theorem 1.5
Definition 8.1.
Let $G$ be a finite group, and consider multisets $\mathcal{H}$ and $\mathcal{H}^{\prime}$ of subgroups in $G$. We write $\mathcal{H}^{\prime}<\mathcal{H}$ if
  • $\mathcal{H}^{\prime}$ is a reduced set; and
  • there is an injection $\Xi\colon\mathcal{H}^{\prime}\hookrightarrow\mathcal{H}^{\operatorname{set}}$ such that $\Xi(H^{\prime})\subset H^{\prime}$ for every $H^{\prime}\in\mathcal{H}^{\prime}$.
We can confirm the following without difficulty.
Lemma 8.2.
Let $G$ be a finite group, and $\mathcal{H}$ a multiset of its subgroups.
  1. (i)
    We have $\mathcal{H}^{\operatorname{red}}<\mathcal{H}$.
  2. (ii)
    Let $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ be reduced sets of subgroups in $G$ such that $\mathcal{H}^{\prime\prime}<\mathcal{H}^{\prime}<\mathcal{H}$. Then, we have $\mathcal{H}^{\prime\prime}<\mathcal{H}$.
Lemma 8.3.
Let $G=G_{p}\times G^{\prime}$, where $G_{p}$ is a $p$-group, and $G^{\prime}$ is a finite group of order coprime to $p$. Consider a multiset $\mathcal{H}$ of subgroups of $G$. Then, there is a reduced set $\mathcal{H}_{0}$ of $G$ such that
  1. (i)
    $\mathcal{H}_{0}<\mathcal{H}$;
  2. (ii)
    $(\mathcal{H}_{0})_{G_{p}}^{\operatorname{set}}=\mathcal{H}_{G_{p}}^{\operatorname{red}}$; and
  3. (iii)
    $J_{G/\mathcal{H}}]=[J_{G/\mathcal{H}_{0}}]$.
Proof.
(i): By Corollary 3.13 (ii) and Lemma 8.2, we may assume that $\mathcal{H}$ is reduced. Note that all elements of $\mathcal{H}_{G_{p}}$ are of the form $G_{p}\cap H$ for some $H\in\mathcal{H}$. Now, pick $H_{0}\in\mathcal{H}$ such that $G_{p}\cap H_{0}\notin\mathcal{H}_{G_{p}}^{\operatorname{red}}$. Take $H_{1}\in\mathcal{H}$ so that $G_{p}\cap H_{0}\subset N_{1}:=G_{p}\cap H_{1}\in\mathcal{H}_{G_{p}}^{\operatorname{red}}$. Then we have $H_{0}N_{1}\notin\mathcal{H}$ since $\mathcal{H}$ is strongly reduced. Moreover, $H_{1}\cap H_{0}N_{1}$ is strictly contained in $H_{0}N_{1}$, which follows from $H_{0}\not\subset H_{1}$. In particular, it is not contained in $\mathcal{H}\cup\{H_{0}N_{1}\}$. Now we consider the multiset $\widetilde{\mathcal{H}}$ of subgroup of $G$ which satisfies
  • $\widetilde{\mathcal{H}}^{\operatorname{set}}=\mathcal{H}\cup\{H_{0}N_{1},H_{1}\cap H_{0}N_{1}\}$;
  • $m_{\widetilde{\mathcal{H}}}(H_{0}N_{1})=2$; and
  • $m_{\widetilde{\mathcal{H}}}(H_{0})=1$ for each $H_{0}\in\widetilde{\mathcal{H}}^{\operatorname{set}}\setminus\{H_{0}N_{1}\}$.
Moreover, we write for $\varphi$ the weight function on $\widetilde{\mathcal{H}}$ defined as follows:
$\varphi(H)=\begin{cases}((H_{0}N_{1}:H_{0}),(H_{0}N_{1}:H_{1}\cap H_{0}N_{1}))&\text{if $H=H_{0}N_{1}$ and $\#H>\#(H_{1}\cap HN_{1})$;}\\ ((H_{0}N_{1}:H_{1}\cap H_{0}N_{1}),(H_{0}N_{1}:H_{0}))&\text{if $H=H_{0}N_{1}$ and $\#H_{0}\leq\#(H_{1}\cap H_{0}N_{1})$; }\\ 1&\text{otherwise. }\end{cases}$
Then Proposition 3.16 gives an isomorphism of $G$-lattices
$J_{G/\widetilde{\mathcal{H}}}\cong\begin{cases}J_{G/\mathcal{H}}\oplus\mathbb{Z}[G/H_{0}N_{1}]&\text{if }H_{1}\subset H_{0}N_{1};\\ J_{G/\mathcal{H}}\oplus\mathbb{Z}[G/H_{0}N_{1}]^{\oplus 2}\oplus\mathbb{Z}[G/(H_{1}\cap H_{0}N_{1})]&\text{if }H_{1}\not\subset HN_{1}.\end{cases}$
On the other hand, since $(H_{0}N_{1}:H_{0})$ is a power of $p$ and $(H_{0}N_{1}:H_{1}\cap H_{0}N_{1})$ is coprime to $p$, we obtain that $d_{\varphi}$ is the constant weight function on $\widetilde{\mathcal{H}}^{\operatorname{set}}$ that takes value $1$. Therefore, if we set $\mathcal{H}^{\dagger}:=\widetilde{\mathcal{H}}^{\operatorname{red}}$, then Lemma 3.10 and Corollary 3.13 (ii) follow that there is an equality
$J_{G/\widetilde{\mathcal{H}}}^{(\varphi)}]=[J_{G/\widetilde{\mathcal{H}}^{\operatorname{red}}}].$
In summary, one has
$J_{G/\mathcal{H}}]=[J_{G/\mathcal{H}^{\dagger}}].$
Note that the inclusions $H_{0}\subset H_{0}N_{1}$ and $H_{1}\cap H_{0}N_{1}\subset H_{0}N_{1}$ imply that $H_{0}$ and $H_{1}\cap H_{0}N_{1}$ are not contained in $\mathcal{H}$$^{\dagger}$. In particular, we have $\widetilde{\mathcal{H}}^{\operatorname{red}}\setminus\{H_{0}N_{1}\}\subset\mathcal{H}$. In addition, the assumption that $\mathcal{H}$ is reduced implies that $H_{0}\cap G^{\prime}$ is contained in $\mathcal{H}_{G^{\prime}}^{\operatorname{red}}$. Hence, $H_{0}N_{1}$ is contained in $\mathcal{H}$$^{\dagger}$. Now, set
$\Xi\colon\widetilde{\mathcal{H}}^{\operatorname{red}}\rightarrow\mathcal{H};\,H\mapsto\begin{cases}H_{0}&\text{if }H=H_{0}N_{1};\\ H&\text{otherwise. }\end{cases}$
Then, it is injective and $\Xi(H)\subset H$ for any $H\in\widetilde{\mathcal{H}}^{\operatorname{red}}$. Therefore, we obtain $\widetilde{\mathcal{H}}^{\operatorname{red}}<\mathcal{H}$. On the other hand, we have $({\mathcal{H}}^{\dagger})_{G_{1}}^{\operatorname{red}}=\mathcal{H}_{G_{1}}^{\operatorname{red}}$ and
$\{H\in\mathcal{H}^{\dagger}\mid G_{p}\cap H\notin\mathcal{H}_{G_{p}}^{\operatorname{red}}\}<\#\{H\in\mathcal{H}\mid G_{p}\cap H\notin\mathcal{H}_{G_{p}}^{\operatorname{red}}\}.$
Applying the above procedure for all $H_{0}\in\mathcal{H}$ with $G_{p}\cap H_{0}\notin\mathcal{H}_{G_{p}}^{\operatorname{red}}$, we obtain a strongly reduced set of subgroups $\mathcal{H}_{0}$ of $G$ satisfying (i), (ii) and (iii). $\blacksquare$
Theorem 8.4.
Let $G$ be a finite nilpotent group, and $\mathcal{H}$ a multiset of its subgroups. Then the following are equivalent:
  1. (i)
    $J_{G/\mathcal{H}}$ is a quasi-permutation $G$-lattice;
  2. (ii)
    $J_{G/\mathcal{H}}$ is a quasi-invertible $G$-lattice;
  3. (iii)
    $J_{G/\mathcal{H}}]=[J_{G/\mathcal{H}^{\prime}}]$, where $\mathcal{H}^{\prime}$ is a reduced set of subgroups of $G$ with $\mathcal{H}^{\prime}<\mathcal{H}$ that satisfies one of the following:
    • (iii-$\alpha$)
      $\mathcal{H}^{\prime}=\{N\}$, $N\triangleleft G$ and $G/N$ is cyclic; or
    • (iii-$\beta$)
      $\mathcal{H}^{\prime}=\{H,H^{\prime}\}$, $G/N^{G}(\mathcal{H}^{\prime})\cong C_{m}\times D_{2^{\nu}}$ for some $m\in\mathbb{Z}\setminus 2\mathbb{Z}$ and ${\nu}\in\mathbb{Z}_{>0}$, $H/N^{G}(\mathcal{H}^{\prime})\cong\langle(1,\tau_{2^{\nu}})\rangle$, and $H^{\prime}/N^{G}(\mathcal{H}^{\prime})\cong\langle(1,\sigma_{2^{\nu}}\tau_{2^{\nu}})\rangle$.
Proof.
(i) $\Rightarrow$ (ii): This is clear.
(ii) $\Rightarrow$ (iii): We may assume that $\mathcal{H}$ is strongly reduced. For each odd prime divisor $p$ of #G\#G, we denote by $G_{p}$ the unique $p$-Sylow subgroup of $G$. Then $J_{G/\mathcal{H}}$ is quasi-invertible as a $G_{p}$-lattice. On the other hand, the set $\mathcal{H}_{G_{p}}^{\operatorname{set}}$ consists of $G_{p}\cap H$ for all $H\in\mathcal{H}$. Hence, by Theorem 5.10, there is $H_{p}\in\mathcal{H}$ so that $N_{p}:=G_{p}\cap H_{p}\triangleleft G_{p}$, $G_{p}/N_{p}$ is cyclic and $G_{p}\cap H\subset N_{p}$ for any $H\in\mathcal{H}$. In particular, the subgroup $N_{p}$ is normal in $G$. Because $G$ is nilpotent, it is a product of $G_{p}$ for all prime divisors $p$ of #G\#G, Lemma 8.3 implies the existence of a strongly reduced set of subgroups $\mathcal{H}^{\prime}$ of $G$ so that
  • $(\mathcal{H}^{\prime})_{G_{p}}^{\operatorname{set}}=\{N_{p}\}$ for any odd prime $p$, where $N_{p}\triangleleft G_{p}$ and $G_{p}/N_{p}$ is cyclic; and
  • $J_{G/\mathcal{H}}]=[J_{G/\mathcal{H}^{\dagger}}]$.
Therefore, by replacing $\mathcal{H}$ and $G$ with $\mathcal{H}^{\dagger}$ and $G/N^{G}(\mathcal{H}^{\dagger})$, respectively, we may assume $G=C_{m}\times G_{2}$ for some odd integer $m$ and #H\#H is a power of $2$ for every $H\in\mathcal{H}$. In this case, the set of subgroups $\mathcal{H}_{G_{2}}^{\operatorname{set}}$ of $G_{2}$ is strongly reduced since $\mathcal{H}$ is so. Hence, the assumption that $J_{G/\mathcal{H}^{\prime\prime}}$ is quasi-invertible implies that
  1. (A)
    $\mathcal{H}_{G_{2}}^{\operatorname{set}}=\{N_{2}\}$ for some $N\triangleleft G_{2}$ so that $G_{2}/N_{2}$ is cyclic; or
  2. (B)
    $\mathcal{H}_{G_{2}}^{\operatorname{set}}=\{H,H^{\prime}\}$, $G_{2}/N^{G_{2}}(\mathcal{H}_{G_{2}}^{\operatorname{set}})\cong D_{2^{\nu}}$ for some $n\in\mathbb{Z}_{>0}$, $H/N^{G_{2}}(\mathcal{H}_{G_{2}}^{\operatorname{set}})\cong\langle\tau_{2^{\nu}}\rangle$ and $H^{\prime}/N^{G_{2}}(\mathcal{H}_{G_{2}}^{\operatorname{set}})\cong\langle\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle$.
If (A) holds, then the set $\mathcal{H}^{\prime}:=\{N_{2}\}$ satisfies (iii-$\alpha$). On the other hand, if (B) is valid, then the set $\mathcal{H}^{\prime}:=\{H,H^{\prime}\}$ is the one that satisfies (iii-$\beta$).
(iii) $\Rightarrow$ (i): This is a consequence of Theorem 6.7. $\blacksquare$
Now, we begin with the proof of Theorem 1.5 using Theorem 8.4. Let $k$ be a field, and $\mathbf{K}=\prod_{i=1}^{r}K_{i}$ a finite étale algebra over $k$. A subalgebra of $\mathbf{K}$ refers to a finite product $\prod_{j=1}^{s}K^{\prime}_{i_{j}}$, where $1\leq i_{1}<\cdots<i_{s}\leq r$ are integers and $K^{\prime}_{i_{j}}$ is an intermediate field of $K_{i_{j}}/k$ for each $j\in\{1,\ldots,s\}$.
Proof of Theorem 1.5.
It suffices to prove (ii) $\Rightarrow$ (iii) $\Rightarrow$ (i). Put $G:=\operatorname{Gal}(L/k)$ and
$\mathcal{H}:=\{\operatorname{Gal}(L/K_{i})<G\mid i\in\{1,\ldots,r\}\}$
(ii) $\Rightarrow$ (iii): Combining the assumption with Proposition 3.1, we obtain that the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-invertible. Hence, there exists a reduced set of subgroups in $G$ with $\mathcal{H}^{\prime}<\mathcal{H}$ and $J_{G/\mathcal{H}^{\prime}}]=[J_{G/\mathcal{H}}]$ that satisfies (iii-$\alpha$) or (iii-$\beta$) in Theorem 8.4. If (iii-$\alpha$) holds, denote by $\mathbf{K}^{\prime}=K^{\prime}$ the intermediate field of $\widetilde{K}/k$ corresponding to $N$. Then, $K^{\prime}$ is a subalgebra of $\mathbf{K}$ satisfying (iii-a) in Theorem 1.5. On the other hand, if (iii-$\beta$) is true, write $\mathcal{H}=\{H,H^{\prime}\}$, and denote by $K^{\prime}_{1}$ and $K^{\prime}_{2}$ the intermediate fields of $L/k$ corresponding to $K^{\prime}_{1}$ and $K^{\prime}_{2}$, respectively. Let $\mathbf{K}^{\prime}:=K^{\prime}_{1}\times K^{\prime}_{2}$, then it is a subalgebra of $\mathbf{K}$ satisfying (iii-b) in Theorem 1.5.
(iii) $\Rightarrow$ (i): We may assume $\mathbf{K}=\mathbf{K}^{\prime}$. If $\mathbf{K}$ satisfies (iii-a), then (iii-$\alpha$) in Theorem 8.4 is valid for $\mathcal{H}$. Hence, $J_{G/\mathcal{H}}$ is quasi-permutation by Proposition 5.1. On the other hand, if $\mathbf{K}$ satisfies (iii-b), then (iii-$\beta$) in Theorem 8.4 holds for $\mathcal{H}$. Consequently, Theorem 6.1 implies that $J_{G/\mathcal{H}}$ is quasi-permutation. Therefore, the assertion follows from Proposition 3.1. $\blacksquare$
9. Proof of Theorem 1.9
In this section, we assume that $k$ is a global field. For a finite étale algebra $\mathbf{K}$ over $k$, let
$\Sha(\mathbf{K}/k):=(\operatorname{N}_{\mathbf{K}/k}(\mathbb{A}_{\mathbf{K}}^{\times})\cap k^{\times})/\operatorname{N}_{\mathbf{K}/k}(\mathbf{K}^{\times}).$
Here, $\mathbb{A}_{\mathbf{K}}^{\times}$ is the product of the idèle groups of all the factors of $\mathbf{K}$. We say that the multinorm principle holds for $\mathbf{K}/k$ if
$\Sha(\mathbf{K}/k)=1.$
One can rephrase $\Sha(\mathbf{K}/k)$ by means of multinorm one tori. For an algebraic torus $T$ over $k$, the Tate–Shafarevich group of $T$ is defined as follows:
$^{1}(k,T):=\operatorname{Ker}\left(H^{1}(k,T)\xrightarrow{(\operatorname{Res}_{k_{v}/k})_{v}}\bigoplus_{v}H^{1}(k_{v},T)\right).$
Here, $v$ runs through all places of $k$, and $k_{v}$ denotes the completion of $k$ at $v$ (see also [PR94, Section 6.3]).
Lemma 9.1 ([LOY24, p. 7, (2.5)]).
Let $k$ be a global field, and $\mathbf{K}$ a finite étale algebra over $k$. Then, there is an isomorphism
[…]
Note that Lemma 9.1 is a generalization of Ono’s theorem, which is given in [Ono63, p. 70].
Proposition 9.2 ([Vos69, p. 1213, Theorem 5]).
Let $k$ be a global field, and $T$ an algebraic torus over $k$. Then, there is an exact sequence
$0\rightarrow A_{k}(T)\rightarrow H^{1}(k,\operatorname{Pic}(\overline{X}))^{\vee}\rightarrow{}^{1}(k,T)\rightarrow 0.$
Here, $\operatorname{Pic}(\overline{X})^{\vee}$ and $A_{k}(T)$ is defined as follows:
  • $\overline{X}:=X\otimes_{k}k^{\operatorname{sep}}$, where $X$ is a smooth compactification of $T$ over $k$;
  • $\operatorname{Pic}(\overline{X})^{\vee}$ is the Pontryagin dual of $\operatorname{Pic}(\overline{X})$;
  • $A_{k}(T):=(\prod_{v}T(k_{v}))/\overline{T(k)}$ is the defect of the weak approximation of $T$.
Proof of Theorem 1.9.
By Theorem 1.8, $T_{\mathbf{K}/k}$ is stably rational over $k$. In particular, if we set $G:=\operatorname{Gal}(K_{1}K_{2}/k)$, then Proposition 2.12 implies that the $G$-lattice $X^{*}(T_{\mathbf{K}/k})$ is quasi-permutation. Hence, we have $H^{1}(k,\operatorname{Pic}(\overline{X}))=0$ by Corollary 2.13. Combining this result with Proposition 9.2, we obtain $^{1}(k,T_{\mathbf{K}/k})=0$. Now, the assertion follows from Lemma 9.1. $\blacksquare$